### v2

# Bourbaki degree of pairs of projective surfaces

Abstract
The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials - defined as the kernel of the associated Jacobian matrix - naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation. As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
The present work focuses on sequences of two homogeneous polynomials in four variables. We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis. The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation. We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants. Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly-free cases. In particular, one of the nearly-free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
1 Introduction
Let $R=\kappa[x_{0},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$ variables with coefficients in an algebraically closed field $\kappa$. For an algebraically independent sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials in $R$ with degrees $d_{1}+1,\ldots,d_{k}+1$, respectively, with $d_{1}\leq\ldots\leq d_{k}$ one can consider the Jacobian matrix as a map of locally free sheaves on $\mathbb{P}^{n}\doteq\operatorname{Proj}R$
$\nabla\sigma=\begin{pmatrix}\nabla f_{1}\\ \nabla f_{2}\\ \vdots\\ \nabla f_{k}\end{pmatrix}:\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i}).$
The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$. It is considered (see [11]) in analogy with the case of the tangent logarithmic sheaf associated to hypersurfaces in $\mathbb{P}^{n}$. A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
In the case $k=1$ of divisors on $\mathbb{P}^{n}$, one has a short exact sequence of the form:
$0\rightarrow{\mathcal{T}}_{f}\rightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\mathcal{I}_{J_{f}}(d)\rightarrow 0,$
where $J_{f}=\langle\partial_{0}f,\ldots,\partial_{n}f\rangle\trianglelefteq R$ is the Jacobian ideal of the homogeneous polynomial $f$. The sheaf ${\mathcal{T}}_{f}$ is the sheaf of $\mathcal{O}_{\mathbb{P}^{n}}$-modules associated to the graded $R$-module $\operatorname{Syz}(J_{f})$ of Jacobian syzygies of $f$. For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated to the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
One should consider sequences rather than the associated subvarieties $X=V(\sigma)$ if $k>1$, because of the following observation: different choices of sequences $\sigma^{\prime},\sigma$ generating the same ideal may have different Jacobian syzygy modules (see [11, Example 2.7]). When $k=1$, however, any different choice $f^{\prime}$ for generator of the ideal $\langle f\rangle$ gives a linear multiple of $\nabla f$, so that ${\mathcal{T}}_{f}\simeq{\mathcal{T}}_{f^{\prime}}$. Similarly, when $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and in particular ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [11, Lemma 2.14]).
We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$. Following the terminology of [11], we say that a sequence $\sigma=(f_{1},\ldots,f_{k})$ is compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$. This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$. Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [11, Lemma 2.8]).
One may also consider the relationship with Bourbaki ideals, as in [17] and [7]. For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{n}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B_{\nu}}(e-d)\rightarrow 0$
where $d=-c_{1}({\mathcal{T}}_{\sigma})$ and $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection, since ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank two. The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice. In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [17]. By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. The focus of this work is to understand the above concept for $n=3$ and $k=2$.
For $k=1$, the singular scheme of $V(f)$, defined by the Jacobian ideal $\Sigma_{f}\doteq V(J_{f})\subset\mathbb{P}^{n}$, plays an important role in this study. If $s=\dim\Sigma_{f}$, the Hilbert polynomial is given by
$H(\mathcal{O}_{J_{f}}(d),t)=H(R/J_{f},t)=\frac{\deg(\Sigma)}{s!}t^{s}+O(t^{s-1}),$
and, by definition, the degree $\deg(\Sigma)$ is the leading coefficient above. Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$. For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [17]):
$\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}),$
where $e=\operatorname{indeg}({\mathcal{T}}_{f})$.
Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for ${\mathcal{T}}_{\sigma}$ to get an ideal sheaf as the cokernel. This problem is addressed in the recent work [7] where the authors consider tame hypersurfaces and other notions of Bourbaki degrees. In this paper, we focus on the case of sequences $\sigma=(f,g)$ on $\mathbb{P}^{3}$, so our logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has also rank two, and we consider the same approach as in [17], since we only have to choose one syzygy to do the construction.
We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$. In [11], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$. In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$, and $c_{1}(\mathcal{Q}_{\sigma})$ coincides with the degree of the greatest common divisor among all $(k\times k)$-minors of $\nabla\sigma$. The generic case is when $c_{1}(\mathcal{Q}_{\sigma})=0$, and this $g.c.d.$ is one, in which case we call $\sigma$ a normal sequence (see 2.1).
If we assume $c_{1}(\mathcal{Q}_{\sigma})=0$, we obtain a Hilbert polynomial of the form
$H(\mathcal{Q}_{\sigma},t)=\frac{\deg(\mathcal{Q}_{\sigma})}{s!}t^{s}+O(t^{s-1}),$
where $s\leq n-2$. We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ has codimension at least three.
In the case $n=3$ and $k=2$, we have sequences $\sigma=(f,g)$ of homogeneous polynomials with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, setting $d\doteq d_{f}+d_{g}$, so our objects are pairs of projective surfaces. We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and thus $c_{1}({\mathcal{T}}_{\sigma})=-d$. Since ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we may reproduce analogous considerations as in [17]. We show a formula for the Bourbaki degree of $\sigma$ given by
$\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma),$
where $m(\sigma)=\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$ is the degree of the sheaf $\mathcal{Q}_{\sigma}$ as defined above. The discrete invariant $m(\sigma)$ plays the role of the Tjurina number in the previous context. We compare the different scheme structures between $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ in Example 2, considering a free sequence $\sigma=(f,g)$ of polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with $d_{f}=1,d_{g}=2$ and so that $m(\sigma)=5$, but $\deg(\Xi_{\sigma})=6$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))=4$ as non-reduced schemes, showing the relationship between $\deg(\mathcal{Q}_{\sigma})$ and $\deg(\operatorname{supp}(\mathcal{Q}_{\sigma}))$ is more complicated than in the case $k=1$, where these two coincide for the sheaf $\mathcal{O}_{J_{f}}(d)$.
About the initial degree and the invariant $m=m(\sigma)$, we show some bounds that are analogous to the case of plane projective curves:
Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
  • (a)
    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
  • (b)
    $m(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$;
  • (c)
    The following are equivalent:
    • (1)
      $m(\sigma)=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$;
    • (2)
      $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$;
    • (3)
      ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$;
    • (4)
      $\sigma$ is compressible;
We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 3.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 9): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable. We also show the bound $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, which is attained for example for regular pencils with singular members having only isolated singularities (see Remark 10).
In Section 3.2, we relate our notion to the notion of the Bourbaki degree for a projective plane curve, by considering a reduced polynomial $g\in\kappa[x_{0},x_{1},x_{2}]$ together with its associated surface $V(g)\subset\mathbb{P}^{3}$ and projective curve $X=V(g)\subset\mathbb{P}^{2}$, to obtain $\operatorname{Bour}(X)=\operatorname{Bour}(g,x_{3})$.
Afterwards, in Section 3.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is a Bourbaki scheme obtained by the choice of a minimal syzygy. Related to the free resolutions, we introduce the notions of nearly-free sequences and $3$-syzygy sequences. We have the following chain of implications:
$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3-\text{syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
and the converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$.
In Section 4, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [11, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections. For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [17, Corollary 2.11]). The main theorem of the section is:
Theorem B. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials with degrees $d_{f}+1,d_{g}+1$, respectively. Then:
  • (a)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, then $\operatorname{Bour}(\sigma)\in\{0,1,2\}$;
  • (b)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then $\operatorname{Bour}(\sigma)\leq 5$.
In Section 5, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$. In these two classes, we classify all free and nearly free cases in terms of their discrete invariants and establish some stability results, which are summarized below:
Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 12$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=12,9$ or $8$, corresponding to $e$ being $0,1$ or $2$, respectively;
  • (c)
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 41), which is strictly $\mu$-semistable;
  • (d)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 16), but the bound above for stability may not be sharp.
Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 7$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=7$ or $5$, and each corresponds to $e$ being 0 or $1$, respectively;
  • (c)
    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 54 and Example 15);
  • (d)
    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Example 15 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable. This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [2]:
Conjecture. If the tangent sheaf of a codimension one foliation on $\mathbb{P}^{3}$ is not split, then it is $\mu$-semistable.
Along the text, we describe examples developed computationally with aid of Macaulay2 software ([12]) for the two families of sequences.
Acknowledgments
FM is financed by the São Paulo Research Foundation (FAPESP), Brasil. Process Number #2021/10550-4. FM is currently a PhD. student under the advice of Marcos Jardim and Daniele Faenzi, with partial funding by the Bridges ”Brazil-France interplays in Gauge Theory, extremal structures and stability” projects ANR-21-CE40-0017 and ANR-17-EURE-0002. We thank Marcos Jardim, Daniele Faenzi and Alan Muniz for fruitful discussions and suggestions regarding this work.
2 Basic setting and $m$-syzygy sequences on $\mathbb{P}^{n}$
We work over an algebraically closed field $\kappa$ of characteristic zero, although most of the definitions can be considered under relaxed assumptions. We denote by $R=\kappa[x_{0},\ldots,x_{n}]$ the graded ring of polynomials over $\kappa$ and $\mathbb{P}^{n}\doteq\mathbb{P}^{n}_{\kappa}=\operatorname{Proj}R$ the projective $n$ space over $\kappa$.
In this section, we recall from [11] the definition of the logarithmic tangent sheaf associated to a sequence of $k$ polynomials, and introduce some notation inspired by the literature about the case $k=1$.
For a sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials with degrees $d_{1}+1,\ldots,d_{k}+1$, we consider the Jacobian matrix $\nabla\sigma$ as a morphism between locally free sheaves:
$\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\xrightarrow{\nabla\sigma}\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i})$
so the kernel sheaf ${\mathcal{T}}_{\sigma}\doteq\ker(\nabla\sigma)$, is called the logarithmic tangent sheaf associated to $\sigma$. We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel sheaf and by $\mathcal{M}_{\sigma}\doteq\operatorname{im}(\nabla\sigma)$ the image sheaf. We denote by $d\doteq\sum_{i=1}^{n}d_{i}$ the total degree of a sequence $\sigma$. Assuming the generic rank of $\operatorname{im}(\nabla\sigma)$ is $k$, it follows that ${\mathcal{T}}_{\sigma}$ is a reflexive sheaf of rank n+1kn+1-k.
2.1 Normal sequences
By construction, the reduced support of the sheaf $\mathcal{Q}\sigma$ coincides with the vanishing locus $\Xi\sigma\doteq V(\bigwedge^{k}\nabla\sigma)$ of the $k\times k$-minors of the matrix $\nabla\sigma$. Assuming $\sigma$ is algebraically independent, the generic rank of $\nabla\sigma$ is $k$, so that $\operatorname{{codim}}\Xi_{\sigma}\geq 1$.
Moreover, since $\mathcal{Q}\sigma$ is a torsion sheaf, the condition $c_{1}(\mathcal{Q}\sigma)=0$ is equivalent to $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$. In terms of the matrix $\nabla\sigma$, $\operatorname{{codim}}\Xi_{\sigma}\geq 2$ if and only if the greatest common divisor among all $(k\times k)$-minors is one. When $k=1$, we have
$\operatorname{{codim}}(Z_{f})=\operatorname{{codim}}(\mathcal{Q}_{f})\geq 2\iff V(f)\text{ is normal}.$
Inspired by the behavior above, we call a sequence $\sigma$ normal when $\operatorname{{codim}}(\operatorname{supp}(\mathcal{Q}_{\sigma}))\geq 2$.
We note that the two schemes $\operatorname{supp}(\mathcal{Q}_{\sigma})$ and $\Xi_{\sigma}$ may have different scheme structures and, in particular, different degrees. On one hand, the scheme structure of the support $\operatorname{supp}(\mathcal{Q}_{\sigma})=V(\operatorname{Ann}\mathcal{Q}_{\sigma})$ is induced by the annihilator ideal sheaf of $\mathcal{Q}_{\sigma}$, defined locally as the annihilator ideal of the associated module. On the other hand, the scheme structure of $\Xi_{\sigma}$ is induced locally by the 0-th Fitting ideal of the $\mathcal{Q}_{\sigma}$. These schemes have the same reduced locus, since the support of these two ideals coincide, but in general $\operatorname{Fitt}_{0}\subseteq\operatorname{Ann}$ (see, for example, [9, Proposition 20.7]).
For $k=1$, when $\sigma=f$, we have an isomorphism $\mathcal{Q}f\simeq\mathcal{O}{Z_{f}}$. This holds because $\mathcal{M}f\simeq\mathcal{I}{Z_{f}}(d)$ is the ideal sheaf of the scheme $Z_{f}$, and whenever $f$ is reduced, the schemes $\Xi_{f}$ and $\operatorname{supp}(\mathcal{Q}{\sigma})$ coincide.
If $\sigma$ is a normal sequence, then the Hilbert polynomial of the cokernel sheaf $\mathcal{Q}_{\sigma}$ is of the form
$\operatorname{Hilb}(\mathcal{Q}_{\sigma},t)=\frac{m(\sigma)}{(n-2)!}t^{n-2}+\ldots$
where $m(\sigma)\geq 0$ is a discrete invariant so that $m(\sigma)=0$ if and only if $\operatorname{{codim}}(\mathcal{Q}_{\sigma})\geq 3$. However, $m(\sigma)$ generally coincides with neither the schematic degree of $\Xi_{\sigma}$ nor that of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, as we explore in Example 2 at the end of this section.
2.2 $m$-syzygy sequences and exponents
We may consider a minimal free resolution of the sheaf ${\mathcal{T}}_{\sigma}$, which will be given in the following form:
$\ldots\rightarrow\bigoplus_{i=1}^{m}\mathcal{O}_{\mathbb{P}^{n}}(-e_{i})\xrightarrow{\rho}{\mathcal{T}}_{\sigma}\rightarrow 0.$
In which case, we say $\sigma$ is an $m$-syzygy sequence.
We note that $m\geq n$ and $\sigma$ is said to be free if and only if ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, or, equivalently, when $m=n$.
The integers $e_{1}\leq\ldots\leq e_{m}$ will be called exponents, and the first $\operatorname{indeg}({\mathcal{T}}_{\sigma})\doteq e_{1}=e$ will be called the initial degree of the sheaf ${\mathcal{T}}_{\sigma}$. This is the minimum integer $e\geq 0$ such that there is a non-trivial syzygy $\nu$ of degree $d$ for the Jacobian matrix $\nabla\sigma$. Since there is an injection ${\mathcal{T}}_{\sigma}\hookrightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}$ and the latter is a $\mu$-semistable sheaf, all exponents satisfy $e_{i}\geq 0$.
A sequence $\sigma=(f_{1},\ldots,f_{k})$ is called compressible if, after a linear change of variables, there is a variable that does not occur in any of the $f_{1},\ldots,f_{k}$. The number of variables omitted in this way will give rise to syzygies of degree zero for $\nabla\sigma$, so the following hold (for proofs, see [11, Lemma 2.7 and Lemma 2.8]):
Lemma 1.
The following quantities coincide:
  • The number of variables omitted in $f_{1},\ldots,f_{k}$;
  • The discrete invariant $h^{0}({\mathcal{T}}_{\sigma})$;
  • The number of 0’s in a vector of exponents for ${\mathcal{T}}_{\sigma}$.
In particular, $e=0$ if and only if $\sigma$ is compressible.
For $n=3$ and when $\sigma=(f,g)$ is a normal sequence, we show the conditions above are equivalent to $m(\sigma)$ attaining its maximum value, see Section 3.1.
Example 2.
Let $\sigma=(2x_{1}x_{3}-x_{1}^{2},3x_{2}x_{3}^{2}-3x_{0}x_{1}x_{3}+x_{1}^{3})$. The Jacobian matrix is of the form
$\begin{pmatrix}0&-2x_{1}+2x_{3}&0&2x_{1}\\ -3x_{1}x_{3}&3x_{1}^{2}-3x_{0}x_{3}&3x_{3}^{2}&-3x_{0}x_{1}+6x_{2}x_{3}\\ \end{pmatrix}$
and the matrix below
$\begin{pmatrix}x_{3}&x_{0}x_{1}-x_{1}^{2}-2x_{2}x_{3}\\ 0&-x_{1}x_{3}\\ x_{1}&-2x_{2}x_{3}\\ 0&-x_{1}x_{3}+x_{3}^{2}\end{pmatrix}$
gives two linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
In this case, the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$ and the 0-th Fitting ideal are different, given by:
$\begin{cases}\operatorname{Ann}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3},x_{0}x_{1}^{2}-x_{1}^{3})=(x_{1},x_{3})^{2}\cap(x_{3},x_{0}-x_{1})\\ \operatorname{Fitt}_{0}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3}^{2},x_{1}^{2}x_{3},x_{0}x_{1}^{2}-x_{1}^{3}-2x_{1}x_{2}x_{3}+2x_{2}x_{3}^{2}),\end{cases}$
so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively. However, from the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$ we conclude $m(\sigma)=5$.
3 The Bourbaki degree of pairs of projective surfaces
In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials. We start with definitions and first results (3.1), followed by a reduction to the case of a projective plane curve (3.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (3.3), in particular introducing the class of nearly-free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
3.1 Framework and first results
By a sequence $\sigma=(f,g)$, unless otherwise stated, we mean an algebraically independent sequence of two homogeneous polynomials in $R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$. We also work only with normal sequences, as defined previously in 2.1. By curve we mean a locally Cohen-Macaulay closed subscheme of $\mathbb{P}^{3}$ of pure dimension one.
Each sequence $\sigma=(f,g)$ induces a morphism of sheaves on $\mathbb{P}^{3}$ by the Jacobian matrix:
$\nabla\sigma:\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d_{f})\oplus\mathcal{O}_{\mathbb{P}^{3}}(d_{g}),$
and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$. We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [11].
The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
Lemma 3.
Let $\mathcal{Q}$ be a coherent sheaf on $\mathbb{P}^{3}$ with $\operatorname{{rk}}(\mathcal{Q})=0$ and $c_{1}(\mathcal{Q})=0$. Then the Hilbert polynomial of $\mathcal{Q}$ is given by
$\mathcal{X}(\mathcal{Q}(t))=\operatorname{ch}_{2}(\mathcal{Q})t+\operatorname{ch}_{3}(\mathcal{Q})+2\operatorname{ch}_{2}(\mathcal{Q}).$
Proof.
Follows from direct application of the Hirzebruch-Riemann-Roch theorem. ∎
We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$. From the previous formula, $m(\sigma)$ is non-negative, and it is zero if and only if the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ is constant, that is, if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
Any saturated syzygy of the Jacobian matrix $\nu$, of degree $e\in\mathbb{Z}$, induces a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described by the Serre correspondence (see [14, Theorem $4.1$]). Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$. Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $e\geq 0$. The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $e$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
Proposition 4.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
For any saturated syzygy $\nu\in H^{0}({\mathcal{T}}(e))$ of degree $e\geq 0$, let $B_{\nu}\subset\mathbb{P}^{3}$ be the associated closed subscheme of pure dimension one. Then, we have the following equation:
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+m_{0}-m(\sigma),$
where $m_{0}\doteq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$.
Proof.
From the hypothesis $c_{1}({\mathcal{Q}}_{\sigma})=0$, we conclude $c_{1}({\mathcal{T}}_{\sigma})=-(d_{f}+d_{g})$ and $\operatorname{ch}_{2}({\mathcal{Q}}_{\sigma})=m(\sigma)$. Moreover, from additivity of $\operatorname{ch}_{2}$ on short exact sequences:
$\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$ $=-\operatorname{ch}_{2}({\mathcal{O}}_{\mathbb{P}^{3}}(d_{f})\oplus{\mathcal{O}}_{\mathbb{P}^{3}}(d_{g}))+\operatorname{ch}_{2}({\mathcal{Q}}_{\sigma})$
$=-\frac{d_{f}^{2}+d_{g}^{2}}{2}+m(\sigma).$
We can relate the Chern character and the Chern classes by the formula
$m(\sigma)$ $=\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})+\frac{d_{f}^{2}+d_{g}^{2}}{2}=\frac{c_{1}^{2}({\mathcal{T}}_{\sigma})-2c_{2}({\mathcal{T}}_{\sigma})}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
$=\frac{(d_{f}+d_{g})^{2}}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}-c_{2}({\mathcal{T}}_{\sigma})$
$=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-c_{2}({\mathcal{T}}_{\sigma}),$
so that
$c_{2}({\mathcal{T}}_{\sigma})=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma).$ (1)
On the other hand, since $c_{2}({\mathcal{T}}_{\sigma}(e))=\deg(B_{\nu})$ and ${\mathcal{T}}_{\sigma}$ is reflexive of rank $2$, we have
$\deg(B_{\nu})=c_{2}({\mathcal{T}}_{\sigma}(e))=c_{2}({\mathcal{T}}_{\sigma})+e\cdot c_{1}({\mathcal{T}}_{\sigma})+e^{2}.$
From this, together with 1, we obtain
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
We define the Bourbaki degree of a sequence $\sigma$, inspired by [17, Definition 2.4].
Definition 5.
Let $\sigma$ be a normal sequence and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. The Bourbaki degree of a normal sequence $\sigma$ is defined by:
$\operatorname{Bour}(\sigma)\doteq\deg(B_{\nu})=e(e-d)+m_{0}-m(\sigma),$
for some non-trivial syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. We note that every non-trivial syzygy of minimal degree is saturated.
Remark 6.
It follows from construction that the sequence $\sigma$ is free if and only if $\operatorname{Bour}(\sigma)=0$.
Remark 7.
From [13, Proposition 4.1], we have the formula
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
relating the third Chern class of ${\mathcal{T}}_{\sigma}$ and the discrete invariants of $B$, when $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $B$ is the zero locus of a non-zero section in $H^{0}({\mathcal{T}}_{\sigma}(e))$. Moreover, dualizing the sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B}(e-d)\rightarrow 0$
we conclude that the singular set of the sheaf ${\mathcal{T}}_{\sigma}$ is contained in $B$.
Now, we obtain some bounds for the quantities $\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ in terms of the degrees $d_{f},d_{g}$.
Theorem A. Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
  • (a)
    We have $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
  • (b)
    $m(\sigma)\leq m_{0}$;
  • (c)
    The following are equivalent:
    • (1)
      $m(\sigma)=m_{0}$;
    • (2)
      $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$;
    • (3)
      $\sigma$ is compressible;
    • (4)
      ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$;
Proof.
To show $(a)$, we build explicit syzygies of the Jacobian matrix $\nabla\sigma$ of degrees $d_{f}+d_{g}$, and at least one of them is nonzero. Writing the Jacobian matrix by
$\nabla\sigma=\begin{pmatrix}\partial_{0}f&\partial_{1}f&\partial_{2}f&\partial_{3}f\\ \partial_{0}g&\partial_{1}g&\partial_{2}g&\partial_{3}g\\ \end{pmatrix}$,
the following vectors
$\nu_{0}$ $=\begin{pmatrix}0\\ \partial_{2}f\partial_{3}g-\partial_{3}f\partial_{2}g\\ -\partial_{1}f\partial_{3}g+\partial_{3}f\partial_{1}g\\ \partial_{1}f\partial_{2}g-\partial_{2}f\partial_{1}g\end{pmatrix};\nu_{1}=\begin{pmatrix}\partial_{2}f\partial_{3}g-\partial_{3}f\partial_{2}g\\ 0\\ -\partial_{0}f\partial_{3}g+\partial_{3}f\partial_{0}g\\ \partial_{0}f\partial_{2}g-\partial_{2}f\partial_{0}g\end{pmatrix}$;
$\nu_{2}$ $=\begin{pmatrix}\partial_{1}f\partial_{3}g-\partial_{3}f\partial_{1}g\\ -\partial_{0}f\partial_{3}g+\partial_{3}f\partial_{0}g\\ 0\\ \partial_{0}f\partial_{1}g-\partial_{1}f\partial_{0}g\end{pmatrix};\nu_{3}=\begin{pmatrix}\partial_{1}f\partial_{2}g-\partial_{2}f\partial_{1}g\\ -\partial_{0}f\partial_{2}g+\partial_{2}f\partial_{0}g\\ \partial_{0}f\partial_{1}g-\partial_{1}f\partial_{0}g\\ 0\end{pmatrix}$.
are a syzygies of degree $d_{f}+d_{g}$. Since there is at least one nonzero $2\times 2$ minor, from the hypothesis of algebraically independent, at least one of the syzygies $\nu_{0},\ldots,\nu_{3}$ is nonzero.
For $(b)$, using that $e\leq d$, we can use that
$0\leq\operatorname{Bour}(\sigma)=e(e-d)+m_{0}-m(\sigma),$
so that $m(\sigma)\leq m_{0}+e(e-d)$, but $e(e-d)\leq 0$ and the claim follows.
For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [11, Lemma 2.7]. Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [11, Lemma 2.8]. The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
To show $(1)\Rightarrow(2)$, let $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. From the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=e(e-d)$. Since $\operatorname{Bour}(\sigma)\geq 0$ and from part (a) we have $e\leq d$, it follows that $e=0$ or $e=d$. Both cases mean $\operatorname{Bour}(\sigma)=0$ and the sequence is free. Since $c_{1}({\mathcal{T}}_{\sigma})=-d$ is additive, it follows that $e=0$, otherwise we would have $c_{1}({\mathcal{T}}_{\sigma})<-d$. ∎
Remark 8.
The bound obtained in 3.1, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory. We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [11, Section 9]). Let $C$ be the one-dimensional part of the singular scheme of this foliation. From the formulas of discrete invariants in [2, Theorem 3.1],
$c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C).$
Using the formula $c_{2}({\mathcal{T}}_{\sigma})=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$ and the equations
$c_{2}({\mathcal{T}}_{\sigma})-d+1=c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C)$
we obtain $m(\sigma)=\deg(C)-d-d_{f}d_{g}-1$, so from the bound above we get
$\deg(C)-d-d_{f}d_{g}-1=m(\sigma)$ $\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g},$
and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [20, Corollary 4.8].
From simple observations about the formula $\operatorname{Bour}(\sigma)$, we are able to obtain the following inequalities related to $\mu$-stability and freeness of the logarithmic sheaves.
Proposition 9.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$. Denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $d=d_{f}+d_{g}$. Then:
  • (a)
    $\operatorname{Bour}(\sigma)\leq m_{0}$, and equality holds if and only if $m(\sigma)=0$ and $e=d$.
  • (b)
    If $\operatorname{Bour}(\sigma)>(d_{f}-1)(d_{f}+d_{g})+d_{g}^{2}+1$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
  • (c)
    If
    $m(\sigma)<\frac{1}{2}\left(\frac{3d_{f}^{2}}{2}+\frac{3d_{g}^{2}}{2}+d_{f}d_{g}\right),$
    then $\sigma$ is not free.
Proof.
The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)\geq 0$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$, from the inequality $e\leq d$. This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility. The converse also follows simply from the formula.
By construction, $m(\sigma)\geq 0$, and therefore for a given value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, the Bourbaki degree of $\sigma$ can be at most
$\operatorname{Bour}(\sigma)\leq e(e-d)+m_{0},$
and this is a function $H=H(e)$ which attains its minimum at $e=d/2$. The function $H$ is decreasing on $e\in\{1,\ldots,d/2\}$, and if $e>d/2$, ${\mathcal{T}}_{\sigma}$ is $\mu$-stable, since $\mu({\mathcal{T}}_{\sigma})=-d/2$. Thus, the maximum value $\operatorname{Bour}(\sigma)$ in the range $e\in\{1,\ldots,d/2\}$ for any possible value of $m(\sigma)$ is $H(1)$, which is the expression above on the right-hand side. If $\operatorname{Bour}(\sigma)$ is higher than this, then $e>d/2$, and thus we obtain $(b)$.
With the same strategy as $(b)$, since the function $H(e)$ attains its minimum at $e=d/2$, for $m(\sigma)$ satisfying the inequality of the claim we obtain that
$\operatorname{Bour}(\sigma)=H(e)\geq H(d/2)=\frac{d^{2}}{4}-\frac{d^{2}}{2}+m_{0}-m(\sigma)>0,$
and therefore $\operatorname{Bour}(\sigma)\neq 0$, independently of the value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. ∎
Remark 10.
If $\sigma=(f,g)$ is a pencil of surfaces of the same degree, say $d_{f}=d_{g}=p$, then $m(\sigma)=0$ if $\sigma$ is a regular pencil and, moreover, every singular member $V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}$ is normal. This follows from the description
$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g))$
obtained in [11, Lemma 2.17], since normal projective surfaces have isolated singularities and regular sequences have only a finite number of singular members, therefore $\dim(\operatorname{supp}(\mathcal{Q}_{\sigma}))=0$.
Example 11.
We consider the sequence
$\sigma=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
where $f$ is a normal singular cubic with an $A_{1}$-singularity at $[0:0:0:1]$ and $g$ is a normal singular cubic with singularity type $2A_{1}A_{2}$.
Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a locally free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
Remark 12.
We have not found any examples of sequences with $m(\sigma)=0$ so that $e<d$, which motivates the conjecture: does $m(\sigma)=0$ implies $e=d$?
3.2 The Bourbaki degree of a plane curve from the Bourbaki degree of a pair of surfaces
In this section, we obtain the formula of the Bourbaki degree of a projective plane curve $X=V(g)\subset\mathbb{P}^{2}$, where $g\in\kappa[x_{0},x_{1},x_{2}]$ is a reduced homogeneous polynomial, introduced in [17], as a special case of our formula, by considering the pair $\sigma=(x_{3},g)$ in $\mathbb{P}^{3}$.
Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$. We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$.
The matrix $\nabla\sigma$ will be given by
$\nabla\sigma=\begin{pmatrix}0&0&0&1\\ \partial_{0}g&\partial_{1}g&\partial_{2}g&0\end{pmatrix}$.
Let us denote by $\nabla\overline{g}=(\partial_{0}g,\partial_{1}g,\partial_{2}g)$ the vector, so we denote by
${\mathcal{T}}_{\overline{g}}\doteq\ker(\nabla\overline{g})\hookrightarrow\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}\xrightarrow{\nabla\overline{g}}\mathcal{O}_{\mathbb{P}^{3}}(d_{g})$
the kernel of the multiplication by this vector. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
${\mathcal{T}}_{\overline{g}}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}$$\mathcal{O}_{\mathbb{P}^{3}}(d)$$\mathcal{O}_{Z_{g}}(d)$${\mathcal{T}}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}$$\mathcal{O}_{\mathbb{P}^{3}}(d)\oplus\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{Q}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{O}_{\mathbb{P}^{3}}$$\scriptstyle{\nabla\overline{g}}$$\scriptstyle{\nabla\sigma}$$\scriptstyle{\cdot 1}$
and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$.
From the discrete invariants considered $(d_{f}=0)$, we have the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
$\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}(\ker\nabla g(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
where $\tau(X)$ is the Tjurina number of the curve $X$, since $d_{f}=0$.
From the construction, we have that the singular scheme of ${\mathcal{T}}_{\sigma}$ (described by $Z_{g}$) intersects $V(x_{3})$ transversely, so we obtain the formula between Chern classes $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:\mathbb{P}^{2}\simeq V(x_{3})\hookrightarrow\mathbb{P}^{3}$ is the inclusion. Therefore,
$1-d[H]+(m_{0}-m(\sigma))[H]^{2}$ $=i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$
$=1-d[H]+(d^{2}-\tau(X))[H]^{2},$
and since $m_{0}=d_{g}^{2}$, we obtain the equality $m(\sigma)=\tau(X)$.
Moreover, doing the analogous comparison for the total classes of the twist ${\mathcal{T}}_{\sigma}(e)$, we obtain
$i^{*}c({\mathcal{T}}_{\sigma}(e))=1+(e-d)[H]+\operatorname{Bour}(\sigma)[H]^{2}=1+(e-d)[H]+\operatorname{Bour}(X)[H]^{2}=c(i^{*}{\mathcal{T}}_{\sigma}(e)),$
and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$.
Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number, and thus it should be the correct generalization for the case of pairs of surfaces.
3.3 Locally free resolutions and the Bourbaki scheme
In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [17, Theorem 2.1, (c)]. This follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$.
Lemma 13.
Let $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ be a non-zero section with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $B\subset\mathbb{P}^{3}$ be the pure codimension $2$ subscheme associated to $\nu$ in a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\xrightarrow{\pi}\mathcal{I}_{B}(e-d)\rightarrow 0.$
Then:
  • (a)
    Every free resolution for $\mathcal{I}_{B}$:
    $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\xrightarrow{\omega}\mathcal{I}_{B}\rightarrow 0$
    lifts for a free resolution of the form
    $0\rightarrow F_{2}(e-d)\rightarrow F_{1}(e-d)\rightarrow F_{0}(e-d)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\omega(e-d),\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0$
    for ${\mathcal{T}}_{\sigma}$.
  • (b)
    For a minimal free resolution of ${\mathcal{T}}_{\sigma}$ including the section $\nu$:
    $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\lambda,\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0,$
    it induces a free resolution for $\mathcal{I}_{B}$ of the form:
    $0\rightarrow F_{2}(d-e)\rightarrow F_{1}(d-e)\rightarrow{}F_{0}(d-e)\xrightarrow{\lambda(d-e)}\mathcal{I}_{B}\rightarrow 0.$
Proof.
To show $(a)$, we apply the functor $\operatorname{Hom}(F_{0}(e-d),-)$ to the short exact sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\xrightarrow{\pi}\mathcal{I}_{B}(e-d)\rightarrow 0,$
to get the exact piece:
$\operatorname{Hom}(F_{0}(e-d),{\mathcal{T}}_{\sigma})\xrightarrow{\pi^{*}}\operatorname{Hom}(F_{0}(e-d),\mathcal{I}_{B}(e-d))\rightarrow\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))=0,$
since $\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(F_{0}^{\vee}(-2e-d))=0$, as $F_{0}^{\vee}$ is a direct sum of line bundles and these have vanishing first cohomology in $\mathbb{P}^{3}$. Thus, $\pi^{*}$ is surjective, and there is a morphism $\tilde{\omega}:F_{0}(e-d)\rightarrow{\mathcal{T}}_{\sigma}$ such that $\pi\circ\omega(e-d)=\tilde{\omega}$. We now consider the map $\omega(e-d)\oplus\nu$ in the following commutative diagram with short exact sequences as the central two columns:
[…]
From the snake lemma, we obtain that $\operatorname{coker}(\tilde{\omega}\oplus\nu)=0$ and that $\ker(\tilde{\omega}\oplus\nu)\simeq\ker(\omega(e-d))$. Thus, we can continue the resolution for $\mathcal{I}_{B}$, twisting by $\mathcal{O}_{\mathbb{P}^{3}}(e-d)$, to obtain the following free resolution:
[…]
for ${\mathcal{T}}_{\sigma}$, as claimed.
To show $(b)$, we consider the diagram with exact rows induced by the fact above to obtain the short exact sequence in cokernels as the third row below:
[…]
Completing to the resolution and twisting accordingly, we obtain the resolution from the claim below.
[…]
Definition 14.
Let $\sigma$ be a non-free normal sequence with degrees $d_{f}+1,d_{g}+1$. We say $\sigma$ is:
  • nearly free if $\operatorname{Bour}(\sigma)=1$.
  • 33-syzygy if there is a minimal free resolution for ${\mathcal{T}}_{\sigma}$ such that $\operatorname{{rk}}(F_{0})=2$ in the notation of Lemma 13, $(b)$.
Example 15 (Nearly free sequence with $d_{f}=1,d_{g}=2$).
We consider the following sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(x_{0}^{2}+x_{3}^{2},x_{0}^{3}+x_{0}x_{1}x_{2}+x_{3}^{3})$
with Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 3x_{0}^{2}+x_{1}x_{2}&x_{0}x_{2}&x_{0}x_{1}&3x_{3}^{2}\end{pmatrix}$.
The 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the $1$-st Fitting ideal has dimension zero.
The minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=1$, $m(\sigma)=4$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=3$.
Example 16 ($3$-syzygy pencil of cubics which is not nearly-free).
Considering the following pencil of cubics:
$\sigma=(x_{2}x_{3}(x_{0}-x_{1}),x_{0}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}))$
where $f$ is a hyperplane arrangement and $g$ is the union of a plane and a smooth quadric, with the Jacobian matrix:
$\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{0}x_{3}-x_{1}x_{3}&x_{0}x_{2}-x_{1}x_{2}\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
In this case, the 0-th Fitting ideal coincides with the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$, and the saturation of the first Fitting ideal is zero. The primary decomposition of $\Xi_{\sigma}$ is described in the following table:
dimension degree radical ideal
1 1 $(x_{2},x_{3})$
1 1 $(x_{3},x_{0}-x_{1})$
1 1 $(x_{2},x_{0}-x_{1})$
1 2 $(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for $\mathcal{I}_{B}$ will be of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}\rightarrow\mathcal{I}_{B}\rightarrow 0,$
presenting $B$ as a complete intersection of two quadric surfaces in $\mathbb{P}^{3}$.
For the rest of the section, we study some aspects of these special classes of sequences.
Proposition 17.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}+1$, $d_{g}+1$. Then:
  • (a)
    $\sigma$ is nearly free if and only if the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
    $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-1)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
    where $d=d_{f}+d_{g}$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
  • (b)
    if $\sigma$ is nearly-free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
Proof.
This follows from Lemma 13, using the minimal free resolution for a line in $l\subset\mathbb{P}^{3}$ as the intersection of two planes:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\rightarrow\mathcal{I}_{l}\rightarrow 0$
since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for the syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. ∎
Remark 18.
The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [5], related to rational cuspidal curves. In [17, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [5] (see [17, Definition 2.17]). Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
The notion of 33-syzygy divisors is also present in a number of works in the field, for example [1], [6] and [8].
Example 19.
Consider $f=x_{0}^{3}+x_{0}x_{1}x_{2}+x_{3}^{3}$ and $g=x_{0}^{k+1}+x_{3}^{k+1}$ for $k\geq 3$. Then $d=d_{f}+d_{g}=k+2$ and:
  • For $k=2$, $(f,g)$ is a free pencil of cubics with $e=1$, $m(\sigma)=9$;
  • For $k>2$, $(f,g)$ is nearly-free with $e=1$.
For $k=2$, we note that the matrix below:
$\begin{pmatrix}0&-x_{0}x_{3}^{2}\\ x_{1}&0\\ -x_{2}&x_{2}x_{3}^{2}\\ 0&x_{0}^{3}\end{pmatrix}$
gives trivializing syzygies such that ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
For $k>2$, we set
$M=\begin{pmatrix}0&-x_{0}x_{2}x_{3}^{k}&-x_{0}x_{1}x_{3}^{k}\\ x_{1}&-3x_{0}^{k}x_{3}^{2}+3x_{0}^{2}x_{3}^{k}&0\\ -x_{2}&x_{2}^{2}x_{3}^{k}&-3x_{0}^{k}x_{3}^{2}+3x_{0}^{2}x_{3}^{k}+x_{1}x_{2}x_{3}^{k}\\ 0&x_{0}^{k+1}x_{2}&x_{0}^{k+1}x_{1}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}^{k}x_{3}^{2}-x_{0}^{2}x_{3}^{k}\\ \frac{1}{3}x_{1}\\ -\frac{1}{3}x_{2}\end{pmatrix}$,
so we obtain a free resolution of ${\mathcal{T}}_{\sigma}$ given by
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-1)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)^{\oplus 2}\xrightarrow{M}\mathcal{T}_{\sigma}\rightarrow 0,$
and in particular $\sigma$ is a nearly-free sequence.
Proposition 20.
If a normal sequence $\sigma=(f,g)$ is a $3$-syzygy, then $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$. Moreover, a sequence $\sigma$ is $3$-syzygy if and only if $B_{\nu}$ is a complete intersection, for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
Proof.
First, if we assume $\sigma$ is $3$-syzygy, then there is a free resolution of the form:
$0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\lambda}{\mathcal{T}}_{\sigma}\rightarrow 0,$
so we split the resolution into two short exact sequences:
$F_{2}\hookrightarrow F_{1}\twoheadrightarrow S\text{ and }S\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma},$
and focus on the second one. The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [14, Proposition 1.1]. Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
$S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma}$
is a free resolution for ${\mathcal{T}}_{\sigma}$, concluding $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$.
For the equivalence stated above, if we start with a sequence $\sigma$ which is $3$-syzygy and then apply Lemma 13, $(b)$, we obtain a resolution for $\mathcal{I}_{B}$ which is of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
since $\operatorname{{rk}}(F_{0})=2$, thus concluding $B$ must be a complete intersection scheme. On the other hand, if $B$ is a complete intersection, then there is a resolution for $\mathcal{I}_{B}$ of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\xrightarrow{r,s}\mathcal{I}_{B}\rightarrow 0,$
induced by the two equations $r,s$ defining $B=V(r,s)\subset\mathbb{P}^{3}$. Thus, applying Lemma 13, $(a)$, we can lift the resolution above for a free resolution for ${\mathcal{T}}_{\sigma}$ of the form
$0\rightarrow F_{1}\rightarrow F_{0}^{\prime}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $\operatorname{{rk}}(F_{0}^{\prime})=3$, thus $\operatorname{{rk}}(F_{0})=2$ and $\sigma$ is $3$-syzygy. ∎
The following proposition follows from the formula in Remark 7 for $c_{3}({\mathcal{T}}_{\sigma})$ in terms of $e,\deg(B)$ and $p_{a}(B)$ for a Bourbaki scheme $B=B_{\nu}$.
Proposition 21.
If $\sigma=(f,g)$ is a nearly-free sequence such that ${\mathcal{T}}_{\sigma}$ is locally free, then $d=d_{f}+d_{g}$ must be even and ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e),$
assuming ${\mathcal{T}}_{\sigma}$ is locally free, we obtain $c_{3}({\mathcal{T}}_{\sigma})=0$. On the other hand, since $\sigma$ is nearly-free, $\deg(B)=1$ and $p_{a}(B)=0$, hence
$e=\frac{d+2}{2},$
which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$. ∎
Remark 22.
We note that there is a chain of implications:
$\sigma\text{ is nearly free }\Rightarrow\sigma\text{ is $3$-syzygy }\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
where the first follows from 20 since $\sigma$ is nearly free iff $B_{\nu}$ is a line for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and every line is a complete intersection of two planes. The converses do not hold, as we explore in the next examples: there are $3$-syzygy pencils of cubics which are not nearly free (Example 16), pencils of cubics which satisfy $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but are not $3$-syzygy (Example 23) and pencils of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (Example 24).
We also construct two pencils of cubics with the same discrete invariants $(m(\sigma),\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma))$ and the same Chern classes, which are distinguished by their homological behavior: one is $3$-syzygy and the other satisfies $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (see Example 23 and Example 25).
Example 23 (pencil of cubics which is not $3$-syzygy and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$).
Considering the following pencil of cubics:
$\sigma=(x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{2}^{3}+x_{1}x_{2}x_{3}+x_{3}^{3})$
with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\\ 0&x_{2}x_{3}&3x_{2}^{2}+x_{1}x_{3}&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$.
We may check computationally that the saturation of the ideal of $(2\times 2)$-minors coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$ and with the saturation of the 0-th Fitting ideal, therefore over this subset $\mathcal{Q}_{\sigma}$ has rank one. The $1$-dimensional part of this scheme has $3$ irreducible components, two lines $L_{1}=V(x_{0},x_{3}),L_{2}=V(x_{2},x_{3})$ and a quadric plane curve $Q=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})$. Moreover, $L_{1}$ and $Q$ have multiplicity two structure, adding to $m(\sigma)=5$.
This can also be obtained from the minimal free resolution of ${\mathcal{T}}_{\sigma}$:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{M}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. This implies that $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but we need $4$ syzygies to generate ${\mathcal{T}}_{\sigma}$.
Example 24 (pencil of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$).
We consider the sequence of cubics $d_{f}=d_{g}=2$ given by:
$\sigma=(x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3},x_{2}x_{3}(x_{2}-x_{1})),$
considered in [11, Theorem 8.1]. From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle. Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{L}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
So we compute $e=3$ and $m(\sigma)=7$. This may also be seen from the primary decomposition of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, which is composed of three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$, where the last line has a multiplicity $5$ structure, and at every line the rank of $\mathcal{Q}_{\sigma}$ is one.
Example 25 (pencils of cubics with the same discrete invariants and different homological behavior).
Considering the following pencil of cubics:
$\sigma=(x_{0}^{3}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{3}^{3}+x_{1}x_{3}^{2}+x_{0}x_{1}x_{3}+x_{0}^{2}x_{2}),$
with the associated Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}+x_{3}^{2}&x_{0}^{2}&x_{0}x_{1}+2x_{1}x_{3}+3x_{3}^{2}\end{pmatrix}$.
Here, although the 0-th Fitting ideal is different from the annihilator ideal of $\mathcal{Q}_{\sigma}$, their one-dimensional component coincide with a multiplicity five structure along the line $V(x_{3},x_{0})$, and thus $m(\sigma)=5$.
Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 6}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$ and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$. A resolution for $\mathcal{I}_{B}$ will be of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 6}\rightarrow\mathcal{I}_{B}\rightarrow 0.$
This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as 16, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
In Section 5 we characterize all nearly free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly free sequences with $d_{f}=1,d_{g}=2$.
4 Extreme cases of low initial degree
In this section, we observe that a non-zero section $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ induces a sub-foliation by curves of degree $e+1$ of the foliation ${\mathcal{T}}_{\sigma}(1)$. We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two. From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 28. We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example [4]).
As explored in [11, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
$0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the foliation. Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
[…]
The middle column of the previous diagram
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(1-e)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow G\rightarrow 0$
defines a foliation by curves of $\mathbb{P}^{3}$ of degree $e$, where $G$ is a rank two torsion-free sheaf and $G^{\vee}$ is called the conormal sheaf of the foliation. Dualizing this short exact sequence, we obtain
$0\rightarrow G^{\vee}\rightarrow\Omega^{1}_{\mathbb{P}^{3}}\rightarrow\mathcal{I}_{W}(e-1)\rightarrow 0,$
defining a subscheme $W\subset\mathbb{P}^{3}$, called the singular scheme of the associated foliation by curves. It has codimension at least two and it is also described by ${\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})\simeq\mathcal{O}_{W}$.
The classification of such foliations by curves of low degree provides the following bounds on their singular schemes:
  • (a)
    ([4, $2.3$, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1. In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
  • (b)
    (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$ or is zero-dimensional.
Theorem B. Let $\sigma=(f,g)$ be a normal sequence of polynomials of degrees $d_{f}+1,d_{g}+1$. Then:
  • (a)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, then $\operatorname{Bour}(\sigma)\in\{0,1,2\}$;
  • (b)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then $\operatorname{Bour}(\sigma)\leq 5$.
Proof.
To show $(a)$, we dualize the following short exact sequence, obtained above for $e=1$:
$0\rightarrow\mathcal{I}_{B}(2-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
to get a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-2)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+2)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}$ is defined by the short exact sequence $0\to\mathcal{U}\to\mathcal{O}_{\Gamma_{\sigma}}\to\mathcal{O}_{Y}\to 0$ and $Y$ is the one-dimensional component of $\Gamma\sigma$. Moreover, since $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$, we may consider the final piece of the long exact sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+2)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$, and from the classification $\deg(W^{\prime})\leq\deg(W)\leq 2$, hence the result follows.
For $(b)$, we proceed analogously, dualizing the sequence
$0\rightarrow\mathcal{I}_{B}(3-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0$
to obtain a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-3)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+1)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}\hookrightarrow\mathcal{O}_{\Gamma_{\sigma}}\rightarrow\mathcal{O}_{Y}$ and $Y$ is the one-dimensional component of $\Gamma_{\sigma}$. Moreover, $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$ and from the final piece of the sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+1)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{O}_{\Gamma_{\sigma}},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$. Since $W^{\prime}$ is a subscheme of $W$, we have $\deg(W^{\prime})\leq\deg(W)$. The classification of foliations by curves of degree $2$ implies $\deg(W)\leq 5$, which completes the proof. ∎
To be able to present the next applications, we will need the following result on the structure of degree $2$ space curves in $\mathbb{P}^{3}$:
Theorem 26.
[18, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree $2$ and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
  • (a)
    $a\geq-1$, and a=1a=-1 if and only if $B$ is planar;
  • (b)
    For $a\geq 1$, $B$ must be a multiplicity two structure at a line $L\subset\mathbb{P}^{3}$, and these satisfy a short exact sequence of the form
    $0\rightarrow\mathcal{O}_{L}(a)\rightarrow\mathcal{O}_{B}\rightarrow\mathcal{O}_{L}\rightarrow 0;$
  • (c)
    For $a\geq 1$, if $B$ is a multiplicity two structure on a line, then $\omega_{B}\simeq\mathcal{O}_{B}(-a-2)$. If $a=0$ and $B$ is a union of two skew lines, then $\omega_{B}\simeq\mathcal{O}_{B}(-2)$.
Proof.
Items $(a)$ and $(b)$ are in the original paper, and to obtain $(c)$, for $a\geq 1$ we can apply the functor $\operatorname{\mathcal{H}\textit{om}}(-,\omega_{\mathbb{P}^{3}})$ to the short exact sequence in $(ii)$, use that $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ and twist by $a+2$ to obtain a short exact sequence
$0\rightarrow\mathcal{O}_{L}(a)\rightarrow\omega_{B}(a+2)\rightarrow\mathcal{O}_{L}\rightarrow 0$
which coincides with the short sequence in $(b)$, hence we obtain the isomorphism in the claim. For the case $a=0$, $B$ must be either a multiplicity two structure as in $(b)$ with $a=0$ or a union of skew lines. In the first case, the proof follows as before. If $B=L_{1}\cup L_{2}$, then $\omega_{B}\simeq\omega_{L_{1}}\oplus\omega_{L_{2}}$ and the claim follows from the fact $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ in each component. ∎
Remark 27.
The results above can be used to restrict the possible values of numerical invariants. For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results above to show that the only possible discrete invariant is $c_{3}({\mathcal{T}}_{\sigma})=8$, although the existence of such an example remains an open question.
Indeed, from [19, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 7 we obtain $c_{3}({\mathcal{T}}_{\sigma})=8-2a$, where $p_{a}(B)=-1-a$. From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(4-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we obtain the Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(4-a))=10-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=10-2a,$
and additivity of Euler characteristic gives a contradiction whenever $a\neq 0$, so $c_{3}({\mathcal{T}}_{\sigma})=8$ if this occurs.
Using the classification of $\mu$-semistable reflexive sheaves on $\mathbb{P}^{3}$ with small $c_{2}$, we are able to obtain a full classification of the cases of pencils of cubics with $m(\sigma)=7$ and $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$ (see 42).
As another application, we are able to show that sequences $\sigma$ with mixed degrees $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$ cannot be unstable.
Proposition 28.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, and in particular ${\mathcal{T}}_{\sigma}$ must be $\mu$-stable.
Proof.
Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$. From Section 3.1, $(c)$, since $m(\sigma)=3\neq 7$, $\sigma$ is incompressible, and from [16, Theorem 1.1], we obtain $c_{3}\leq 4$ (for a more general formula, see 45). On the other hand, using Remark 7 we have $c_{3}({\mathcal{T}}_{\sigma})=6-2a$, so $a\in\{1,2,3\}$.
From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(3-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we get Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(3-a))=8-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=8-2a,$
which cannot coincide for the values $a\in\{1,2,3\}$, hence we obtain a contradiction. ∎
5 Pencils of cubics and degree 6 curves inside quadric surfaces
In this final section, we show some classification results for pencils of cubics and sequences $\sigma=(f,g)$ with $d_{f}=1$, $d_{g}=2$, which correspond to degree $6$ curves inside quadric surfaces.
The results are derived from the previous sections, Section 3.1 and Section 4, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [14], [21], [15], [3] and [16].
5.1 Pencils of cubics
In this section, we focus on the case of pencils of cubics, i.e., $d_{f}=d_{g}=2$. The Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ is given by the formula
$\operatorname{Bour}(\sigma)=e(e-4)+12-m(\sigma).$
When $e=1$, the formula above implies $m(\sigma)\leq 9$, from the inequality $\operatorname{Bour}(\sigma)\geq 0$.
The assumption $\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ enables us to apply the result [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}(1)$ and obtain the following upper bounds for $m(\sigma)$ and the third Chern class $c_{3}({\mathcal{T}}_{\sigma})$:
Proposition 29.
If $\sigma=(f,g)$ is a normal pencil of cubics with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then $m(\sigma)\leq 8$.
Proposition 30.
If $\sigma=(f,g)$ is a normal pencil of cubics and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then
  • (a)
    if $7\leq m(\sigma)\leq 8$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
  • (b)
    if $0\leq m(\sigma)<7$, then
    $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72.$
The following result is obtained using [15, Theorem 0.1] and the bound obtained in Section 3.1, $(a)$:
Proposition 31.
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 4$, and we have the following table of sharper bounds for each possibility of $m=m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
6 $e\leq 3$
7 $e\leq 3$
8 $e\leq 2$
The following result gives a picture of the generic case of a pencil of cubics.
Proposition 32.
Let $\sigma=(f,g)$ be a general pencil of cubics in $\mathbb{P}^{3}$. Then $m(\sigma)=0$, the number of singular members is $32$, and all singular members have one singular point, in particular $m(\sigma)=0$. Moreover $c_{3}({\mathcal{T}}_{\sigma})=32$. (see Example 11).
Proof.
This follows from intersection theory for the bundle of principal parts (see [10], Proposition $7.1$ and Proposition $7.4$). ∎
Rewriting Section 3.1, $(d)$ for pencils of cubics, we obtain:
Proposition 33.
Let $\sigma$ be a normal pencil of cubics. Then $\sigma$ is compressible if and only if $m(\sigma)=12$ (see Example 34).
Example 34 (Free, compressible pencil of cubics).
Consider the sequence $\sigma=(x_{0}^{3}+x_{1}^{3}+x_{0}x_{1}x_{3},x_{0}x_{1}x_{3})$. This sequence is independent of the variable $x_{2}$, with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&3x_{1}^{2}+x_{0}x_{3}&0&x_{0}x_{1}\\ x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}\end{pmatrix}$
so that there are two linearly independent syzygies, one of degree zero and one of degree four, in the following matrix:
$\nu=\begin{pmatrix}0&-x_{0}x_{1}^{3}\\ 0&x_{0}^{3}x_{1}\\ 1&0\\ 0&-x_{0}^{3}x_{3}+x_{1}^{3}x_{3}\end{pmatrix}$
Here, $m(\sigma)=12$, the annihilator ideal and the 0-th Fitting ideal coincide, with the support being three lines $V(x_{0},x_{3}),V(x_{1},x_{3})$ and $V(x_{0},x_{1})$, where the last one has multiplicity $10$ and the other two are simple.
Using Section 4, we obtain the following bounds for $\mu$-semistability of ${\mathcal{T}}_{\sigma}$ in terms of $m(\sigma)$:
Proposition 35.
Let $\sigma$ be a normal pencil of cubics. Then:
  • (a)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable;
  • (b)
    If $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$. Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 33, 36 and 38. Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. Since $\sigma$ is not free, we can apply the result Section 4, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$. On the other hand, we have
$\operatorname{Bour}(\sigma)$ $=1-d+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$
$=1-4+4+4+4-m(\sigma)$
$=9-m(\sigma)>2,$
since $m(\sigma)\leq 6$. This contradicts the bound $\operatorname{Bour}(\sigma)\leq 2$ established earlier.
Moreover, for $(b)$, if we assume $m(\sigma)\leq 2$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then
$\operatorname{Bour}(\sigma)$ $=2-2d+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$
$=4-8+12-m(\sigma)$
$=8-m(\sigma)>5$
if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Section 4, $(b)$. ∎
Proposition 36.
Let $\sigma=(f,g)$ be an incompressible normal pencil of cubics. Then $m(\sigma)\leq 9$, and $m(\sigma)=9$ if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$ (see Example 37).
Proof.
The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 29 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$. From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$. On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, then $e=1$, and the equation $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=9$. ∎
Example 37 (Free, unstable incompressible pencil of cubics ($m(\sigma)=9$)).
Consider the sequence $\sigma=(x_{1}(x_{2}^{2}-x_{1}^{2}),x_{3}x_{2}(x_{0}-x_{1}))$. Then the matrix $\nabla\sigma$ is given by:
$\nabla\sigma=\begin{pmatrix}0&-3x_{1}^{2}+x_{2}^{2}&2x_{1}x_{2}&0\\ x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\end{pmatrix}$
and it admits two linearly independent syzygies, one of degree one and one of degree $3$:
$\nu\doteq\begin{pmatrix}x_{0}-x_{1}&2x_{1}x_{2}^{2}\\ 0&2x_{1}x_{2}^{2}\\ 0&3x_{1}^{2}x_{2}-x_{2}^{3}\\ -x_{3}&-3x_{1}^{2}x_{3}+x_{2}^{2}x_{3}\end{pmatrix}$
Thus, we conclude ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, and in particular $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. In this case, $m(\sigma)=9$.
Proposition 38.
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ if and only if $m(\sigma)=8$ (see Example 39).
Proof.
If $m(\sigma)=8$, by 31, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 2$. If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [3, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$. If we assume $e=1$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 7 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain $c_{3}({\mathcal{T}}_{\sigma})=4$, on the other hand, 30 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction and this case cannot happen.
On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$. ∎
Example 39 (Free, incompressible and $\mu$-semistable pencil of cubics ($m(\sigma)=8$)).
Consider the sequence $\sigma=(x_{0}^{2}x_{1}+x_{3}^{3},x_{0}^{3}+x_{0}x_{2}x_{3}+x_{3}^{3})$. The Jacobian matrix $\nabla\sigma$ is given by:
$\nabla\sigma=\begin{pmatrix}2x_{0}x_{1}&x_{0}^{2}&0&3x_{3}^{2}\\ 3x_{0}^{2}+x_{2}x_{3}&0&x_{0}x_{3}&x_{0}x_{2}+3x_{2}^{2}\end{pmatrix}$
and it admits two linearly independent syzygies of degree $2$:
$\nu\doteq\begin{pmatrix}-x_{0}x_{3}&-x_{0}x_{2}\\ 2x_{1}x_{3}&2x_{1}x_{2}-9x_{3}^{2}\\ 3x_{0}^{2}+x_{2}x_{3}&x_{2}^{2}-9x_{0}x_{3}\\ 0&3x_{0}^{2}\end{pmatrix}$
Thus, we conclude ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$.
Proposition 40.
Let $\sigma=(f,g)$ be a nearly free pencil of cubics. Then, the only possible discrete invariants are $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $m(\sigma)=7$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 41).
Proof.
Using Remark 7 for $p_{a}(B)=0$ and $\deg(B)=1$, we obtain
$c_{3}({\mathcal{T}}_{\sigma})$ $=2p_{a}(B)-2+\deg(B)(4+d-2e)$
$=-2+8-2e=2(3-e).$
On the other hand, if $m(\sigma)=7$, then $e=2$ and the formula for $c_{3}$ in Remark 7 yields $c_{3}({\mathcal{T}}_{\sigma})=2$. As we have observed in Example 39 and before, we must have $m(\sigma)\leq 7$ to be able to obtain $\operatorname{Bour}(\sigma)=1$. On the other hand, assuming $m(\sigma)\leq 6$, we obtain
$\operatorname{Bour}(\sigma)=12-m(\sigma)+e(e-4)\geq 6+e(e-4)\geq 2$
hence $\operatorname{Bour}(\sigma)\neq 1$. ∎
Example 41 (Nearly free pencil of Cubics).
We consider the following sequence of cubics:
$\sigma=(x_{0}^{2}(x_{1}-x_{2})+x_{2}^{2}(x_{1}-x_{0}+x_{3}),-x_{1}x_{2}x_{3}+x_{2}^{2}x_{3})$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&x_{0}^{2}+x_{2}^{2}&-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})&x_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$
Using Macaulay2 software, we compute $e=2$ and $m(\sigma)=7$, so that $\operatorname{Bour}(\sigma)=1$.
Proposition 42.
Let $\sigma=(f,g)$ be a normal pencil of cubics such that $m(\sigma)=7$. Then, the only possible $\mu$-semistable cases are:
  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 40);
  • $e=3$ and $\sigma$ is locally free, $\operatorname{Bour}(\sigma)=2$ and $B$ is a pair of skew lines (or their degeneration) (see Example 24).
Proof.
For $e\geq 2$, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [3, Lemma 2.1], we conclude that the only possible cases are $c_{3}=0$ (stable case) or $c_{3}=2$ (strictly semistable case). Thus, for $e=2$, we must have $c_{3}({\mathcal{T}}_{\sigma})=2$, and thus we obtain the second case. For the third case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [22] fitting in a sequence of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}(2)\rightarrow\mathcal{I}_{B}(1)\rightarrow 0$
where $B$ is a pair of skew lines or their degeneration, thus $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=3$. ∎
As we mentioned in Remark 27, the only possible case with $m(\sigma)=7$ for which we do not have examples is when $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=8$. We summarize the results of this subsection in the following theorem:
Theorem C. Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 12$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=12,9$ or $8$, corresponding to $e$ being $0,1$ or $2$, respectively;
  • (c)
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 41), which is strictly $\mu$-semistable;
  • (d)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Item $(a)$ is 33 and item $(b)$ with 36 and Example 39. Item $(c)$ follows from 42 and item $(d)$ is 35. ∎
To finish this study, we consider a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
Example 43 ($m(\sigma)=4$, $e=2$, $\operatorname{Bour}(\sigma)=4$, $3$-syzygy).
Considering the following pencil of cubics $(d_{f}=d_{g}=2)$, where the first one is smooth:
$\sigma=(x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3},x_{0}^{3}+x_{1}^{3}+x_{2}x_{3}^{2}),$
with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}&3x_{1}^{2}&3x_{2}^{2}&3x_{3}^{2}\\ 3x_{0}^{2}&3x_{1}^{2}&x_{3}^{2}&2x_{2}x_{3}\end{pmatrix}$.
The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$ with multiplicity $4$. The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
where the first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2,\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=4$. This is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.
5.2 Degree $6$ curves inside quadric surfaces
Let us focus on the case of normal sequences $\sigma=(f,g)$ with $d_{f}=1,d_{g}=2$. Here, the Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ is given by:
$\operatorname{Bour}(\sigma)=e(e-3)+7-m(\sigma).$
Assuming $\sigma$ is incompressible, as $c_{1}({\mathcal{T}}_{\sigma})=-3$, we can apply [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}$ without further restrictions, and obtain the following upper bounds for $m(\sigma)$ and $c_{3}({\mathcal{T}}_{\sigma})$:
Proposition 44.
If $\sigma=(f,g)$ is an incompressible normal sequence with $d_{f}=1,d_{g}=2$, then $m(\sigma)\leq 5$.
Proposition 45.
Let $\sigma=(f,g)$ be an incompressible normal sequence with $d_{f}=1$ and $d_{g}=2$. Then, the following hold:
  • (a)
    If $4\leq m(\sigma)\leq 5$, then $c_{3}\leq 5-m(\sigma))$. In particular, when $m(\sigma)=5$, $\sigma$ is locally free;
  • (b)
    If $0\leq m(\sigma)<4$, then $c_{3}\leq m(\sigma)^{2}-10m(\sigma)+25.$
From Section 3.1, $(d)$, we obtain:
Proposition 46.
Let $\sigma$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $\sigma$ is compressible if and only if $m(\sigma)=7$ (see Example 47).
Example 47 (Free and compressible sequence with $d_{f}=1,d_{g}=2$).
Considering the sequence
$\sigma=(x_{0}(x_{1}-x_{2}),x_{0}^{3}+x_{1}^{3}+x_{2}^{3}),$
which is independent of the variable $x_{3}$. The matrix
$\begin{pmatrix}0&-x_{0}(x_{1}^{2}+x_{2}^{2})\\ 0&x_{0}^{3}+x_{1}x_{2}^{2}-x_{2}^{3}\\ 0&x_{0}^{3}-x_{1}^{3}+x_{1}^{2}x_{2}\\ 1&0\end{pmatrix}$
gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
The following result is obtained using ([15], $0.1$):
Proposition 48.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$, and we have the following table of bounds for the following possibilities of $m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
4 $e\leq 2$
5 $e\leq 1$
As an easy consequence of Remark 7, we obtain:
Proposition 49.
Let $\sigma=(f,g)$ be a normal sequence which is not a pencil, (i.e., with $d_{f}\neq d_{g}$) such that $d$ is odd. If $\operatorname{Bour}(\sigma)$ is odd, then $\sigma$ is not locally free.
Proof.
Assuming ${\mathcal{T}}_{\sigma}$ is locally free, we get $c_{3}({\mathcal{T}}_{\sigma})=0$, and therefore
$2p_{a}(B)$ $=2-\operatorname{Bour}(\sigma)(4+d-2e)$
$p_{a}(B)$ $=1-\operatorname{Bour}(\sigma)(2+\frac{d}{2}-e)=1-(2-e)\operatorname{Bour}(\sigma)+\frac{d\operatorname{Bour}(\sigma)}{2},$
which is not an integer if $\operatorname{Bour}(\sigma)$ is odd, so we get a contradiction. ∎
Proposition 50.
Let $\sigma$ be an incompressible sequence with $d_{f}=1,d_{g}=2$ such that $m(\sigma)$ is even, that is, $m(\sigma)\in\{0,2,4\}$. Then $\sigma$ is not locally free.
Proof.
To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 49.
From 4, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
Assuming $e=2k$ is even, we obtain
$e^{2}-3e=4k^{2}-6k=2(2k^{2}-3k),$
an even number. On the other hand, when $e=2k+1$ is odd, then
$e^{2}-3e=4k^{2}+4k+1-6k-3=4k^{2}-2k-2=2(2k^{2}-k-1),$
which is also even. ∎
Proposition 51.
Let $\sigma=(f,g)$ be an incompressible normal sequence with degrees $d_{f}=1$, $d_{g}=2$. Then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ if and only if $m(\sigma)=5$ (see Example 52).
Proof.
Assuming $\sigma$ is non-compressible, by 48, we obtain that $e=1$. But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free. On the other hand, if we assume ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=1$ and $\operatorname{Bour}(\sigma)=0$ give $m(\sigma)=5$. ∎
Example 52 (Free and incompressible sequence, $m(\sigma)=5$).
Considering the sequence
$\sigma=(x_{0}x_{1},x_{3}x_{2}(x_{0}-x_{1})),$
of arrangements of hyperplanes, with Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&0&0\\ x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\end{pmatrix}$
The matrix
$\begin{pmatrix}0&x_{0}(x_{0}-x_{1})\\ 0&-x_{1}(x_{0}-x_{1})\\ x_{2}&0\\ -x_{3}&-x_{3}(x_{0}+x_{1})\end{pmatrix}$
gives linearly-independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ and $m(\sigma)=5$.
Proposition 53.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=4$, then $\sigma$ is nearly free, and we have two possible cases:
  • (a)
    ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=1$ (see Example 54);
  • (b)
    ${\mathcal{T}}_{\sigma}$ is unstable with $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 15);
Furthermore, these are the only two possibilities of numerical invariants for nearly free sequences with $d_{f}=1,d_{g}=2$.
Proof.
Using 48, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$. Using Remark 7 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples. To conclude the last claim, we note that if $m(\sigma)\leq 3$, then
$\operatorname{Bour}(\sigma)=7-m(\sigma)+e(e-3)\geq 4+e(e-3)\geq 3$
for $e\in\{1,2,3\}$. ∎
Example 54 (Nearly free sequence with $d_{f}=1,d_{g}=2$ and $e=2$).
We consider the following normal sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(x_{0}x_{1}-x_{2}x_{3},x_{1}x_{3}(x_{0}-x_{2})),$
with Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&-x_{3}&-x_{2}\\ x_{1}x_{3}&x_{3}(x_{0}-x_{2})&-x_{1}x_{3}&x_{1}(x_{0}-x_{2})\end{pmatrix}$.
In this case, we have equality of schemes $\Xi_{\sigma}=\operatorname{supp}(\mathcal{Q}_{\sigma})$. The Jacobian scheme has a structure of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$, which is outside the three lines, and therefore $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$. The free resolution of ${\mathcal{T}}_{\sigma}$ is (obtained computationally):
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 3}\xrightarrow{M}{\mathcal{T}}_{\sigma}\rightarrow 0$
given by matrices
$M=\begin{pmatrix}x_{0}x_{1}+x_{3}(x_{2}-x_{0})&x_{0}^{2}&x_{0}x_{3}\\ -x_{1}^{2}+x_{1}x_{3}&-x_{0}x_{1}&0\\ x_{1}x_{2}&x_{2}^{2}&x_{0}x_{1}+x_{3}(x_{2}-x_{1})\\ 0&-x_{2}x_{3}&x_{1}x_{3}-x_{3}^{2}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}\\ x_{3}-x_{1}\\ -x_{2}\end{pmatrix}$.
So we obtain $m(\sigma)=4,e=2$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=1$ corresponding to the point $p$, which is an irreducible component of $\Xi_{\sigma}$ of codimension three.
Proposition 55.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $\operatorname{Bour}(\sigma)=2$ and we may have $c_{3}=0,2,4$. We have examples for the cases $c_{3}=2$ (see Example 56) and $c_{3}=4$ (see Example 57).
Proof.
First, the bound $c_{3}\leq 4$ is obtained from 45 for $m=m(\sigma)=3$. As we shown in 28, $e\geq 2$. We claim that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ in this case.
We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$. Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$. This follows from [21, Proposition 1.1] for $c_{3}=0$, [3, Lemma 2.4] for $c_{3}=2$ and [14, Lemma 9.6] for $c_{3}=4$. ∎
Example 56 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=2$).
We consider the following sequence:
$\sigma=(x_{3}(x_{0}-x_{1}),x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{3}^{3})$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$.
Here, the saturation of the annihilator ideal coincides with the saturation of the 0-th Fitting ideal, and the first Fitting ideal has codimension three. The support of $\mathcal{Q}_{\sigma}$ in codimension two consists of two lines $V(x_{3},x_{0})\cup V(x_{3},x_{0}-x_{1})$, and the first one has multiplicity two. Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
so we get $\operatorname{Bour}(\sigma)=2,m(\sigma)=3$ and $c_{3}({\mathcal{T}}_{\sigma})=2$.
Example 57 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=4$).
We consider the following sequence:
$\sigma=(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2},x_{3}(x_{2}-x_{3})(x_{0}-x_{1}))$
with corresponding Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}2x_{0}&2x_{1}&2x_{2}&2x_{3}\\ x_{2}x_{3}-x_{3}^{2}&x_{3}^{2}-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})+2x_{3}(x_{1}-x_{0})\end{pmatrix}$.
Here, the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $R$. The codimension two locus of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of three simple lines. Thus, $m(\sigma)=3$. Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $\operatorname{Bour}(\sigma)=2$ and $c_{3}({\mathcal{T}}_{\sigma})=4$.
Corollary 58.
Let $\sigma$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since $\mu({\mathcal{T}}_{\sigma})=-3/2$, then ${\mathcal{T}}_{\sigma}$ is stable if and only if $e>1$. From the result above, the only possibilities of $\operatorname{Bour}(\sigma)\leq 2$ are when $m(\sigma)\geq 3$, and moreover if $m(\sigma)=3$ we cannot have $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. For $m(\sigma)<3$, $\operatorname{Bour}(\sigma)>2$ and by Section 4, $(a)$, $e\neq 1$. ∎
Proposition 59.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)=2$, then $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, $c_{3}\in\{1,3,5,7\}$ and $c_{3}=7$ if and only if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, with $\operatorname{Bour}(\sigma)=3$ and $p_{a}(B)=0$ (see Example 60).
Proof.
If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [14, Section 9]. By the proof of [14, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
From Remark 7, we obtain $p_{a}(B)=1$, and thus $B$ must be a plane cubic curve, with a resolution of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
which by Lemma 13 yields a resolution
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [14, Lemma $9.3$]. If $c_{3}=7$, then ${\mathcal{T}}_{\sigma}(1)$ is stable with Chern classes (1,3,7)(-1,3,7). From [3, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$. ∎
Example 60 ($m(\sigma)=2,e=2$, $\operatorname{Bour}(\sigma)=3$ , $c_{3}=7$).
Consider the following sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(-x_{0}x_{1}+x_{1}x_{2}-x_{2}x_{3},x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3}),$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}-x_{1}&x_{2}-x_{0}&x_{1}-x_{3}&-x_{2}\\ x_{1}^{2}&2x_{0}x_{1}&3x_{2}^{2}+2x_{2}x_{3}&x_{2}^{2}\end{pmatrix}$.
Here the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the codimension two part of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of a line $V(x_{1},x_{2})$ with multiplicity two. The saturation of the first Fitting ideal is $(0)$ in this case. The sheaf ${\mathcal{T}}_{\sigma}$ admits the following free resolution:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $e=2$, $\operatorname{Bour}(\sigma)=3$ and $m(\sigma)=2$, with $c_{3}({\mathcal{T}}_{\sigma})=7$.
Example 61 ($m(\sigma)=2$, $e=3$, $\operatorname{Bour}(\sigma)=5$, $c_{3}=3$).
We consider the following sequence:
$\sigma=(x_{2}x_{3}-x_{0}x_{1},x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{2}x_{3}^{2}+x_{3}^{3})$
with corresponding Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}-x_{1}&-x_{0}&x_{3}&x_{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}+x_{3}^{2}&x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2}\end{pmatrix}$.
Here, the saturation of the 0-th Fitting ideal coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $(1)$. Moreover, $m(\sigma)=2$ with $\operatorname{supp}(\mathcal{Q}_{\sigma})$ being a double line structure at $V(x_{0},x_{3})$. Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 3}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so we obtain $\operatorname{Bour}(\sigma)=5$ and $c_{3}({\mathcal{T}}_{\sigma})=3$.
We summarize the results of this subsection in the following theorem:
Theorem D. Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 7$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=7$ or $5$, and each corresponds to $e$ being 0 or $1$, respectively;
  • (c)
    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 54 and Example 15);
  • (d)
    If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Item $(a)$ is 46 and item $(b)$ follows with 51. Item $(c)$ is shown in 53, and the stability result in $(d)$ is in 58. ∎