### v3

# Bourbaki degree of pairs of projective surfaces

Abstract
The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials — defined as the kernel of the associated Jacobian matrix — naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation. As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
The present work focuses on sequences of two homogeneous polynomials in four variables. We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis. The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation of the degrees. We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants. Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly free cases. In particular, one of the nearly free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
1 Introduction
Let $R=\kappa[x_{0},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$ variables with coefficients in an algebraically closed field $\kappa$. For an algebraically independent sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials in $R$ with degrees $d_{1}+1,\ldots,d_{k}+1$, respectively, with $d_{1}\leq\ldots\leq d_{k}$ one can consider the Jacobian matrix as a map of locally free sheaves on $\mathbb{P}^{n}\doteq\operatorname{Proj}R$
$\nabla\sigma=\begin{pmatrix}\nabla f_{1}\\ \nabla f_{2}\\ \vdots\\ \nabla f_{k}\end{pmatrix}:\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\bigoplus_{i=1}^{k}\mathcal{O}_{\mathbb{P}^{n}}(d_{i}).$
The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$. This definition is in analogy (see [undefj]) with the case of the tangent logarithmic sheaf associated to divisors in $\mathbb{P}^{n}$, namely for $k=1$. A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
In the case of divisors on $\mathbb{P}^{n}$, where $k=1$, one has a short exact sequence of the form:
$0\rightarrow{\mathcal{T}}_{f}\rightarrow\mathcal{O}_{\mathbb{P}^{n}}^{\oplus(n+1)}\rightarrow\mathcal{I}_{J_{f}}(d)\rightarrow 0,$
where $J_{f}=\langle\partial_{0}f,\ldots,\partial_{n}f\rangle\trianglelefteq R$ is the Jacobian ideal of the homogeneous polynomial $f$. The sheaf ${\mathcal{T}}_{f}$ is the sheaf of $\mathcal{O}_{\mathbb{P}^{n}}$-modules associated to the graded $R$-module $\operatorname{Syz}(J_{f})$ of Jacobian syzygies of $f$. For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated to the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
For the generalization of $k>1$, one should consider sequences rather than the associated ideals $I=(f_{1},\ldots,f_{k})$, as pointed out by [undefj, Example 2.7]. When $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and furthermore ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [undefj, Lemma 2.14]).
We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$, called the initial degree. Following the terminology of [undefj], we call a sequence $\sigma=(f_{1},\ldots,f_{k})$ compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$. This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$. Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [undefj, Lemma 2.8]).
One may also consider the relationship with Bourbaki ideals, as in [undefr] and [undeff]. For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{n}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B_{\nu}}(e-d)\rightarrow 0$
where $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection. The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice. In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [undefr]. By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. The focus of this work is to understand the concept above for $n=3$ and $k=2$.
For $k=1$, the singular scheme of $V(f)$, defined by the Jacobian ideal $\Sigma_{f}\doteq V(J_{f})\subset\mathbb{P}^{n}$, plays an important role in this study. If $s=\dim\Sigma_{f}$, the Hilbert polynomial is given by
$H(\mathcal{O}_{J_{f}}(d),t)=H(R/J_{f},t)=\frac{\deg(\Sigma)}{s!}t^{s}+O(t^{s-1}),$
and, by definition, the degree $\deg(\Sigma)$ is the leading coefficient above. Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$. For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [undefr]):
$\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}),$
where $e=\operatorname{indeg}({\mathcal{T}}_{f})$.
Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for the logarithmic sheaf ${\mathcal{T}}_{f}$ to get an ideal sheaf as the cokernel. This problem is addressed in a recent work [undeff] where the authors consider another kind of Bourbaki sequences, recently developed in [undefq], by choosing generators of maximal degree.
We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel of the Jacobian matrix, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$. In [undefj], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$. In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$. The generic case is when $\dim(\mathcal{Q}_{\sigma})\leq n-2$, and in particular $c_{1}(\mathcal{Q}_{\sigma})=0$. In this case, we call $\sigma$ a normal sequence.
For now on, we assume $n=3$ and $k=2$, that is, sequences $\sigma=(f,g)$ of homogeneous polynomials $f,g\in R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, and we set $d\doteq d_{f}+d_{g}$. We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and $c_{1}({\mathcal{T}}_{\sigma})=-d$. Since the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
If $\sigma$ is a normal sequence, $\dim(\mathcal{Q}_{\sigma})\leq 1$ and the Hilbert polynomial is of the form
$H(\mathcal{Q}_{\sigma},t)=\deg(\mathcal{Q}_{\sigma})t+O(1),$
We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ is zero-dimensional. From the associativity formula applied to the graded homogeneous module $\mathcal{Q}$ associated to $\mathcal{Q}_{\sigma}$, we obtain the expression (see, for example, [undefs, Theorem $14.7$])
$m(\sigma)=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(R/P)=2}\operatorname{{length}}_{R_{P}}\mathcal{Q}_{P}\cdot\deg(R/P).$
Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we reproduce analogous considerations as in [undefr]. We show a formula for the Bourbaki degree of $\sigma$ given by
$\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma),$
so that $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. A first result of the paper is some bounds for the discrete quantities defined above.
Theorem A(Theorem A) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. Then:
  • (a)
    $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d$;
  • (b)
    $m(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$;
  • (c)
    The following are equivalent:
    • (1)
      $m(\sigma)=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$;
    • (2)
      $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$;
    • (3)
      ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$;
    • (4)
      $\sigma$ is compressible;
We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 2.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 8): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable. In general, the Bourbaki degree is bounded by $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, and the maximum is attained precisely by the matrices $\nabla\sigma$ for which the Buchsbaum–Rim complex is a minimal resolution, or equivalently $m(\sigma)=0$. Geometrically, all regular pencils ($d_{f}=d_{g}$) with isolated singularities satisfy this property (see Remark 9).
In Section 2.2, we relate our notion to the notion of the Bourbaki degree for a projective plane curve, by considering a reduced polynomial $g\in\kappa[x_{0},x_{1},x_{2}]$ together with its associated surface $V(g)\subset\mathbb{P}^{3}$ and projective curve $X=V(g)\subset\mathbb{P}^{2}$, to obtain $\operatorname{Bour}(X)=\operatorname{Bour}(g,x_{3})$.
Afterwards, in Section 2.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is the Bourbaki scheme obtained by the choice of a minimal syzygy. Related to the free resolutions, we introduce the notions of nearly free sequences and $3$-syzygy sequences, with the following chain of implications
$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3\text{-syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1.$
The converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$. We point out nearly free sequences are never locally free, as we show in 19.
In Section 3, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [undefj, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections. For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [undefr, Corollary 2.11]). The main theorem of the section is:
Theorem B.(Theorem B) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials with degrees $d_{f}+1,d_{g}+1$, respectively. Then:
  • (a)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, then $\operatorname{Bour}(\sigma)\in\{0,1,2\}$;
  • (b)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then $\operatorname{Bour}(\sigma)\leq 5$.
In Section 4, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$. To do so, we use classical results on bounds for Chern classes of reflexive sheaves in $\mathbb{P}^{3}$. We classify all free and nearly free cases in these classes in terms of their discrete invariants and establish some stability results, which are summarized below:
Theorem C.(Theorem C)Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 12$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=12,9$ or $8$, corresponding to $e$ being $0,1$ or $2$, respectively;
  • (c)
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42), which is strictly $\mu$-semistable;
  • (d)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 15), but the bound given above for stability may not be sharp.
Theorem D.(Theorem D) Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 7$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=7$ or $5$, and each corresponds to $e$ being 0 or $1$, respectively;
  • (c)
    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 56 and Example 14);
  • (d)
    If $m(\sigma)=3$ then $\operatorname{Bour}(\sigma)=2$, with a possible unstable case $(c_{3}=6)$ and stable cases with $c_{3}\in\{0,2,4\}$. Among these, we have examples for the stable cases with $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
  • (e)
    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Example 14 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable. This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [undefa]:
Conjecture. If the tangent sheaf of a codimension one foliation on $\mathbb{P}^{3}$ is not split, then it is $\mu$-semistable.
Along the text, we describe examples developed computationally with aid of Macaulay2 software ([undefl]) for the two families of sequences.
Acknowledgments
FM is financed by the São Paulo Research Foundation (FAPESP), Brasil. Process Number #2021/10550-4. FM is currently a PhD. student under the advice of Marcos Jardim and Daniele Faenzi, with partial funding by the Bridges ”Brazil-France interplays in Gauge Theory, extremal structures and stability” projects ANR-21-CE40-0017 and ANR-17-EURE-0002. We thank Marcos Jardim, Daniele Faenzi, Alan Muniz, Victor Cordeiro and Abbas Nejad for fruitful discussions and suggestions regarding this work.
2 The Bourbaki degree of pairs of projective surfaces
In this section, we develop the concept of the Bourbaki degree of pairs of projective surfaces on $\mathbb{P}^{3}$, determined by normal sequences $\sigma=(f,g)$ of homogeneous polynomials. We start with definitions and first results (Section 2.1), followed by a reduction to the case of a projective plane curve (Section 2.2) and finish with comparison results relating the geometry of the Bourbaki scheme and the associated logarithmic sheaf, using free resolutions (Section 2.3), in particular introducing the class of nearly free sequences $\sigma$, which are characterized by $\operatorname{Bour}(\sigma)=1$.
2.1 Framework and first results
By a sequence $\sigma=(f,g)$, unless otherwise stated, we mean an algebraically independent sequence of two homogeneous polynomials in $R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$. By curve we mean a locally Cohen-Macaulay closed subscheme of $\mathbb{P}^{3}$ of pure dimension one.
Each sequence $\sigma=(f,g)$ induces a morphism of sheaves on $\mathbb{P}^{3}$ by the Jacobian matrix:
$\nabla\sigma:\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d_{f})\oplus\mathcal{O}_{\mathbb{P}^{3}}(d_{g}),$
and we denote the kernel by ${\mathcal{T}}_{\sigma}$, the image by $\mathcal{M}_{\sigma}$ and the cokernel by $\mathcal{Q}_{\sigma}$. We say a sequence $\sigma$ is free whenever ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles, following [undefj].
The following lemma relates the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ and its Chern characters.
Lemma 1.
Let $\mathcal{Q}$ be a coherent sheaf on $\mathbb{P}^{3}$ with $\operatorname{{rk}}(\mathcal{Q})=0$ and $c_{1}(\mathcal{Q})=0$. Then the Hilbert polynomial of $\mathcal{Q}$ is given by
$\mathcal{X}(\mathcal{Q}(t))=\operatorname{ch}_{2}(\mathcal{Q})t+\operatorname{ch}_{3}(\mathcal{Q})+2\operatorname{ch}_{2}(\mathcal{Q}).$
Proof.
Follows from direct application of the Hirzebruch-Riemann-Roch theorem. ∎
We denote by $m(\sigma)\doteq\operatorname{ch}_{2}(\mathcal{Q}_{\sigma})$. From the previous formula, $m(\sigma)$ is non-negative, and it is zero if and only if the Hilbert polynomial of $\mathcal{Q}_{\sigma}$ is constant, that is, if and only if $\mathcal{Q}_{\sigma}$ is a zero-dimensional sheaf.
As in the introduction, the Jacobian scheme of $\nabla\sigma$ is defined as
$\Xi_{\sigma}\doteq V\left(\bigwedge^{2}\nabla\sigma\right)=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
the zero-locus of the $2\times 2$ minors of $\nabla\sigma$ or, in other words, the $0$-th Fitting scheme of $\nabla\sigma$. From the general theory of Fitting ideals (see, for example, [undefh, Chapter 20]), the annihilator ideal of $\mathcal{Q}_{\sigma}$ contains the Fitting ideal, and they have the same support, namely
$\operatorname{supp}(\mathcal{Q}_{\sigma})_{\text{red}}=V(\operatorname{Fitt}_{0}(\nabla\sigma))$
as closed subsets, and therefore the irreducible components of the schematic support $\operatorname{supp}(\mathcal{Q}_{\sigma})$ coincide with the irreducible components of $\Xi_{\sigma}$. In general, however, their schematic structure may be different, as we explore in the following example. Denote by $(\Xi_{\sigma})_{1}$ the one-dimensional part of the scheme $\Xi_{\sigma}$.
Example 2.
Let $\sigma=(2x_{1}x_{3}-x_{1}^{2},3x_{2}x_{3}^{2}-3x_{0}x_{1}x_{3}+x_{1}^{3})$. The Jacobian matrix is of the form
$\begin{pmatrix}0&-2x_{1}+2x_{3}&0&2x_{1}\\ -3x_{1}x_{3}&3x_{1}^{2}-3x_{0}x_{3}&3x_{3}^{2}&-3x_{0}x_{1}+6x_{2}x_{3}\\ \end{pmatrix}$
and the matrix below
$\begin{pmatrix}x_{3}&x_{0}x_{1}-x_{1}^{2}-2x_{2}x_{3}\\ 0&-x_{1}x_{3}\\ x_{1}&-2x_{2}x_{3}\\ 0&-x_{1}x_{3}+x_{3}^{2}\end{pmatrix}$
gives two linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$.
In this case, the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$ and the 0-th Fitting ideal are different, given by:
$\begin{cases}\operatorname{Ann}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3},x_{0}x_{1}^{2}-x_{1}^{3})=(x_{1},x_{3})^{2}\cap(x_{3},x_{0}-x_{1})\\ \operatorname{Fitt}_{0}(\mathcal{Q}_{\sigma})&=(x_{3}^{2},x_{1}x_{3}^{2},x_{1}^{2}x_{3},x_{0}x_{1}^{2}-x_{1}^{3}-2x_{1}x_{2}x_{3}+2x_{2}x_{3}^{2}),\end{cases}$
so both schemes $\Xi_{\sigma}$ and $\operatorname{supp}(\mathcal{Q}_{\sigma})$ are non-reduced, with degrees $4$ and $6$, respectively. From the formula for $\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$, since ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ we conclude $m(\sigma)=5$. The right approach to computing $m(\sigma)$ in general is using the associativity formula, as we will in the next examples.
Any saturated syzygy of the Jacobian matrix $\nu$, of degree $e\in\mathbb{Z}$, induces a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\rightarrow{\mathcal{I}}_{B_{\nu}}(p)\rightarrow 0,$
where $B_{\nu}\subset\mathbb{P}^{3}$ is the curve associated by $\nu$, described as the Serre correspondence (see [undefn, Theorem $4.1$]) between the sheaf ${\mathcal{T}}_{\sigma}$ and the scheme $B$. Moreover, we have $\deg(B_{\nu})=c_{2}({\mathcal{T}}(e))$. Since ${\mathcal{T}}_{\sigma}\hookrightarrow{\mathcal{O}}_{\mathbb{P}^{3}}^{\oplus 4}$ and the latter is a $\mu$-semistable sheaf, $e\geq 0$. The following proposition describes a formula for the degree $\deg(B_{\nu})$ in terms of the discrete invariants $e$, $d_{f}$, $d_{g}$ and $m(\sigma)$.
Proposition 3.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1$, $\deg(g)=d_{g}+1$.
For any saturated syzygy $\nu\in H^{0}({\mathcal{T}}(e))$ of degree $e\geq 0$, let $B_{\nu}\subset\mathbb{P}^{3}$ be the associated closed subscheme of pure dimension one. Then, we have the following equation:
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+m_{0}-m(\sigma),$
where $m_{0}\doteq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$.
Proof.
From the hypothesis $c_{1}({\mathcal{Q}}_{\sigma})=0$, we conclude $c_{1}({\mathcal{T}}_{\sigma})=-(d_{f}+d_{g})$ and $\operatorname{ch}_{2}({\mathcal{Q}}_{\sigma})=m(\sigma)$. Moreover, from additivity of $\operatorname{ch}_{2}$ on short exact sequences:
$\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})$ $=-\operatorname{ch}_{2}({\mathcal{O}}_{\mathbb{P}^{3}}(d_{f})\oplus{\mathcal{O}}_{\mathbb{P}^{3}}(d_{g}))+\operatorname{ch}_{2}({\mathcal{Q}}_{\sigma})$
$=-\frac{d_{f}^{2}+d_{g}^{2}}{2}+m(\sigma).$
We can relate the Chern character and the Chern classes by the formula
$m(\sigma)$ $=\operatorname{ch}_{2}({\mathcal{T}}_{\sigma})+\frac{d_{f}^{2}+d_{g}^{2}}{2}=\frac{c_{1}^{2}({\mathcal{T}}_{\sigma})-2c_{2}({\mathcal{T}}_{\sigma})}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}$
$=\frac{(d_{f}+d_{g})^{2}}{2}+\frac{d_{f}^{2}+d_{g}^{2}}{2}-c_{2}({\mathcal{T}}_{\sigma})$
$=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-c_{2}({\mathcal{T}}_{\sigma}),$
so that
$c_{2}({\mathcal{T}}_{\sigma})=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma).$ (1)
On the other hand, since $c_{2}({\mathcal{T}}_{\sigma}(e))=\deg(B_{\nu})$ and ${\mathcal{T}}_{\sigma}$ is reflexive of rank $2$, we have
$\deg(B_{\nu})=c_{2}({\mathcal{T}}_{\sigma}(e))=c_{2}({\mathcal{T}}_{\sigma})+e\cdot c_{1}({\mathcal{T}}_{\sigma})+e^{2}.$
From this, together with 1, we obtain
$\deg(B_{\nu})=e^{2}-e(d_{f}+d_{g})+(d_{f}^{2}+d_{g}^{2}+d_{f}d_{g})-m(\sigma).$
We define the Bourbaki degree of a sequence $\sigma$, inspired by the construction in [undefr, Definition 2.4].
Definition 4.
Let $\sigma$ be a normal sequence and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. The Bourbaki degree of a normal sequence $\sigma$ is defined by:
$\operatorname{Bour}(\sigma)\doteq\deg(B_{\nu})=e(e-d)+m_{0}-m(\sigma),$
for some non-trivial syzygy $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. We note that every non-trivial syzygy of minimal degree is saturated.
Remark 5.
It follows from construction that the sequence $\sigma$ is free if and only if $\operatorname{Bour}(\sigma)=0$, since every short exact sequence of the form
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-e)\to{\mathcal{T}}_{\sigma}\to\mathcal{O}_{\mathbb{P}^{3}}(e-d)\to 0$
splits.
Remark 6.
From [undefm, Proposition 4.1], we have the formula
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e)$
relating the third Chern class of ${\mathcal{T}}_{\sigma}$ and the discrete invariants of $B$, when $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $B$ is the zero locus of a non-zero section in $H^{0}({\mathcal{T}}_{\sigma}(e))$. Moreover, dualizing the sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow\mathcal{I}_{B}(e-d)\rightarrow 0$
we conclude that the singular set of the sheaf ${\mathcal{T}}_{\sigma}$ is contained in $B$.
Now, we obtain some bounds for the quantities $\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ in terms of the degrees $d_{f},d_{g}$.
Theorem A.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $\kappa[x_{0},\ldots,x_{3}]$, with degrees $d_{f}+1,d_{g}+1$ respectively. Then:
  • (a)
    We have $\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq d_{f}+d_{g}$;
  • (b)
    $m(\sigma)\leq m_{0}$;
  • (c)
    The following are equivalent:
    • (1)
      $m(\sigma)=m_{0}$;
    • (2)
      $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$;
    • (3)
      $\sigma$ is compressible;
    • (4)
      ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$;
Proof.
To show $(a)$, we build explicit syzygies of the Jacobian matrix $\nabla\sigma$ of degrees $d_{f}+d_{g}$, and at least one of them is nonzero. Writing the Jacobian matrix by
$\nabla\sigma=\begin{pmatrix}\partial_{0}f&\partial_{1}f&\partial_{2}f&\partial_{3}f\\ \partial_{0}g&\partial_{1}g&\partial_{2}g&\partial_{3}g\\ \end{pmatrix}$,
the following vectors
$\nu_{0}$ $=\begin{pmatrix}0\\ \partial_{2}f\partial_{3}g-\partial_{3}f\partial_{2}g\\ -\partial_{1}f\partial_{3}g+\partial_{3}f\partial_{1}g\\ \partial_{1}f\partial_{2}g-\partial_{2}f\partial_{1}g\end{pmatrix};\nu_{1}=\begin{pmatrix}\partial_{2}f\partial_{3}g-\partial_{3}f\partial_{2}g\\ 0\\ -\partial_{0}f\partial_{3}g+\partial_{3}f\partial_{0}g\\ \partial_{0}f\partial_{2}g-\partial_{2}f\partial_{0}g\end{pmatrix}$;
$\nu_{2}$ $=\begin{pmatrix}\partial_{1}f\partial_{3}g-\partial_{3}f\partial_{1}g\\ -\partial_{0}f\partial_{3}g+\partial_{3}f\partial_{0}g\\ 0\\ \partial_{0}f\partial_{1}g-\partial_{1}f\partial_{0}g\end{pmatrix};\nu_{3}=\begin{pmatrix}\partial_{1}f\partial_{2}g-\partial_{2}f\partial_{1}g\\ -\partial_{0}f\partial_{2}g+\partial_{2}f\partial_{0}g\\ \partial_{0}f\partial_{1}g-\partial_{1}f\partial_{0}g\\ 0\end{pmatrix}$.
are a syzygies of degree $d_{f}+d_{g}$. Since there is at least one nonzero $2\times 2$ minor, from the hypothesis of algebraically independent, at least one of the syzygies $\nu_{0},\ldots,\nu_{3}$ is nonzero.
For $(b)$, using that $e\leq d$, we can use that
$0\leq\operatorname{Bour}(\sigma)=e(e-d)+m_{0}-m(\sigma),$
so that $m(\sigma)\leq m_{0}+e(e-d)$, but $e(e-d)\leq 0$ and the claim follows.
For $(c)$, we start by pointing out that the equivalence $(2)\iff(3)$ is the content of [undefj, Lemma 2.7]. Moreover, $(3)\Rightarrow(4)$ using that $c_{1}({\mathcal{T}}_{\sigma})=-d$ and [undefj, Lemma 2.8]. The implication $(4)\Rightarrow(1)$ can be obtained using the Bourbaki degree formula, since $\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$ and $\operatorname{Bour}(\sigma)=0$, as $\sigma$ is free.
To show $(1)\Rightarrow(2)$, let $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. From the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=e(e-d)$. Since $\operatorname{Bour}(\sigma)\geq 0$ and from part (a) we have $e\leq d$, it follows that $e=0$ or $e=d$. Both cases mean $\operatorname{Bour}(\sigma)=0$ and the sequence is free. Since $c_{1}({\mathcal{T}}_{\sigma})=-d$ is additive, it follows that $e=0$, otherwise we would have $c_{1}({\mathcal{T}}_{\sigma})<-d$. ∎
Remark 7.
The bound obtained in Theorem A, $(b)$ for $m(\sigma)$ relates to a known bound in foliation theory. We will use that ${\mathcal{T}}_{\sigma}(1)$ is the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$ of degree $d=d_{f}+d_{g}$ (see [undefj, Section 9]). Let $C$ be the one-dimensional part of the singular scheme of this foliation. From the formulas of discrete invariants in [undefa, Theorem 3.1] given for generic foliations,
$c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C).$
Using the formula $c_{2}({\mathcal{T}}_{\sigma})=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$ and the equations
$c_{2}({\mathcal{T}}_{\sigma})-d+1=c_{2}({\mathcal{T}}_{\sigma}(1))=d^{2}+2-\deg(C)$
we obtain $m(\sigma)=\deg(C)-d-d_{f}d_{g}-1$, so from the bound above we get
$\deg(C)-d-d_{f}d_{g}-1=m(\sigma)$ $\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g},$
and therefore $\deg(C)\leq d^{2}+d+1$, a bound that can be found in more generality for foliations in [undefv, Corollary 4.8].
From simple observations about the formula $\operatorname{Bour}(\sigma)$, we are able to obtain the following inequalities related to $\mu$-stability and freeness of the logarithmic sheaves.
Proposition 8.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $\kappa[x_{0},\ldots,x_{3}]$. Denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $d=d_{f}+d_{g}$. Then:
  • (a)
    $\operatorname{Bour}(\sigma)\leq m_{0}$, and equality holds if any of the equivalent facts hold:
    • (1)
      $m(\sigma)=0$;
    • (2)
      A minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
      $0\to\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{1})\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d-d_{2})\xrightarrow{\varphi}\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}(-d)\xrightarrow{\psi}{\mathcal{T}}_{\sigma}\to 0,$
      a Buchsbaum–Rim resolution, where
      $\psi=\begin{pmatrix}0&\Delta_{12}&\Delta_{13}&\Delta_{14}\\ -\Delta_{12}&0&\Delta_{23}&\Delta_{24}\\ -\Delta_{13}&-\Delta_{23}&0&\Delta_{34}\\ -\Delta_{14}&-\Delta_{24}&-\Delta_{34}&0\end{pmatrix},\quad\phi=\begin{pmatrix}\partial_{0}f&-\partial_{0}g\\[2.0pt] -\partial_{1}f&\partial_{1}g\\[2.0pt] \partial_{2}f&-\partial_{2}g\\[2.0pt] -\partial_{3}f&\partial_{3}g\end{pmatrix}$,
      and each $\Delta_{ij}$ is the $(i,j)$-minor of the matrix $\nabla\sigma$. In particular, $e=d$.
  • (b)
    If $\operatorname{Bour}(\sigma)>(d_{f}-1)(d_{f}+d_{g})+d_{g}^{2}+1$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
  • (c)
    If
    $m(\sigma)<\frac{1}{2}\left(\frac{3d_{f}^{2}}{2}+\frac{3d_{g}^{2}}{2}+d_{f}d_{g}\right),$
    then $\sigma$ is not free.
Proof.
The claim $(a)$ follows from the formula $\operatorname{Bour}(\sigma)\geq 0$, since $m(\sigma)\geq 0$ and $e(e-d)\leq 0$, from the inequality $e\leq d$. This also shows equality occurs whenever e(ed)e(e-d) and $m(\sigma)$ are both zero, therefore $e=d$, since $e=0$ means $\operatorname{Bour}(\sigma)=0$ from compressibility. The converse also follows simply from the formula, assuming that $m(\sigma)=0$ and $e=d$.
To show the equivalences above, we start by noting that $m(\sigma)=0$ implies that the grade of the ideal of $(2\times 2)$-minors is three, and therefore the Buchsbaum–Rim complex is a minimal free resolution for $\nabla\sigma$. From this fact, we obtain that $e=d$. For the converse, since $e=d$ and the Buchsbaum–Rim complex is a minimal resolution, we obtain that the grade of the ideal of minors of $\nabla\sigma$ is at least three, hence $\dim(\mathcal{Q}_{\sigma})=0$ and thus $m(\sigma)=0$.
By construction, $m(\sigma)\geq 0$, and therefore for a given value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, the Bourbaki degree of $\sigma$ can be at most
$\operatorname{Bour}(\sigma)\leq e(e-d)+m_{0},$
and this is a function $H=H(e)$ which attains its minimum at $e=d/2$. The function $H$ is decreasing on $e\in\{1,\ldots,d/2\}$, and if $e>d/2$, ${\mathcal{T}}_{\sigma}$ is $\mu$-stable, since $\mu({\mathcal{T}}_{\sigma})=-d/2$. Thus, the maximum value $\operatorname{Bour}(\sigma)$ in the range $e\in\{1,\ldots,d/2\}$ for any possible value of $m(\sigma)$ is $H(1)$, which is the expression above on the right-hand side. If $\operatorname{Bour}(\sigma)$ is higher than this, then $e>d/2$, and thus we obtain $(b)$.
With the same strategy as $(b)$, since the function $H(e)$ attains its minimum at $e=d/2$, for $m(\sigma)$ satisfying the inequality of the claim we obtain that
$\operatorname{Bour}(\sigma)=H(e)\geq H(d/2)=\frac{d^{2}}{4}-\frac{d^{2}}{2}+m_{0}-m(\sigma)>0,$
and therefore $\operatorname{Bour}(\sigma)\neq 0$, independently of the value of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$. ∎
Remark 9.
Let $\sigma=(f,g)$ be a sequence of surfaces of the same degree, say $d_{f}=d_{g}=p$, called a pencil of surfaces. It defines a 11-parameter family of projective surfaces as:
$V_{\sigma}\doteq\{V(z_{0}f+z_{1}g)\subset\mathbb{P}^{3}:z=[z_{0}:z_{1}]\in\mathbb{P}^{1}\},$
and each hypersurface $V(z_{0}f+z_{1}g)$ is called a member of the pencil $V_{\sigma}$. The set where each surface is singular is the vanishing locus of a polynomial $p=p(z)$ (the discriminant) inside $\mathbb{P}^{1}$, so it must be either a finite subset or the whole family. When this set is finite, the pencil is called a regular pencil, otherwise is called irregular.
In [undefj, Lemma 2.17], the authors describe the support of the sheaf $\mathcal{Q}_{\sigma}$ for pencils of surfaces as the union of singular loci:
$(\Xi_{\sigma})_{\text{red}}=\bigcup_{[z_{0}:z_{1}]\in\mathbb{P}^{1}}\operatorname{Sing}(V(z_{0}f+z_{1}g)).$
Hence, whenever a normal sequence $\sigma=(f,g)$ is a regular pencil where every singular member has isolated singularities, we conclude $\dim(\Xi_{\sigma})=0$ from the description above, and thus $m(\sigma)=0$.
Example 10.
Consider the sequence
$\sigma=(x_{3}(x_{0}x_{2}-x_{1}^{2})-(x_{0}-2x_{1})(3x_{1}-x_{0}-2x_{2})(x_{1}-2x_{2}),x_{3}(x_{0}x_{2}-x_{1}^{2})-x_{1}^{2}(x_{0}-x_{1}))$
where $f$ is a normal singular cubic with an $A_{1}$-singularity at $[0:0:0:1]$ and $g$ is a normal singular cubic with singularity type $2A_{1}A_{2}$.
Here, $m(\sigma)=0$, $\operatorname{Bour}(\sigma)=12$ and $c_{3}({\mathcal{T}}_{\sigma})=32$. Moreover, ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
given by the Buchsbaum–Rim complex.
2.2 Plane curves as pairs of surfaces
Let $g\in\kappa[x_{0},x_{1},x_{2}]$ be a square-free polynomial, so that the projective curve $X=V(g)\subset\mathbb{P}^{2}$ is reduced with isolated singularities, and the algebraically independent sequence $\sigma=(x_{3},g)$, denoting the plane $H=V(x_{3})\simeq\mathbb{P}^{2}$. We consider $S=V(g)\subset\mathbb{P}^{3}$ as a projective surface whose singular locus consists of possibly non-reduced lines, with the same multiplicity at each corresponding intersection point with $H$. Denote $Z_{g}\doteq\operatorname{Sing}(S)$ the associated singular scheme, given as the zero locus of the Jacobian ideal of $g$.
The matrix $\nabla\sigma$ will be given by
$\nabla\sigma=\begin{pmatrix}0&0&0&1\\ \partial_{0}g&\partial_{1}g&\partial_{2}g&0\end{pmatrix}$.
Let us denote by $\nabla\overline{g}=(\partial_{0}g,\partial_{1}g,\partial_{2}g)$ and
${\mathcal{T}}_{\overline{g}}\doteq\ker(\nabla\overline{g})\hookrightarrow\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}\xrightarrow{\nabla\overline{g}}\mathcal{O}_{\mathbb{P}^{3}}(d_{g})$
the kernel of the multiplication. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
${\mathcal{T}}_{\overline{g}}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 3}$$\mathcal{O}_{\mathbb{P}^{3}}(d)$$\mathcal{O}_{Z_{g}}(d)$${\mathcal{T}}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 4}$$\mathcal{O}_{\mathbb{P}^{3}}(d)\oplus\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{Q}_{\sigma}$$\mathcal{O}_{\mathbb{P}^{3}}$$\mathcal{O}_{\mathbb{P}^{3}}$$\scriptstyle{\nabla\overline{g}}$$\scriptstyle{\nabla\sigma}$$\scriptstyle{\cdot 1}$
and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$. Denoting $i:H\hookrightarrow\mathbb{P}^{2}$ the inclusion, it is clear that $i^{*}({\mathcal{T}}_{\sigma})\simeq{\mathcal{T}}_{g}$ as a logarithmic sheaf over $\mathbb{P}^{2}$, since $H$ and $Z_{g}$ intersect transversely.
From the discrete invariants considered $(d_{f}=0)$, we have the following formulas for the Bourbaki degrees of $\sigma$ and of $X=V(g)\subset\mathbb{P}^{2}$:
$\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}({\mathcal{T}}_{g}(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
where $\tau(X)$ is the Tjurina number of the curve $X$, since $d_{f}=0$.
From the transversality, we obtain $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:H\hookrightarrow\mathbb{P}^{3}$ and thus
$1-d[H]+(m_{0}-m(\sigma))[H]^{2}$ $=i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$
$=1-d[H]+(d^{2}-\tau(X))[H]^{2},$
and since $m_{0}=d_{g}^{2}$, we obtain the equality $m(\sigma)=\tau(X)$.
Moreover, doing the analogous comparison for the total classes of the twist ${\mathcal{T}}_{\sigma}(e)$, we obtain
$i^{*}c({\mathcal{T}}_{\sigma}(e))=1+(e-d)[H]+\operatorname{Bour}(\sigma)[H]^{2}=1+(e-d)[H]+\operatorname{Bour}(X)[H]^{2}=c(i^{*}{\mathcal{T}}_{\sigma}(e)),$
and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$. Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number does.
A more algebraic way to see this is through the associativity formula. Note that $\mathcal{Q}_{\sigma}\otimes\mathcal{O}_{H}\simeq\mathcal{O}_{\operatorname{Sing}(X)}(d)$, with the effect of adding a variable to the graded ring of the plane $\mathbb{P}^{2}\hookrightarrow\mathbb{P}^{3}$, and the associated map of modules is given by the quotient map
$0\to(x_{3})\to\kappa[x_{0},x_{1},x_{2},x_{3}]\to\kappa[x_{0},x_{1},x_{2}]\to 0.$
Then, for any associated prime $P$ of the Jacobian ideal of $X$ at $\kappa[x_{0},x_{1},x_{2}]$, the extension will satisfy
$\dim R/(R\cdot P)$ $=\dim(\kappa[x_{0},x_{1},x_{2}]/P)+1$
$\deg R/(R\cdot P)$ $=\deg(\kappa[x_{0},x_{1},x_{2}]/P).$
since the generators do not include $x_{3}$. Moreover, the length of the modules over $P$ will be preserved, as the element $x_{3}\not\in P$ will be invertible. Together with the associativity formula, we obtain:
$\tau(X)$ $=\sum_{P\in\operatorname{Ann}(\mathcal{O}_{\operatorname{Sing}X}(d)),\dim(\kappa[x_{0},x_{1},x_{2}]/P)=1}\operatorname{{length}}(\mathcal{O}_{\operatorname{Sing}X}(d))_{P}\cdot\deg(\kappa[x_{0},x_{1},x_{2}]/P)$
$=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(\kappa[x_{0},x_{1},x_{2},x_{3}]/P)=2}\operatorname{{length}}(\mathcal{Q}_{\sigma})_{P}\cdot\deg(R/P)=m(\sigma).$
The following example of the nature above is a sequence $\sigma=(x_{3},g)$ of degree $d$ such that $e=d$ but $m(\sigma)\neq 0$, showing that a converse for 8, $(a)$ is not possible without assuming the Buchsbaum–Rim resolution.
Example 11.
Let $f=x_{3}$ and $g=x_{0}x_{1}x_{2}^{2}+x_{0}^{4}+x_{1}^{4}$. The quartic plane curve $X=V(g)\subset\mathbb{P}^{3}_{[x_{0}:x_{1}:x_{2}]}$ is nodal and the minimum degree for a syzygy of $\nabla g$ is $e=d$.
Hence, $m(\sigma)=\tau(X)=1\neq 0$, accounting for the nodal singularity, but $e=d=3$.
2.3 Free resolutions and the Bourbaki scheme
In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [undefr, Theorem 2.1, (c)]. The next result follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$, as it is the case for $\mathbb{P}^{2}$.
Lemma 12.
Let $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ be a non-zero section with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and let $B\subset\mathbb{P}^{3}$ be the pure codimension $2$ subscheme associated to $\nu$ in a short exact sequence:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\xrightarrow{\pi}\mathcal{I}_{B}(e-d)\rightarrow 0.$
Then:
  • (a)
    Every free resolution for $\mathcal{I}_{B}$:
    $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\xrightarrow{\omega}\mathcal{I}_{B}\rightarrow 0$
    lifts for a free resolution of the form
    $0\rightarrow F_{2}(e-d)\rightarrow F_{1}(e-d)\rightarrow F_{0}(e-d)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\omega(e-d),\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0$
    for ${\mathcal{T}}_{\sigma}$.
  • (b)
    For a minimal free resolution of ${\mathcal{T}}_{\sigma}$ including the section $\nu$:
    $0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{(\lambda,\nu)}{\mathcal{T}}_{\sigma}\rightarrow 0,$
    it induces a free resolution for $\mathcal{I}_{B}$ of the form:
    $0\rightarrow F_{2}(d-e)\rightarrow F_{1}(d-e)\rightarrow{}F_{0}(d-e)\xrightarrow{\lambda(d-e)}\mathcal{I}_{B}\rightarrow 0.$
Proof.
To show $(a)$, we apply the functor $\operatorname{Hom}(F_{0}(e-d),-)$ to the short exact sequence
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\nu}{\mathcal{T}}_{\sigma}\xrightarrow{\pi}\mathcal{I}_{B}(e-d)\rightarrow 0,$
to get the exact piece:
$\operatorname{Hom}(F_{0}(e-d),{\mathcal{T}}_{\sigma})\xrightarrow{\pi^{*}}\operatorname{Hom}(F_{0}(e-d),\mathcal{I}_{B}(e-d))\rightarrow\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))=0,$
since $\operatorname{{Ext}}^{1}(F_{0}(e-d),\mathcal{O}_{\mathbb{P}^{3}}(-e))\simeq H^{1}(F_{0}^{\vee}(-2e-d))=0$, as $F_{0}^{\vee}$ is a direct sum of line bundles and these have vanishing first cohomology in $\mathbb{P}^{3}$. Thus, $\pi^{*}$ is surjective, and there is a morphism $\tilde{\omega}:F_{0}(e-d)\rightarrow{\mathcal{T}}_{\sigma}$ such that $\pi\circ\omega(e-d)=\tilde{\omega}$. We now consider the map $\omega(e-d)\oplus\nu$ in the following commutative diagram with short exact sequences as the central two columns:
[…]
From the snake lemma, we obtain that $\operatorname{coker}(\tilde{\omega}\oplus\nu)=0$ and that $\ker(\tilde{\omega}\oplus\nu)\simeq\ker(\omega(e-d))$. Thus, we can continue the resolution for $\mathcal{I}_{B}$, twisting by $\mathcal{O}_{\mathbb{P}^{3}}(e-d)$, to obtain the following free resolution:
[…]
for ${\mathcal{T}}_{\sigma}$, as claimed.
To show $(b)$, we consider the diagram with exact rows induced by the fact above to obtain the short exact sequence in cokernels as the third row below:
[…]
Completing to the resolution and twisting accordingly, we obtain the resolution from the claim below.
[…]
Definition 13.
Let $\sigma$ be a non-free normal sequence with degrees $d_{f}+1,d_{g}+1$. We say $\sigma$ is:
  • nearly free if $\operatorname{Bour}(\sigma)=1$.
  • 33-syzygy if there is a minimal free resolution for ${\mathcal{T}}_{\sigma}$ such that $\operatorname{{rk}}(F_{0})=2$ in the notation of Lemma 12, $(b)$.
Example 14 (Nearly free sequence with $d_{f}=1,d_{g}=2$).
We consider the following sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(x_{0}^{2}+x_{3}^{2},x_{0}^{3}+x_{0}x_{1}x_{2}+x_{3}^{3})$
with Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 3x_{0}^{2}+x_{1}x_{2}&x_{0}x_{2}&x_{0}x_{1}&3x_{3}^{2}\end{pmatrix}$.
The scheme $(\Xi_{\sigma})_{1}$ has three primary components in dimension one, given by the prime ideals $\mathfrak{p}_{1}=(x_{0},x_{3})$, $\mathfrak{p}_{2}=(x_{0},x_{2})$ and $\mathfrak{p}_{3}=(x_{0},x_{1})$. We also note that there is an easy syzygy of degree one, namely $\nu=(-x_{3},0,0,x_{0})^{T}$, hence $e=1$.
To show $m(\sigma)=4$, we study the length of the cokernel module $Q$ associated to the sheaf $\mathcal{Q}_{\sigma}$ over each prime $\mathfrak{p}_{i}$, $i=1,2,3$, and use the associativity formula.
Over $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that $x_{1}x_{2}\in R_{P}^{\times}$ is a unit, and $3x_{0}^{2}\in\mathfrak{p}_{P}$ is in the maximal ideal, hence $v=3x_{0}^{2}+x_{1}x_{2}$ is also invertible, denote $u=v^{-1}$. From this, we may rewrite the matrix as
$(\nabla\sigma)_{\mathfrak{p}_{1}}=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 1&ux_{0}x_{2}&ux_{0}x_{1}&3ux_{3}^{2}\\ \end{pmatrix}\sim\begin{pmatrix}0&-2x_{0}^{2}x_{2}u&-2ux_{0}^{2}x_{1}&2x_{3}-6ux_{0}x_{3}^{2}\\ 1&x_{0}x_{2}u&x_{0}x_{1}u&3x_{3}^{2}u\\ \end{pmatrix}$,
first by multiplying the second row by $u$, and then adding the second row scaled by $2x_{0}$ to the first row. From the final form of the matrix above, it is easy to see that it sends the fourth basis vector to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and thus we may compute the cokernel as the remaining entries of the first row:
$\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{(-2x_{0}^{2}x_{2}u,-2ux_{0}^{2}x_{1},2x_{3}-6ux_{0}x_{3}^{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{0}^{2},x_{3})},$
a $R_{\mathfrak{p}_{1}}$-module of length two.
Over the prime $\mathfrak{p}_{2}=(x_{0},x_{2})$, the element $x_{3}\in R_{\mathfrak{p}_{2}}$ is an unit. We denote by $u=x_{3}^{-1}\in R_{\mathfrak{p}_{2}}$ its inverse, and we rewrite the matrix using elementary operations
$(\nabla\sigma)_{\mathfrak{p}_{2}}=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 3x_{0}^{2}+x_{1}x_{2}-3x_{0}x_{3}&x_{0}x_{2}&x_{0}x_{1}&0\\ \end{pmatrix}\sim\begin{pmatrix}ux_{0}&0&0&1\\ 3x_{0}^{2}+x_{1}x_{2}-3x_{0}x_{3}&x_{0}x_{2}&x_{0}x_{1}&0\\ \end{pmatrix}$,
first by adding the first row scaled by $(-\frac{3}{2}x_{3})$ to the second row, and then multiplying the first row by $u/2$. Now, it is clear that the matrix sends the fourth basis vector to $(1,0)\in R_{P}^{2}$, and thus
$\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0}x_{1},x_{0}x_{2},3x_{0}^{2}+x_{1}x_{2}-x_{0}x_{3})}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{0}x_{2},3x_{0}^{2}+x_{1}x_{2}-x_{0}x_{3})}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{0}x_{2},x_{1}x_{2})}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{2})}\simeq\kappa$
since $x_{1}\in R_{\mathfrak{p}_{2}}^{\times}$. Thus, it follows $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$. For the ideal $\mathfrak{p}_{3}$, we note that the approach is analogous as the previous one, using the same elementary operations, and obtaining the same isomorphism at the end $\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{3}}\simeq\kappa$, so that from the associativity formula:
$m(\sigma)=\sum_{i=1}^{3}\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{i}})\cdot\deg(R/\mathfrak{p}_{i})=2+1+1=4.$
Using the Bourbaki formula with $m_{0}=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}=7$, $d=3$ and the obtained discrete invariants $m(\sigma)=4$ and $e=1$, we conclude that $\operatorname{Bour}(\sigma)=1$.
From the minimal free resolution of the ideal of a line in $\mathbb{P}^{3}$, using Lemma 12, the minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
From the resolution, one obtains $c_{3}({\mathcal{T}}_{\sigma})=3$.
Example 15 ($3$-syzygy pencil of cubics which is not nearly-free).
Considering the following pencil of cubics:
$\sigma=(x_{2}x_{3}(x_{0}-x_{1}),x_{0}(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}))$
where $f$ is a hyperplane arrangement and $g$ is the union of a plane and a smooth quadric, with the Jacobian matrix:
$\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines and a plane quadric, given by the following prime ideals of $R$:
$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
$\mathfrak{p}_{2}$ $=(x_{3},x_{0}-x_{1})$
$\mathfrak{p}_{3}$ $=(x_{2},x_{0}-x_{1})$
$\mathfrak{p}_{4}$ $=(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
We claim that $m(\sigma)=5$, a quantity which will be obtained using the associativity formula. For $\mathfrak{p}_{1}=(x_{2},x_{3})$, we start by noting that the quantities:
$h$ $=3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$
$v$ $=x_{0}x_{1}$
$t$ $=(x_{0}-x_{1})$
are all invertible inside the ring $R_{\mathfrak{p}_{1}}$, where the first one is the sum of an invertible element with an element inside the maximal ideal $(x_{2},x_{3})_{\mathfrak{p}_{1}}$. Then, we do two elementary operations:
$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}ux_{2}x_{3}+\frac{1}{2}hx_{2}x_{3}&0&utx_{3}+x_{0}x_{2}^{2}x_{3}&utx_{2}+x_{0}x_{2}x_{3}^{2}\\ h&2v&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$
$\sim\begin{pmatrix}ux_{2}x_{3}+\frac{1}{2}hx_{2}x_{3}&0&utx_{3}+x_{0}x_{2}^{2}x_{3}&utx_{2}+x_{0}x_{2}x_{3}^{2}\\ \frac{1}{2}uh&1&ux_{0}x_{2}&ux_{0}x_{3}\end{pmatrix}$.
The first is to add $\frac{1}{2}x_{2}x_{3}$ times the second row to the first row, scaled by $u=v^{-1}$. Then, we scale the second row by $\frac{1}{2}u$. Since the matrix sends the second basis vector into $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, it suffices to look at the first row of the remaining matrix, so that
$\operatorname{coker}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{((u+\frac{1}{2}h)x_{2}x_{3},utx_{3}+x_{0}x_{2}(x_{2}x_{3}),utx_{2}+x_{0}x_{3}(x_{2}x_{3}))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2}x_{3},utx_{3}+x_{0}x_{2}(x_{2}x_{3}),utx_{2}+x_{0}x_{3}(x_{2}x_{3}))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2}x_{3},utx_{3},utx_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2},x_{3})}\simeq\kappa,$
and hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
For $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, note that $v=x_{0}x_{2}\in R_{\mathfrak{p}_{2}}$ is a unit and denote $u=v^{-1}$. We consider some elementary operations:
$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})h&-x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{1}&0&x_{2}(x_{0}-x_{1})-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{3}\\ h&2x_{0}x_{1}&2v&2x_{0}x_{3}\end{pmatrix}$
$\sim\begin{pmatrix}x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})h&-x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{1}&0&x_{2}(x_{0}-x_{1})-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{3}\\ \frac{1}{2}uh&ux_{0}x_{1}&1&ux_{0}x_{3}\end{pmatrix}$.
The first transformation is to add the scaled second row (by $\frac{1}{2}x_{3}(x_{0}-x_{1})u$) to the first one, and the second transformation is rescaling the second row by $\frac{1}{2}u$. Then, we see that the matrix sends the third basis vector to $(0,1)\in R_{\mathfrak{p}_{2}}^{2}$, so we may compute the cokernel as the first row of the remaining matrix:
$\operatorname{coker}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3}(x_{2}-\frac{1}{2}u(x_{0}-x_{1})h),-x_{3}(x_{2}+\frac{1}{2}u(x_{0}-x_{1})x_{0}x_{1}),(x_{0}-x_{1})(x_{2}-\frac{1}{2}ux_{0}x_{3}^{2})))}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{x_{3},-x_{3},(x_{0}-x_{1})}\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{0}-x_{1})}\simeq\kappa,$
since for example the quantity
$x_{2}-\frac{1}{2}u(x_{0}-x_{1})h$
is of the form a unit $x_{2}$ plus something inside the maximal ideal $(x_{3},x_{0}-x_{1})_{\mathfrak{p}_{2}}$, it is also a unit. Thus, we obtain $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. An analogous consideration also holds for the case of the prime $\mathfrak{p}_{3}=(x_{2},x_{0}-x_{1})$, with similar elementary operations, now to trivialize the fourth column as $(0,1)$, and then to show that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa$.
For the prime $\mathfrak{p}_{4}=(x_{0},h\doteq x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$, we note that $v=x_{2}x_{3}$ is invertible in $R_{\mathfrak{p}_{4}}$, denote $u=v^{-1}$ and do the elementary operations
$(\nabla\sigma)_{\mathfrak{p}_{4}}$ $\sim\begin{pmatrix}v&-v&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\\ h+2x_{0}x_{1}&0&2x_{0}x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})&2x_{0}x_{2}x_{3}(x_{0}-x_{1})\\ \end{pmatrix}$
$\sim\begin{pmatrix}-1&1&-ux_{3}(x_{0}-x_{1})&-ux_{2}(x_{0}-x_{1})\\ h+2x_{0}x_{1}&0&2x_{0}x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})&2x_{0}x_{2}x_{3}(x_{0}-x_{1})\\ \end{pmatrix}$,
first adding the scaled first row $(by2x_{0}x_{1}u)$ to the second row, and then scaling the first row by u-u. The matrix sends the second basis vector to $(1,0)\in R_{\mathfrak{p}_{4}}^{2}$, and thus to compute the cokernel we may consider only the second row of the remaining matrix, so that
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}$ $\simeq\frac{R_{\mathfrak{p}_{4}}}{(h+2x_{0}x_{1},x_{0}(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})),2x_{0}x_{2}x_{3}(x_{0}-x_{1}))}$
$\simeq\frac{R_{\mathfrak{p}_{4}}}{(h+2x_{0}x_{1},x_{0},x_{0})}$
$\simeq\frac{R_{\mathfrak{p}_{4}}}{(x_{0},h)}\simeq\kappa,$
noting that the factors $(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1}))$ and $2x_{2}x_{3}(x_{0}-x_{1})$ are units in the local ring. Thus, we obtain that the length at each prime is one, but one of them has multiplicity two, therefore $m(\sigma)=5$ from the associativity formula. The main result of the next section is:
$e=1$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 2$
$e=2$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 5,$
so for $m_{0}=12$ we conclude that these two correspond to $m(\sigma)\geq 7$ and $m(\sigma)\geq 6$, respectively. Since $m(\sigma)=5$, we conclude automatically that $e\geq 3$.
Using Macaulay2, we obtain a minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the 33-syzygy kind, given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so $e=3$ and, with $m(\sigma)=5$, we conclude $\operatorname{Bour}(\sigma)=4$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for the Bourbaki ideal $\mathcal{I}_{B}$ will be of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}\rightarrow\mathcal{I}_{B}\rightarrow 0,$
presenting $B$ as a complete intersection of two quadric surfaces in $\mathbb{P}^{3}$.
For the rest of the section, we study some aspects of these special classes of sequences.
Proposition 16.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}+1$, $d_{g}+1$. Then:
  • (a)
    $\sigma$ is nearly free if and only if the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
    $0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(e-d-1)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
    where $d=d_{f}+d_{g}$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
  • (b)
    if $\sigma$ is nearly free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
Proof.
This follows from Lemma 12, using the minimal free resolution for a line in $l\subset\mathbb{P}^{3}$ as the intersection of two planes:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)^{\oplus 2}\rightarrow\mathcal{I}_{l}\rightarrow 0$
since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for a choice of syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. The isolated zeros of the minors will be in the support of the sheaf $\operatorname{supp}({\mathcal{T}}_{\sigma})$, and these will be inside the line $l$ from the resolution obtained. ∎
Remark 17.
The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [undefd], related to rational cuspidal curves. In [undefr, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [undefd] (see [undefr, Definition 2.17]). Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
The notion of 33-syzygy divisors is also present in a number of previous works, for example [undef], [undefe] and [undefg].
Example 18.
Consider $f=x_{0}^{3}+x_{0}x_{1}x_{2}+x_{3}^{3}$ and $g=x_{0}^{k+1}+x_{3}^{k+1}$ for $k\geq 2$. Then $d=d_{f}+d_{g}=k+2$ and:
  • For $k=2$, $(f,g)$ is a free pencil of cubics with $e=1$, $m(\sigma)=9$;
  • For $k>2$, $(f,g)$ is nearly free with $e=1$.
The Jacobian matrix for $\sigma=(f,g)$ is
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{2}&x_{0}x_{2}&x_{0}x_{1}&3x_{3}^{2}\\ (k+1)x_{0}^{k}&0&0&(k+1)x_{3}^{k}\end{pmatrix}$.
For $k=2$, we note that the matrix below:
$\begin{pmatrix}0&-x_{0}x_{3}^{2}\\ x_{1}&0\\ -x_{2}&x_{2}x_{3}^{2}\\ 0&x_{0}^{3}\end{pmatrix}$
gives trivializing syzygies such that ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
For $k>2$, we set
$M=\begin{pmatrix}0&-x_{0}x_{2}x_{3}^{k}&-x_{0}x_{1}x_{3}^{k}\\ x_{1}&-3x_{0}^{k}x_{3}^{2}+3x_{0}^{2}x_{3}^{k}&0\\ -x_{2}&x_{2}^{2}x_{3}^{k}&-3x_{0}^{k}x_{3}^{2}+3x_{0}^{2}x_{3}^{k}+x_{1}x_{2}x_{3}^{k}\\ 0&x_{0}^{k+1}x_{2}&x_{0}^{k+1}x_{1}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}^{k}x_{3}^{2}-x_{0}^{2}x_{3}^{k}\\ \frac{1}{3}x_{1}\\ -\frac{1}{3}x_{2}\end{pmatrix}$,
so we obtain a free resolution of ${\mathcal{T}}_{\sigma}$ given by
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-1)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)^{\oplus 2}\xrightarrow{M}\mathcal{T}_{\sigma}\rightarrow 0,$
and in particular $\sigma$ is a nearly free sequence.
Proposition 19.
If $\sigma=(f,g)$ is a nearly free normal sequence, then ${\mathcal{T}}_{\sigma}$ is not locally free.
Proof.
Since
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e),$
assuming ${\mathcal{T}}_{\sigma}$ is locally free, we obtain $c_{3}({\mathcal{T}}_{\sigma})=0$. On the other hand, since $\sigma$ is nearly free, $\deg(B)=1$ and $p_{a}(B)=0$, hence
$e=\frac{d+2}{2},$
which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$, hence ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable.
If we denote by $E\doteq{\mathcal{T}}_{\sigma}(d/2)$ the normalized vector bundle, we must have $c_{1}(E)=0$ and
$c_{2}(E)$ $=c_{2}({\mathcal{T}}_{\sigma})-\frac{d^{2}}{2}+\frac{d^{2}}{4}$
$=m_{0}-m(\sigma)-\frac{d^{2}}{4}.$
On the other hand, using the Bourbaki degree formula with $\operatorname{Bour}(\sigma)=1$ and $e=\frac{d+2}{2}$, we obtain
$1=\operatorname{Bour}(\sigma)$ $=\frac{(d+2)^{2}}{4}-\frac{d(d+2)}{2}+m_{0}-m(\sigma)$
$=\frac{d^{2}}{4}-\frac{d^{2}}{2}-d+\frac{4d}{4}+1+m_{0}-m(\sigma),$
which yields
$m_{0}-m(\sigma)-\frac{d^{2}}{4}=0,$
so that $c_{2}(E)=0$. But a $\mu$-semistable reflexive sheaf with $c_{1}(E)=c_{2}(E)=0$ must be $E\simeq\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 2}$ ([undefn, Lemma 9.7]), implying $\sigma$ is free, contradicting the hypothesis. ∎
Proposition 20.
If a normal sequence $\sigma=(f,g)$ is a $3$-syzygy, then $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$. Moreover, a sequence $\sigma$ is $3$-syzygy if and only if $B_{\nu}$ is a complete intersection, for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
Proof.
First, if we assume $\sigma$ is $3$-syzygy, then there is a free resolution of the form:
$0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\lambda}{\mathcal{T}}_{\sigma}\rightarrow 0,$
so we split the resolution into two short exact sequences:
$F_{2}\hookrightarrow F_{1}\twoheadrightarrow S\text{ and }S\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma},$
and focus on the second one. The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [undefn, Proposition 1.1]. Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
$S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma}$
is a free resolution for ${\mathcal{T}}_{\sigma}$, concluding $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$.
For the equivalence stated above, if we start with a sequence $\sigma$ which is $3$-syzygy and then apply Lemma 12, $(b)$, we obtain a resolution for $\mathcal{I}_{B}$ which is of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
since $\operatorname{{rk}}(F_{0})=2$, thus concluding $B$ must be a complete intersection scheme. On the other hand, if $B$ is a complete intersection, then there is a resolution for $\mathcal{I}_{B}$ of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\xrightarrow{r,s}\mathcal{I}_{B}\rightarrow 0,$
induced by the two equations $r,s$ defining $B=V(r,s)\subset\mathbb{P}^{3}$. Thus, applying Lemma 12, $(a)$, we can lift the resolution above for a free resolution for ${\mathcal{T}}_{\sigma}$ of the form
$0\rightarrow F_{1}\rightarrow F_{0}^{\prime}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $\operatorname{{rk}}(F_{0}^{\prime})=3$, thus $\operatorname{{rk}}(F_{0})=2$ and $\sigma$ is $3$-syzygy. ∎
Remark 21.
We note that there is a chain of implications:
$\sigma\text{ is nearly free }\Rightarrow\sigma\text{ is $3$-syzygy }\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
where the first follows from 20 since $\sigma$ is nearly free iff $B_{\nu}$ is a line for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and every line is a complete intersection of two planes. The converses do not hold, as we explore in the next examples: there are $3$-syzygy pencils of cubics which are not nearly free (Example 15), pencils of cubics which satisfy $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but are not $3$-syzygy (Example 22) and pencils of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (Example 23).
We also construct two pencils of cubics with the same discrete invariants $(m(\sigma),\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma))$ and the same Chern classes, which are distinguished by their homological behavior: one is $3$-syzygy and the other satisfies $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (see Example 22 and Example 24).
Example 22 (pencil of cubics which is not $3$-syzygy and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$).
Considering the following pencil of cubics:
$\sigma=(x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{2}^{3}+x_{1}x_{2}x_{3}+x_{3}^{3})$
with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\\ 0&x_{2}x_{3}&3x_{2}^{2}+x_{1}x_{3}&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$.
The scheme $(\Xi_{\sigma})_{1}$ has three irreducible components, given by two lines and a plane quadric curve:
$L_{1}$ $=V(x_{0},x_{3})\doteq V(\mathfrak{p}_{1})$
$L_{2}$ $=V(x_{2},x_{3})=V(\mathfrak{p}_{2})$
$Q$ $=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})=V(\mathfrak{p}_{3}),$
Now, we proceed to computing the length of the cokernel module $Q$ over each prime ideal $\mathfrak{p}_{i}$, in order to show that $m(\sigma)=5$ using the associativity formula.
Starting with the prime ideal $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that the element $3x_{2}^{2}\in R_{\mathfrak{p}_{1}}$ is invertible, and hence the sum $v=3x_{2}^{2}+x_{1}x_{3}$ is also invertible, as $x_{1}x_{3}\in\mathfrak{p}_{1}$. Let $u=v^{-1}$, so we perform two elementary operations on the matrix $\nabla\sigma$ over $\mathfrak{p}_{1}$:
$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&x_{2}x_{3}&v&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$
$\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&ux_{2}x_{3}&1&u(x_{1}x_{2}+3x_{3}^{2})\end{pmatrix}$.
The first transform is to add the scaled second row (by $ux_{0}^{2}$) to the first, and then we scale the second row by $u$. Then, the third basis vector goes via $\nabla\sigma$ to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and therefore we may compute the cokernel from the quotient
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(2x_{0}x_{2}+x_{1}x_{3},x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3},x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2}))}\simeq\frac{R_{\mathfrak{p}_{1}}}{I},$
where $I=(f_{1},f_{2},f_{3})$ are the generators above. We may factor $f_{2}=x_{0}x_{3}(1-ux_{0}x_{2})$, so that since $(1-ux_{0}x_{2})$ is a unit in the local ring, we conclude $x_{0}x_{3}\in I$. From this, multiplying $x_{3}f_{1}$ we conclude that $x_{1}x_{3}^{2}\in I$, and since $x_{1}$ is also a unit, we obtain $x_{3}^{2}\in I$, and analogously taking $x_{0}f_{1}$ we obtain $x_{0}^{2}x_{2}\in I$, and since $x_{2}$ is a unit we get $x_{0}^{2}\in I$. Thus
$(x_{0}^{2},x_{0}x_{3},x_{3}^{2})\subset I.$
Now, from this, looking at $f_{3}$:
$f_{3}=x_{0}x_{1}+3x_{3}^{2}-x_{0}^{2}u(x_{1}x_{2}+3x_{3}^{2})\in I\Rightarrow x_{0}x_{1}\in I\Rightarrow x_{0}\in I,$
and looking at $f_{1}$, we obtain
$f_{1}=2x_{0}x_{2}+x_{1}x_{3}\in I\Rightarrow x_{1}x_{3}\in I\Rightarrow x_{3}\in I,$
so that $I=(x_{0},x_{3})$ and thus $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\kappa$, with $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
For the case $\mathfrak{p}_{2}=(x_{2},x_{3})$, let $u=(x_{0}^{2})^{-1}$, $q=(3x_{2}^{2}+x_{1}x_{3})$ and consider the elementary operations given by $R_{2}^{\prime}=R_{2}-uqR_{1}$ and then $R_{1}^{\prime}=uR_{1}$:
$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&u(x_{0}x_{1}+3x_{3}^{2})\\ -uq(2x_{0}x_{2}+x_{1}x_{3})&x_{2}x_{3}-uqx_{0}x_{3}&0&x_{1}x_{2}+3x_{3}^{2}-uq(3x_{3}^{2}+x_{0}x_{1})\end{pmatrix}$.
The Jacobian matrix sends the third basis vector to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, and thus we may compute:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I}$
where
$f_{1}$ $=(3x_{2}^{2}+x_{1}x_{3})(2x_{0}x_{2}+x_{1}x_{3})$
$f_{2}$ $=x_{3}(x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3})$
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3ux_{3}^{2}(3x_{2}^{2}+x_{1}x_{3})-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})$
First, we consider
$f_{2}-\frac{x_{3}}{x_{1}}f_{3}=\frac{3x_{3}^{3}}{x_{1}}(1-u(3x_{2}^{2}+x_{1}x_{3}))\in I,$
and thus $x_{3}^{3}\in I$. Then
$x_{3}f_{2}=x_{3}^{2}x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3}^{3}\in I\Rightarrow x_{3}^{2}x_{2}\in I,$
and moreover
$x_{3}f_{1}$ $=x_{2}x_{3}(6x_{0}x_{2}^{2}+3x_{1}x_{2}x_{3}+2x_{0}x_{1}x_{3})+x_{1}^{2}x_{3}^{3}\in I$
$\Rightarrow 6x_{0}x_{2}^{3}x_{3}+x_{3}^{2}x_{2}(3x_{1}x_{2}+2x_{0}x_{1})\in I$
$\Rightarrow 6x_{0}x_{2}^{3}x_{3}\in I$
$\Rightarrow x_{2}^{3}x_{3}\in I.$
On the other hand,
$x_{2}x_{3}f_{3}$ $=-3ux_{3}^{3}q+3x_{2}x_{3}^{3}+x_{2}x_{3}(x_{1}x_{2}-ux_{0}x_{1}q)\in I$
$\Rightarrow x_{1}x_{2}^{2}x_{3}-3ux_{0}x_{1}x_{2}^{3}x_{3}-ux_{0}x_{1}^{2}x_{2}x_{3}^{2}\in I$
$\Rightarrow x_{1}x_{2}^{2}x_{3}\in I\Rightarrow x_{2}^{2}x_{3}\in I.$
Now, consider
$x_{3}f_{3}$ $=x_{1}x_{2}x_{3}+3x_{3}^{3}-uqx_{3}^{3}-uqx_{0}x_{1}x_{3}\in I$
$\Rightarrow x_{3}x_{2}(x_{2}-3ux_{0}x_{2}^{2}-ux_{0}x_{1}x_{3})\in I$
$\Rightarrow x_{1}x_{3}(x_{2}-ux_{0}x_{1}x_{3})\doteq A\in I,$
and
$x_{3}A\in I\Rightarrow x_{1}x_{3}^{2}x_{2}-ux_{0}x_{1}^{2}x_{3}^{2}\in I\Rightarrow x_{3}^{2}\in I.$
This simplifies
$f_{2}=x_{3}(x_{2}-uqx_{0})\in I\Rightarrow x_{2}x_{3}(1-3ux_{2})\in I\Rightarrow x_{2}x_{3}\in I.$
Considering
$f_{1}$ $=6x_{0}x_{2}^{3}+x_{2}x_{3}(3x_{1}x_{2}+2x_{0}x_{1})+x_{1}^{2}x_{3}^{2}\in I$
$\Rightarrow x_{2}^{3}\in I,$
and then
$x_{2}f_{3}$ $=x_{1}x_{2}^{2}+x_{2}(3x_{3}^{2}-3ux_{3}^{2}q)-uqx_{0}x_{1}x_{2}\in I$
$\Rightarrow x_{1}x_{2}^{2}-3ux_{0}x_{1}x_{2}^{3}-ux_{0}x_{1}^{2}x_{2}x_{3}\in I$
$\Rightarrow x_{1}x_{2}^{2}\in I\Rightarrow x_{2}^{2}\in I,$
simplifying
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3uqx_{3}^{2}-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})\in I$
$\Rightarrow x_{1}(x_{2}-ux_{0}x_{1}^{2}x_{3})\in I\Rightarrow x_{2}-ux_{0}x_{1}x_{3}\in I,$
so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa\langle 1,x_{2}\rangle$, meaning $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=2$, since the equations $f_{1},f_{2}$ are trivialized after $x_{3}^{2},x_{2}^{2},x_{2}x_{3}\in I$.
Let us consider the case $\mathfrak{p}_{3}=(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})\doteq(g_{1},g_{2})$. One may be tempted to choose $u=(2x_{0}x_{2}+x_{1}x_{3})^{-1}$ and do the elementary operations $R_{1}^{\prime}=uR_{1}$, but
$2x_{0}x_{2}+x_{1}x_{3}=2x_{0}g_{1}+g_{2}.$
Take $u\doteq(x_{0}^{2})^{-1}\in R_{\mathfrak{p}_{3}}$. If $x_{0}^{2}\in\mathfrak{p}_{3}$, then:
$x_{0}\in\mathfrak{p}_{3}\Rightarrow x_{2}\in\mathfrak{p}_{3}\Rightarrow x_{1}x_{3}\in\mathfrak{p}_{3}$
and $\text{height}(\mathfrak{p}_{3})\geq 3$ would give a contradiction. Thus, we may perform the same elementary operations as we did for $\mathfrak{p}_{2}$, and thus
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{I}$
with generators $I=(f_{1},f_{2},f_{3})$ given as before. By the same reasoning, $(3x_{2}^{2}+x_{1}x_{3})\not\in\mathfrak{p}_{3}$, and thus
$f_{1}\in I\Rightarrow 2x_{0}x_{2}+x_{1}x_{3}\in I.$
From the previous description of the element above, it suffices to show $g_{i}\in I$ for some $i\in\{1,2\}$ to conclude $I=\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$. Let $A\doteq(g_{2}+x_{2}^{2})$, so that $u=A^{-1}$. We can rewrite $f_{3}$ as:
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-u(g_{2}+x_{2}^{2})(3x_{3}^{2}+x_{0}x_{1})$
$=3x_{3}^{2}(1-u(g_{2}+x_{2}^{2}))+x_{1}(x_{2}-ux_{0}(g_{2}+x_{2}^{2}))$
$=3x_{3}^{2}(1-1)+x_{1}(x_{2}-ux_{0})$
$=-x_{1}g_{1}\in I,$
and since $x_{1}\notin\mathfrak{p}_{4}$, $g_{1}\in I$ and $I=\mathfrak{p}_{4}R_{\mathfrak{p}_{4}}$, concluding $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=1$. From the associativity formula, we obtain
$m(\sigma)=1\cdot 1+2\cdot 1+1\cdot 2=5.$
Since $m(\sigma)\leq 6$, we conclude using the bounds from the next section that $e\geq 3$ The minimal free resolution of ${\mathcal{T}}_{\sigma}$ given using Macaulay2 is
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{M}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$ and $\operatorname{Bour}(\sigma)=4$. This means $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$, but we need $4$ syzygies to generate ${\mathcal{T}}_{\sigma}$ or, equivalently, the degree four Bourbaki scheme is not a complete intersection.
Example 23 (pencil of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$).
We consider the sequence of cubics $d_{f}=d_{g}=2$ given by:
$\sigma=(x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3},x_{2}x_{3}(x_{2}-x_{1})),$
considered in [undefj, Theorem 8.1]. From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle. Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
So $e=3$ and $m(\sigma)=7$. The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$. Localizing the cokernel at the first two lines gives length one, so over last line should have length $5$.
Example 24 (pencils of cubics with the same discrete invariants and different homological behavior).
Considering the following pencil of cubics:
$\sigma=(x_{0}^{3}+x_{0}x_{1}x_{3}+x_{3}^{3},x_{3}^{3}+x_{1}x_{3}^{2}+x_{0}x_{1}x_{3}+x_{0}^{2}x_{2}),$
with the associated Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}+x_{3}^{2}&x_{0}^{2}&x_{0}x_{1}+2x_{1}x_{3}+3x_{3}^{2}\end{pmatrix}$.
Here, the unique irreducible component of $(\Xi_{\sigma})_{1}$ is the line $V(x_{0},x_{3})$.
Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 6}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$ and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$. A resolution for $\mathcal{I}_{B}$ will be of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 7}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 6}\rightarrow\mathcal{I}_{B}\rightarrow 0.$
This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as Example 15, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
In Section 4 we characterize all nearly free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly free sequences with $d_{f}=1,d_{g}=2$.
3 Extreme cases of low initial degree
In this section, we observe that a non-zero section $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ induces a sub-foliation by curves of degree $e+1$ of the foliation ${\mathcal{T}}_{\sigma}(1)$. We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two. From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 27. We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example, [undefc]).
As explored in [undefj, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
$0\rightarrow{\mathcal{T}}_{\sigma}(1)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the associated foliation. Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
[…]
The middle column of the previous diagram
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(1-e)\rightarrow\mathbb{T}\mathbb{P}^{3}\rightarrow G\rightarrow 0$
defines a foliation by curves of $\mathbb{P}^{3}$ of degree $e$, where $G$ is a rank two torsion-free sheaf and $G^{\vee}$ is called the conormal sheaf of the foliation. Dualizing this short exact sequence, we obtain
$0\rightarrow G^{\vee}\rightarrow\Omega^{1}_{\mathbb{P}^{3}}\rightarrow\mathcal{I}_{W}(e-1)\rightarrow 0,$
defining a subscheme $W\subset\mathbb{P}^{3}$, called the singular scheme of the associated foliation by curves. It has codimension at least two and it is also described by ${\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})\simeq\mathcal{O}_{W}$.
The classification of such foliations by curves of low degree provides the following bounds on their singular schemes:
  • (a)
    ([undefk, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1. In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
  • (b)
    (in preparation, V. Cordeiro) If $e=2$, then $\deg(W)\leq 5$ or is zero-dimensional.
For completeness, we include an argument for $(b)$, from V. Cordeiro: using the short exact sequence
$0\to G^{\vee}\to\Omega^{1}_{\mathbb{P}^{3}}\to\mathcal{I}_{W}(1)\to 0,$
one obtains $4\leq c_{2}(G^{\vee})=11-\deg(W)\leq 11$ (see, for example, [undefc, 4.1, 4.2]). Now, if $G^{\vee}$ is stable, Bogomolov’s inequality says $c_{2}(G^{\vee})\geq 7$. Now, assuming that $G^{\vee}$ is not stable, $h^{0}(G^{\vee}(2))\neq 0$ and one may choose a section to form a sequence
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\to G^{\vee}\to\mathcal{I}_{Y}(-3)\to 0$
where $Y$ is the vanishing locus of this section. Hence, from additivity one gets $0\leq\deg(Y)=c_{2}(\mathcal{I}_{Y})=c_{2}(G^{\vee})-6$. From the first inequality, we conclude $c_{2}(G^{\vee})\geq 6$, and thus $\deg(W)\leq 5$.
Theorem B.
Let $\sigma=(f,g)$ be a normal sequence of polynomials of degrees $d_{f}+1,d_{g}+1$. Then:
  • (a)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, then $\operatorname{Bour}(\sigma)\in\{0,1,2\}$;
  • (b)
    If $\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then $\operatorname{Bour}(\sigma)\leq 5$.
Proof.
To show $(a)$, we dualize the following short exact sequence, obtained above for $e=1$:
$0\rightarrow\mathcal{I}_{B}(2-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0,$
to get a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-2)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+2)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}$ is defined by the short exact sequence $0\to\mathcal{U}\to\mathcal{O}_{\Gamma_{\sigma}}\to\mathcal{O}_{Y}\to 0$ and $Y$ is the one-dimensional component of $\Gamma\sigma$. Moreover, since $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$, we may consider the final piece of the long exact sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+2)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$, and from the classification $\deg(W^{\prime})\leq\deg(W)\leq 2$, hence the result follows.
For $(b)$, we proceed analogously, dualizing the sequence
$0\rightarrow\mathcal{I}_{B}(3-d)\rightarrow G\rightarrow\mathcal{I}_{\Gamma_{\sigma}}(d+2)\rightarrow 0$
to obtain a long exact sequence, after simplifying, of the form:
$0$ $\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-d-2)\rightarrow G^{\vee}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(d-3)\rightarrow$
$\rightarrow\omega_{Y}(-d+2)\rightarrow\mathcal{O}_{W}\rightarrow\omega_{B}(d+1)\rightarrow$
$\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $\mathcal{U}\hookrightarrow\mathcal{O}_{\Gamma_{\sigma}}\rightarrow\mathcal{O}_{Y}$ and $Y$ is the one-dimensional component of $\Gamma_{\sigma}$. Moreover, $\mathcal{O}_{W}\simeq{\mathcal{E}}{\it xt}^{1}(G,\mathcal{O}_{\mathbb{P}^{3}})$ and from the final piece of the sequence
$0\rightarrow\mathcal{O}_{W^{\prime}}\rightarrow\omega_{B}(d+1)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
where $W^{\prime}\subset W$ is a pure one-dimensional subscheme. Since the support of ${\mathcal{E}}{\it xt}^{3}(\mathcal{O}_{\Gamma_{\sigma}},\mathcal{O}_{\mathbb{P}^{3}})$ is a zero-dimensional scheme, comparing the supports we conclude $\deg(B)=\deg(W^{\prime})$. Since $W^{\prime}$ is a subscheme of $W$, we have $\deg(W^{\prime})\leq\deg(W)$. The classification of foliations by curves of degree $2$ implies $\deg(W)\leq 5$, which completes the proof. ∎
To be able to present the next applications, we will need the following result on the structure of degree $2$ space curves in $\mathbb{P}^{3}$:
Theorem 25.
[undeft, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree 22 and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
  • (a)
    $a\geq-1$, and a=1a=-1 if and only if $B$ is planar;
  • (b)
    For $a\geq 1$, $B$ must be a multiplicity two structure at a line $L\subset\mathbb{P}^{3}$, and these satisfy a short exact sequence of the form
    $0\rightarrow\mathcal{O}_{L}(a)\rightarrow\mathcal{O}_{B}\rightarrow\mathcal{O}_{L}\rightarrow 0;$
  • (c)
    For $a\geq 1$, if $B$ is a multiplicity two structure on a line, then $\omega_{B}\simeq\mathcal{O}_{B}(-a-2)$. If $a=0$ and $B$ is a union of two skew lines, then $\omega_{B}\simeq\mathcal{O}_{B}(-2)$.
Proof.
Items $(a)$ and $(b)$ are in the original paper, and to obtain $(c)$, for $a\geq 1$ we can apply the functor $\operatorname{\mathcal{H}\textit{om}}(-,\omega_{\mathbb{P}^{3}})$ to the short exact sequence in $(ii)$, use that $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ and twist by $a+2$ to obtain a short exact sequence
$0\rightarrow\mathcal{O}_{L}(a)\rightarrow\omega_{B}(a+2)\rightarrow\mathcal{O}_{L}\rightarrow 0$
which coincides with the short sequence in $(b)$, hence we obtain the isomorphism in the claim. For the case $a=0$, $B$ must be either a multiplicity two structure as in $(b)$ with $a=0$ or a union of skew lines. In the first case, the proof follows as before. If $B=L_{1}\cup L_{2}$, then $\omega_{B}\simeq\omega_{L_{1}}\oplus\omega_{L_{2}}$ and the claim follows from the fact $\omega_{\mathbb{P}^{1}}\simeq\mathcal{O}_{\mathbb{P}^{1}}(-2)$ in each component. ∎
Remark 26.
The results above can be used to restrict the possible values of numerical invariants. For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results to show the only possible third Chern class is $c_{3}({\mathcal{T}}_{\sigma})=8$.
Indeed, from [undefu, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 6 we obtain $c_{3}({\mathcal{T}}_{\sigma})=2(p_{a}(B)+5)=8-2a$, where $p_{a}(B)=-1-a$. From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(4-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we obtain the Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(4-a))=10-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=10-2a,$
and additivity of Euler characteristic gives a contradiction whenever $a\neq 0$, so $c_{3}({\mathcal{T}}_{\sigma})=8$ if this occurs.
In the next section, we show the above cannot happen either (see 43).
As another application, we are able to show that sequences $\sigma$ with mixed degrees $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$ cannot be unstable.
Proposition 27.
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$ and $m(\sigma)=3$. If ${\mathcal{T}}_{\sigma}$ is unstable, then the possible discrete invariants are $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=6$.
Proof.
Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$. From Theorem A Theorem A, $\sigma$ is incompressible. Using Remark 6 we have for $\deg(B)=2$, $p_{a}(B)=-1-a$ yields $c_{3}({\mathcal{T}}_{\sigma})=6-2a\geq 0$, so $a\in\{-1,0,1,2,3\}$, but from
From the short exact sequence
$0\rightarrow\mathcal{O}_{W}\rightarrow\mathcal{O}_{B}(3-a)\rightarrow{\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}})\rightarrow 0,$
we get Euler characteristics $\mathcal{X}(\mathcal{O}_{B}(3-a))=8-a$ and
$\mathcal{X}(\mathcal{O}_{W})+\mathcal{X}({\mathcal{E}}{\it xt}^{3}(\mathcal{U},\mathcal{O}_{\mathbb{P}^{3}}))=2+c_{3}({\mathcal{T}}_{\sigma})=8-2a,$
which can only coincide for $a=0$. ∎
We note that we have no explicit construction for the unstable case above to occur as an example.
4 Pencils of cubics and degree 6 curves inside quadric surfaces
In this final section, we show some classification results for pencils of cubics and sequences $\sigma=(f,g)$ with $d_{f}=1$, $d_{g}=2$, which correspond to degree $6$ curves inside quadric surfaces.
The results are derived from the previous sections, Section 2.1 and Section 3, and also from general results for reflexive sheaves of rank two on $\mathbb{P}^{3}$, found in the classical works [undefn], [undefw], [undefo], [undefb] and [undefp].
Two major results which are used are the following ones, both assuming that $\kappa$ is an algebraically closed field of characteristic zero.
Theorem 28 ([undefp], Theorem 1.1).
Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$. Assume $c_{1}\geq-3$ and $h^{0}(\mathcal{E})=0$. Define the integers:
$A$ $\doteq\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil,\text{resp. ditto +1 if }c_{1}=1,3;$
$B$ $\doteq\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil,\text{resp. ditto +1 if }c_{1}=2,4;\text{ditto -1 if }c_{1}=-3.$
Then $c_{2}\geq A$. Furthermore:
  • (a)
    If $A\leq c_{2}\leq B$, then
    $c_{3}\leq(c_{1}+4)c_{2}-2\binom{c_{1}+3}{3}-2.$
  • (b)
    If $c_{2}>B$, then
    $c_{3}\leq c_{2}^{2}-c_{2}(2B-c_{1}-5)+B^{2}-B-2\binom{c_{1}+3}{3}-2.$
With this theorem, we obtain an upper bound for $m(\sigma)$ lower than $m_{0}$ when $\sigma$ is incompressible, so that $h^{0}({\mathcal{T}}_{\sigma})=0$, in the case $d_{f}=1,d_{g}=2$. In the case of pencils of cubics, however, we apply this to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, so we must assume $\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ and treat the case $\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ separately.
Another important result is the following, which enables us to obtain a lower bound for the existence of a syzygy in some cases:
Theorem 29 ([undefo], Theorem 0.1).
Let $\mathcal{E}$ be a rank two reflexive sheaf on $\mathbb{P}^{3}$ with $c_{1}=0$ of $c_{1}=-1$ and with $c_{2}\geq 0$. Let $t\in\mathbb{Z}$ such that either
  • (a)
    $c_{1}=0$ and $t>\sqrt{3c_{2}+1}-2$, or
  • (b)
    $c_{1}=-1$ and $t>\sqrt{3c_{2}+\frac{1}{4}}-\frac{3}{2}$.
Then $H^{0}(\mathcal{E}(t))\neq 0$.
4.1 Pencils of cubics
Now, assume $d_{f}=d_{g}=2$. The Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ is given by the formula
$\operatorname{Bour}(\sigma)=e(e-4)+12-m(\sigma).$
When $e=1$, the formula above implies $m(\sigma)\leq 9$, from the inequality $\operatorname{Bour}(\sigma)\geq 0$.
Applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, assuming $e>1$, we get
$c_{1}(\mathcal{E})=c_{1}({\mathcal{T}}_{\sigma}(1))=-d+2=-2$
and the quantities of the Theorem become
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(4-4+3)\right\rceil=1$
$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil=\left\lceil\frac{1}{3}(4-12+8)\right\rceil=0.$
Moreover, another important invariant is:
$c_{2}(\mathcal{E})$ $=c_{2}({\mathcal{T}}_{\sigma})-d+1$
$=m_{0}-m(\sigma)-d+1$
$=9-m(\sigma).$
Direct application of the result shows the next pair of propositions:
Proposition 30.
If $\sigma=(f,g)$ is a normal pencil of cubics with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then $m(\sigma)\leq 9$.
Proposition 31.
If $\sigma=(f,g)$ is a normal pencil of cubics and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$, then
  • (a)
    if $8\leq m(\sigma)\leq 9$, then $c_{3}\leq 16-2m(\sigma)$. In particular, when $m(\sigma)=8$, $\sigma$ is locally free.
  • (b)
    if $0\leq m(\sigma)<7$, then
    $c_{3}\leq m(\sigma)^{2}-17m(\sigma)+72.$
Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(2)$, a sheaf with $c_{1}=0$ and $c_{2}(\mathcal{E})=9-m(\sigma)\geq 0$ from the previous computations, assuming $e>1$. Then, the Theorem says that, for
$t$ $>\sqrt{3c_{2}+1}-2=\sqrt{27-3m(\sigma)+1}-2$
we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+2))\neq 0$.
Direct application of this bound for high values of $m(\sigma)=6,7,8$ gives minimum $t=2,1,1$ satisfying the inequality above, so we obtain
Proposition 32.
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 4$, and we have the following table of sharper bounds for each possibility of $m=m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
7 $e\leq 3$
8 $e\leq 3$
The following result gives a picture of the generic case of a pencil of cubics.
Proposition 33.
Let $\sigma=(f,g)$ be a general pencil of cubics in $\mathbb{P}^{3}$. Then $m(\sigma)=0$, the number of singular members is $32$, and all singular members have one singular point, in particular $m(\sigma)=0$. Moreover $c_{3}({\mathcal{T}}_{\sigma})=32$ (see Example 10), and the minimal free resolution for the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ is of the form
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to{\mathcal{T}}_{\sigma}\to 0,$
given by the Buchsbaum–Rim complex.
Proof.
This follows from intersection theory for the bundle of principal parts (see [undefi], Proposition $7.1$ and Proposition $7.4$). ∎
Rewriting Theorem A Theorem A, $(d)$ for pencils of cubics, we obtain:
Proposition 34.
Let $\sigma$ be a normal pencil of cubics. Then $\sigma$ is compressible if and only if $m(\sigma)=12$ (see Example 35), and in this case ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)$.
Example 35 (Free, compressible pencil of cubics).
Consider the sequence $\sigma=(x_{0}^{3}+x_{1}^{3}+x_{0}x_{1}x_{3},x_{0}x_{1}x_{3})$. This sequence is independent of the variable $x_{2}$, with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{3}&3x_{1}^{2}+x_{0}x_{3}&0&x_{0}x_{1}\\ x_{1}x_{3}&x_{0}x_{3}&0&x_{0}x_{1}\end{pmatrix}$
so that there are two linearly independent syzygies, one of degree zero and one of degree four, in the following matrix:
$\nu=\begin{pmatrix}0&-x_{0}x_{1}^{3}\\ 0&x_{0}^{3}x_{1}\\ 1&0\\ 0&-x_{0}^{3}x_{3}+x_{1}^{3}x_{3}\end{pmatrix}$
Here, $m(\sigma)=12$, with the irreducible components of $(\Xi_{\sigma})_{1}$ being three lines $V(x_{0},x_{3}),V(x_{1},x_{3})$ and $V(x_{0},x_{1})$, where the last one has multiplicity $10$ and the other two are simple.
Using Theorem B Theorem B, we obtain the following bounds for $\mu$-semistability of ${\mathcal{T}}_{\sigma}$ in terms of $m(\sigma)$:
Proposition 36.
Let $\sigma$ be a normal pencil of cubics. Then:
  • (a)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable;
  • (b)
    If $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since $\mu({\mathcal{T}}_{\sigma})=-2$, we show for $m(\sigma)\leq 6$ that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$. Since $m(\sigma)\leq 6$, we conclude that $\sigma$ is neither compressible nor free, from the previous results 34, 37 and 39. Let us suppose that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. Since $\sigma$ is not free, we can apply the result Theorem B Theorem B, $(a)$, and conclude $\operatorname{Bour}(\sigma)\leq 2$. On the other hand, we have
$\operatorname{Bour}(\sigma)$ $=1-d+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$
$=1-4+4+4+4-m(\sigma)$
$=9-m(\sigma)>2,$
since $m(\sigma)\leq 6$. This contradicts the bound $\operatorname{Bour}(\sigma)\leq 2$ established earlier.
Moreover, for $(b)$, if we assume $m(\sigma)\leq 2$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, then
$\operatorname{Bour}(\sigma)$ $=2-2d+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma)$
$=4-8+12-m(\sigma)$
$=8-m(\sigma)>5$
if $m(\sigma)\leq 2$, we got a contradiction with the bound $\operatorname{Bour}(\sigma)\leq 5$ established earlier in Theorem B Theorem B, $(b)$. ∎
Proposition 37.
Let $\sigma=(f,g)$ be an incompressible normal pencil of cubics. Then $m(\sigma)\leq 9$, and $m(\sigma)=9$ if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$ (see Example 38).
Proof.
The bound $m(\sigma)\leq 8$ is obtained for $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})>1$ in 30 for pencils of cubics, and $m(\sigma)\leq 9$ holds for $e\geq 1$, thus $m(\sigma)=9$ only if $e=1$. From the formula of the Bourbaki degree, we obtain $\operatorname{Bour}(\sigma)=0$ in this case, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$. On the other hand, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, then $e=1$, and the equation $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=9$. ∎
Example 38 (Free, unstable and incompressible pencil of cubics).
Consider the sequence $\sigma=(x_{1}(x_{2}^{2}-x_{1}^{2}),x_{3}x_{2}(x_{0}-x_{1}))$. Then the matrix $\nabla\sigma$ is given by:
$\nabla\sigma=\begin{pmatrix}0&-3x_{1}^{2}+x_{2}^{2}&2x_{1}x_{2}&0\\ x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\end{pmatrix}$
and it admits two linearly independent syzygies, one of degree one and one of degree $3$:
$\nu\doteq\begin{pmatrix}x_{0}-x_{1}&2x_{1}x_{2}^{2}\\ 0&2x_{1}x_{2}^{2}\\ 0&3x_{1}^{2}x_{2}-x_{2}^{3}\\ -x_{3}&-3x_{1}^{2}x_{3}+x_{2}^{2}x_{3}\end{pmatrix}$
Thus, we conclude ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$, and in particular $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. In this case, $m(\sigma)=9$.
Proposition 39.
Let $\sigma=(f,g)$ be a normal pencil of cubics. Then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ if and only if $m(\sigma)=8$ (see Example 40).
Proof.
If $m(\sigma)=8$, by 32, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$. If $e=2$, we note that $c_{2}({\mathcal{T}}_{\sigma}(2))=0$ and ${\mathcal{T}}_{\sigma}(2)$ is strictly $\mu$-semistable with $c_{1}=0$, thus it follows from [undefb, Lemma 2.0, (a)] that ${\mathcal{T}}_{\sigma}(2)\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}$. If we assume $e\in{1,3}$, we obtain $\operatorname{Bour}(\sigma)=1$ and, using Remark 6 for $B$ being a line (that is, $p_{a}(B)=0$ and $\deg(B)=1$), we obtain:
$e=1$ $\Rightarrow c_{3}=4$
$e=3$ $\Rightarrow c_{3}=0$
On the other hand, 31 gives $c_{3}({\mathcal{T}}_{\sigma})=0$ for $m(\sigma)=8$, so we obtain a contradiction with the case $e=1$. For the case $e=3$, we would have a nearly-free locally free pencil, which cannot happen by 19.
Conversely, if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $\operatorname{Bour}(\sigma)=0$ implies $m(\sigma)=8$. ∎
Example 40 (Free, incompressible and $\mu$-semistable pencil of cubics ($m(\sigma)=8$)).
Consider the sequence $\sigma=(x_{0}^{2}x_{1}+x_{3}^{3},x_{0}^{3}+x_{0}x_{2}x_{3}+x_{3}^{3})$. The Jacobian matrix $\nabla\sigma$ is given by:
$\nabla\sigma=\begin{pmatrix}2x_{0}x_{1}&x_{0}^{2}&0&3x_{3}^{2}\\ 3x_{0}^{2}+x_{2}x_{3}&0&x_{0}x_{3}&x_{0}x_{2}+3x_{2}^{2}\end{pmatrix}$
and it admits two linearly independent syzygies of degree $2$:
$\nu\doteq\begin{pmatrix}-x_{0}x_{3}&-x_{0}x_{2}\\ 2x_{1}x_{3}&2x_{1}x_{2}-9x_{3}^{2}\\ 3x_{0}^{2}+x_{2}x_{3}&x_{2}^{2}-9x_{0}x_{3}\\ 0&3x_{0}^{2}\end{pmatrix}$
Thus, we conclude ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$.
Proposition 41.
Let $\sigma=(f,g)$ be a nearly free pencil of cubics. Then, the only possible discrete invariants are $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $m(\sigma)=7$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42).
Proof.
Using Remark 6 for $p_{a}(B)=0$ and $\deg(B)=1$, we obtain
$c_{3}({\mathcal{T}}_{\sigma})$ $=2p_{a}(B)-2+\deg(B)(4+d-2e)$
$=-2+8-2e=2(3-e).$
On the other hand, if $m(\sigma)=7$, then $e=2$ and the formula for $c_{3}$ in Remark 6 yields $c_{3}({\mathcal{T}}_{\sigma})=2$. As we have observed in Example 40 and before, we must have $m(\sigma)\leq 7$ to be able to obtain $\operatorname{Bour}(\sigma)=1$. On the other hand, assuming $m(\sigma)\leq 6$, we obtain
$\operatorname{Bour}(\sigma)=12-m(\sigma)+e(e-4)\geq 6+e(e-4)\geq 2$
hence $\operatorname{Bour}(\sigma)\neq 1$. ∎
Example 42 (Nearly free pencil of Cubics).
We consider the following sequence of cubics:
$\sigma=(x_{0}^{2}(x_{1}-x_{2})+x_{2}^{2}(x_{1}-x_{0}+x_{3}),-x_{1}x_{2}x_{3}+x_{2}^{2}x_{3})$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&x_{0}^{2}+x_{2}^{2}&-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})&x_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$.
There are four irreducible components in $(\Xi_{\sigma})_{1}$, all lines, corresponding to the four prime ideals below:
$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
$\mathfrak{p}_{2}$ $=(x_{1}-x_{2},x_{3})$
$\mathfrak{p}_{3}$ $=(x_{1},x_{2})$
$\mathfrak{p}_{4}$ $=(x_{0},x_{2})$
For the prime $\mathfrak{p}_{1}$, we may consider the element $u=(2x_{0}(x_{1}-x_{2})-x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{1}}$ and the elementary operation $R_{1}^{\prime}=uR_{1}$, to obtain the matrix
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}1&u(x_{0}^{2}+x_{2}^{2})&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&-x_{2}x_{3}&x_{3}(2x_{2}-x_{1})&x_{2}(x_{2}-x_{1})\end{pmatrix}$,
which sends the first basis vector to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$ and therefore we may compute the cokernel by considering the second line of the remaining matrix, namely
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(-x_{2}x_{3},x_{3}(2x_{2}-x_{1}),x_{2}(x_{2}-x_{1}))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(-x_{2}x_{3},x_{3},x_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}}\simeq\kappa,$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
For the prime $\mathfrak{p}_{2}=(x_{1}-x_{2},x_{3})$, the same element $u\in R_{\mathfrak{p}_{2}}$ is invertible, so that we compute the cokernel analogously to obtain:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(-x_{2}x_{3},x_{3}(2x_{2}-x_{1}),x_{2}(x_{2}-x_{1}))}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{1}-x_{2})}\simeq\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
For the prime $\mathfrak{p}_{3}=(x_{1},x_{2})$, we take $u=(x_{0}^{2}+x_{2}^{2})^{-1}\in R_{\mathfrak{p}_{3}}^{\times}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}+u[x_{2}x_{3}]R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to obtain
$(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}u(2x_{0}(x_{1}-x_{2})-x_{2}^{2})&1&u(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&ux_{2}^{2}\\ 0&0&x_{3}(2x_{2}-x_{1})-ux_{2}x_{3}(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))&x_{2}(x_{2}-x_{1})-ux_{2}^{3}x_{3}\end{pmatrix}$,
which sends $e_{2}\in R_{\mathfrak{p}_{3}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{3}}^{2}$. Thus, we may compute the cokernel by considering the remaining columns of the second row:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{(f_{1},f_{2})}$
where
$f_{1}$ $=x_{3}(2x_{2}-x_{1})+ux_{2}x_{3}(-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}))$
$f_{2}$ $=x_{2}(x_{2}-x_{1})+ux_{2}^{3}x_{3}.$
Considering $f_{1}\in I$, we obtain
$2x_{2}(1+ux_{2}(x_{1}-x_{0}+x_{3})-ux_{0}^{2})\equiv x_{1}\mod I.$
Substituting this into $f_{2}\in I$, we get:
$0\equiv x_{2}^{2}\left(1-(2(1+ux_{2}(x_{1}-x_{0}+x_{3}))-ux_{0}^{2})+ux_{2}x_{3}\right)\mod I,$
and the element multiplying $x_{2}^{2}$ is an unit in $R_{\mathfrak{p}_{3}}$, we conclude that $x_{2}^{2}\in I$. Hence $x_{1}^{2}\in I$ by $f_{1}\in I$, and we may rewrite $f_{1}$ and $f_{2}$ mod $I$ as:
$f_{1}$ $=x_{2}(2-ux_{0}^{2})-x_{1}\mod I$
$f_{2}$ $=-x_{1}x_{2}\mod I,$
and thus the ideal $I=(x_{1}^{2},x_{2}^{2},x_{1}x_{2},x_{2}v-x_{1})$ where $v\in R_{\mathfrak{p}_{3}}$ is a unit. Hence, one may write
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa\langle 1,x_{1},x_{2}\rangle,$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=3$.
For the prime $\mathfrak{p}_{4}=(x_{0},x_{2})$, we take $u=(x_{3}(2x_{2}-x_{1}))^{-1}\in R_{\mathfrak{p}_{4}}$ and the elementary operations $R_{1}^{\prime}=R_{1}-u[-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3})]R_{2}\doteq R_{1}-uQR_{2}$ and $R_{2}^{\prime}=uR_{2}$, to obtain:
$(\nabla\sigma)_{\mathfrak{p}_{4}}\sim\begin{pmatrix}2x_{0}(x_{1}-x_{2})-x_{2}^{2}&(x_{0}^{2}+x_{2}^{2})+uQx_{2}x_{3}&0&x_{2}^{2}-uQx_{2}(x_{2}-x_{1})\\ 0&-ux_{2}x_{3}&1&ux_{2}(x_{2}-x_{1})\end{pmatrix}$,
sending $e_{3}\in R_{\mathfrak{p}_{4}}^{4}$ to $(0,1)\in R_{\mathfrak{p}_{4}}^{2}$. Hence, we compute the cokernel as:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}\simeq\frac{R_{\mathfrak{p}_{4}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{4}}}{I}$
with
$f_{1}$ $=2x_{0}x_{1}-x_{2}(2x_{0}+x_{2})=2x_{0}(x_{1}-x_{2})-x_{2}^{2}$
$f_{2}$ $=(x_{0}^{2}+x_{2}^{2})+uQx_{2}x_{3}$
$f_{3}$ $=x_{2}^{2}-uQx_{2}(x_{2}-x_{1})$
$u$ $=(x_{3}(2x_{2}-x_{1}))^{-1}$
$Q$ $=-x_{0}^{2}+2x_{2}(x_{1}-x_{0}+x_{3}).$
Thus,
$f_{2}$ $=x_{0}^{2}+x_{2}^{2}-ux_{3}x_{0}^{2}x_{2}+2u(x_{1}-x_{0}+x_{3})x_{3}x_{2}^{2}$
$=x_{0}^{2}(1-ux_{3}x_{2})+x_{2}^{2}(1+2u(x_{1}-x_{0}+x_{3})x_{3})$
$f_{3}$ $=x_{2}^{2}+u(x_{2}-x_{1})x_{0}^{2}x_{2}-2u(x_{2}-x_{1})(x_{1}-x_{0}+x_{3})x_{2}^{2},$
where the elements $(x_{2}-x_{1}),(x_{1}-x_{0}+x_{3}),x_{3}\in R_{\mathfrak{p}_{4}}^{\times}$. Since $f_{1}\in I$, we have
$x_{2}^{2}\equiv 2x_{0}(x_{1}-x_{2})\mod I,$
and thus
$f_{2}\equiv x_{0}(x_{0}(1-ux_{3}x_{2})+2(x_{1}-x_{2})+2u(x_{1}-x_{2})(x_{1}-x_{0}+x_{3})x_{3})\mod I,$
but since the element inside the parenthesis is invertible in $R_{\mathfrak{p}_{4}}$, we conclude that $x_{0}\in I$. From $f_{1}\in I$, we obtain that $x_{2}^{2}\in I$. Taking this into account, we conclude that
$f_{1}$ $\equiv-x_{2}^{2}\mod I$
$f_{2}$ $\equiv x_{2}^{2}\mod I$
$f_{3}$ $\equiv x_{2}^{2}(1-2ux_{2}(x_{1}+x_{3})+2ux_{1}(x_{1}+x_{3}))\equiv x_{2}^{2}\mod I,$
and therefore
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}\simeq\frac{R_{\mathfrak{p}_{4}}}{(x_{0},x_{2}^{2})}\simeq\kappa\langle 1,x_{2}\rangle,$
with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=2$. Using the associativity formula to compute $m(\sigma)$, we obtain:
$m(\sigma)=1\cdot 1+1\cdot 1+3\cdot 1+2\cdot 1=7.$
On the other hand, $e(\sigma)=2$, so using the Bourbaki formula we obtain $\operatorname{Bour}(\sigma)=1$, and conclude also the minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the form
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0.$
Proposition 43.
Let $\sigma=(f,g)$ be a normal, $\mu$-semistable pencil of cubics such that $m(\sigma)=7$. Then, the only possible cases are:
  • $e=2$ and $\sigma$ is a nearly free sequence with $c_{3}({\mathcal{T}}_{\sigma})=2$ (see 41);
  • $e=3$ and $\sigma$ is locally free, $\operatorname{Bour}(\sigma)=2$ and $B$ is a pair of skew lines (or their degeneration) (see Example 23).
Proof.
From 32, we know that $e\leq 3$. For $\mu$-semistable sheaves, $e\geq 2$ and, since $c_{2}({\mathcal{T}}_{\sigma}(2))=1$ and $c_{1}({\mathcal{T}}_{\sigma}(2))=0$, from [undefb, Lemma 2.1], we conclude that the only possible cases are $c_{3}=2$ (strictly semistable case) or $c_{3}=0$ (stable case). For the strictly semistable case $e=2$ we obtain $\operatorname{Bour}(\sigma)=1$, so we get the first case.
For the second case, we must have a stable bundle ${\mathcal{T}}_{\sigma}(2)$ with Chern classes $(0,1,0)$, which are precisely null correlation bundles described in [undefx] fitting in a sequence of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}(2)\rightarrow\mathcal{I}_{B}(1)\rightarrow 0$
where $B$ is a pair of skew lines or their degeneration, thus $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=3$. ∎
Theorem C.
Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 12$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=12,9$ or $8$, corresponding to $e$ being $0,1$ or $2$, respectively;
  • (c)
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42), which is strictly $\mu$-semistable;
  • (d)
    If $m(\sigma)\leq 6$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable, and if $m(\sigma)\leq 2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Item $(a)$ is 34 and item $(b)$ with 37 and Example 40. Item $(c)$ follows from 43 and item $(d)$ is 36. ∎
To finish this study, we consider a strictly $\mu$-semistable pencil of cubics with $m(\sigma)=4$.
Example 44 ($m(\sigma)=4$, $e=2$, $\operatorname{Bour}(\sigma)=4$, $3$-syzygy).
Considering the following pencil of cubics $(d_{f}=d_{g}=2)$, where the first one is smooth:
$\sigma=(x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3},x_{0}^{3}+x_{1}^{3}+x_{2}x_{3}^{2}),$
with Jacobian matrix
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}&3x_{1}^{2}&3x_{2}^{2}&3x_{3}^{2}\\ 3x_{0}^{2}&3x_{1}^{2}&x_{3}^{2}&2x_{2}x_{3}\end{pmatrix}$.
The scheme structures $\operatorname{supp}(\mathcal{Q}_{\sigma})=\Xi_{\sigma}$ coincide, with support at a line $V(x_{2},x_{3})$. Over the prime ideal $\mathfrak{p}=(x_{2},x_{3})$, we perform some elementary operations for the matrix $\nabla\sigma$, first $R_{1}^{\prime}=R_{1}-R_{2}$, and then $R_{2}^{\prime}=u/3\cdot R_{2}$, where $u=(x_{1}^{2})^{-1}\in R_{\mathfrak{p}}$, to obtain the form:
$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}0&0&3x_{2}^{2}-x_{3}^{2}&3x_{3}^{2}-2x_{2}x_{3}\\ ux_{0}^{2}&1&(u/3)x_{3}^{2}&(2/3)ux_{2}x_{3}\end{pmatrix}$,
so it sends the second basis vector to $(0,1)\in R_{\mathfrak{p}}^{2}$, and then we may compute the cokernel using the first line and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ $\simeq\frac{R_{P}}{(3x_{2}^{2}-x_{3}^{2},3x_{3}^{2}-2x_{2}x_{3})}$
$\simeq\kappa\langle 1,x_{2},x_{3},x_{2}^{2}\rangle,$
and thus we conclude $\operatorname{{length}}(\mathcal{Q})_{\mathfrak{p}}=4$. Since $\deg(R/\mathfrak{p})=1$, using the associativity formula we obtain $m(\sigma)=4$. A first syzygy is $\nu=(-x_{1}^{2},x_{0}^{2},0,0)^{T}$, so we obtain $e=2$, and thus $\operatorname{Bour}(\sigma)=4$.
The sheaf ${\mathcal{T}}_{\sigma}$ admits a minimal free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
obtained using Macaulay2, so this is an example of a $3$-syzygy sequence, with $c_{3}({\mathcal{T}}_{\sigma})=16$.
To finish this subsection, we explore in an example that the Bourbaki degree of a pencil of cubics does not only depend on the singularity type of each cubic, but rather how these align when we consider the associated pencil.
Example 45.
Consider the cubic polynomials:
$f$ $=x_{3}(x_{1}^{2}-x_{0}x_{2})+x_{1}^{2}(x_{0}-x_{1})$
$g$ $=x_{3}(x_{1}^{2}-x_{0}x_{2})+x_{1}^{3}$
$g^{\prime}$ $=x_{0}(x_{3}^{2}-x_{1}x_{2})+x_{3}^{3}.$
Here, $V(f)$ is a cubic surface with singularity type $4A_{1}$, with primary ideals $(x_{0},x_{1},x_{2})$, $(x_{0},x_{1}^{2},x_{3})$ and $(x_{1},x_{2},x_{3})$. On the other hand, the surfaces $V(g)$ and $V(g^{\prime})$ are both of type $A_{1}2A_{2}$ and are isomorphic, via the change of coordinates
$x_{0}$ $\mapsto x_{1}$
$x_{1}$ $\mapsto x_{3}$
$x_{3}$ $\mapsto x_{0}$
and fixing $x_{3}$. The sequence $\sigma=(f,g)$ is non-normal with $V(x_{1})$ in the divisorial component of the Jacobian scheme. On the other hand, $\sigma^{\prime}=(f,g^{\prime})$ is a normal, generic pencil with $m(\sigma^{\prime})=0$.
4.2 Degree 6 curves inside quadric surfaces
Let us focus on the case of normal sequences $\sigma=(f,g)$ with $d_{f}=1,d_{g}=2$. Here, the Bourbaki degree of a sequence $\sigma$ in terms of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ and $m(\sigma)$ is given by:
$\operatorname{Bour}(\sigma)=e(e-3)+7-m(\sigma).$
We start by applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}$, assuming $\sigma$ is incompressible. The quantities of the theorem become
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(9-6+3)\right\rceil=0$
$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil-1=\left\lceil\frac{1}{3}(9-9+8)\right\rceil-1=1.$
Moreover, another important invariant is: $c_{2}(\mathcal{E})=7-m(\sigma)$. Direct application of the result shows the next pair of propositions:
Proposition 46.
If $\sigma=(f,g)$ is an incompressible normal sequence with $d_{f}=1,d_{g}=2$, then $m(\sigma)\leq 5$.
Proof.
Let us assume that $e\in[1,d-2]$ and $d\geq 3$. Then, from the Bourbaki formula and the fact that $\operatorname{Bour}(\sigma)\geq 0$, we obtain
$m(\sigma)\leq m_{0}+e(e-d)\leq m_{0}+e-d\leq m_{0}-2,$
which gives the claim. This covers the case $e=1$. If we now assume $e>1$ and apply Theorem 28 to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a reflexive sheaf with $c_{1}=-1$, we obtain the constant is
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(1-2+3)\right\rceil=0.$
Hence, $5-m(\sigma)=c_{2}({\mathcal{T}}_{\sigma}(1))=c_{2}(\mathcal{E})\geq A=0$ and thus $m(\sigma)\leq 5$. ∎
Proposition 47.
Let $\sigma=(f,g)$ be an incompressible normal sequence with $d_{f}=1$ and $d_{g}=2$. Then, the following hold:
  • (a)
    If $6\leq m(\sigma)\leq 7$, then $c_{3}\leq 5-m(\sigma)$.
  • (b)
    If $0\leq m(\sigma)<6$, then $c_{3}\leq m(\sigma)^{2}-14m(\sigma)+47$.
From Theorem A Theorem A, $(d)$, we obtain:
Proposition 48.
Let $\sigma$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $\sigma$ is compressible if and only if $m(\sigma)=7$ (see Example 49).
Example 49 (Free and compressible sequence with $d_{f}=1,d_{g}=2$).
Considering the sequence
$\sigma=(x_{0}(x_{1}-x_{2}),x_{0}^{3}+x_{1}^{3}+x_{2}^{3}),$
which is independent of the variable $x_{3}$. The matrix
$\begin{pmatrix}0&-x_{0}(x_{1}^{2}+x_{2}^{2})\\ 0&x_{0}^{3}+x_{1}x_{2}^{2}-x_{2}^{3}\\ 0&x_{0}^{3}-x_{1}^{3}+x_{1}^{2}x_{2}\\ 1&0\end{pmatrix}$
gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$.
Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a sheaf with $c_{1}=-1$ and $c_{2}(\mathcal{E})=5-m(\sigma)\geq 0$. Then, the theorem says that, for
$t$ $>\sqrt{3c_{2}+\frac{1}{4}}-\frac{3}{2}=\sqrt{15-3m(\sigma)+\frac{1}{4}}-\frac{3}{2}$
we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+1))\neq 0$.
Direct application of this bound for high values of $m(\sigma)=4,5$ we get $t=0,t=1$, respectively. This shows the following proposition.
Proposition 50.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1$, $d_{g}=2$. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\leq 3$, and we have the following table of bounds for the following possibilities of $m(\sigma)$:
$m(\sigma)$ $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$
4 $e\leq 2$
5 $e\leq 1$
As an easy consequence of Remark 6, we obtain:
Proposition 51.
Let $\sigma=(f,g)$ be a normal sequence which is not a pencil, (i.e., with $d_{f}\neq d_{g}$) such that $d$ is odd. If $\operatorname{Bour}(\sigma)$ is odd, then $\sigma$ is not locally free.
Proof.
Assuming ${\mathcal{T}}_{\sigma}$ is locally free, we get $c_{3}({\mathcal{T}}_{\sigma})=0$, and therefore
$2p_{a}(B)$ $=2-\operatorname{Bour}(\sigma)(4+d-2e)$
$p_{a}(B)$ $=1-\operatorname{Bour}(\sigma)(2+\frac{d}{2}-e)=1-(2-e)\operatorname{Bour}(\sigma)+\frac{d\operatorname{Bour}(\sigma)}{2},$
which is not an integer if $\operatorname{Bour}(\sigma)$ is odd, so we get a contradiction. ∎
Proposition 52.
Let $\sigma$ be an incompressible sequence with $d_{f}=1,d_{g}=2$ such that $m(\sigma)$ is even, that is, $m(\sigma)\in\{0,2,4\}$. Then $\sigma$ is not locally free.
Proof.
To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 51.
From 3, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
Assuming $e=2k$ is even, we obtain
$e^{2}-3e=4k^{2}-6k=2(2k^{2}-3k),$
an even number. On the other hand, when $e=2k+1$ is odd, then
$e^{2}-3e=4k^{2}+4k+1-6k-3=4k^{2}-2k-2=2(2k^{2}-k-1),$
which is also even. ∎
Proposition 53.
Let $\sigma=(f,g)$ be an incompressible normal sequence with degrees $d_{f}=1$, $d_{g}=2$. Then ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$ if and only if $m(\sigma)=5$ (see Example 54).
Proof.
Assuming $\sigma$ is non-compressible, by 50, we obtain that $e=1$. But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free. On the other hand, if we assume ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=1$ and $\operatorname{Bour}(\sigma)=0$ give $m(\sigma)=5$. ∎
Example 54 (Free and incompressible sequence, $m(\sigma)=5$).
Considering the sequence
$\sigma=(x_{0}x_{1},x_{3}x_{2}(x_{0}-x_{1})),$
of arrangements of hyperplanes, with Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&0&0\\ x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\end{pmatrix}$
The matrix
$\begin{pmatrix}0&x_{0}(x_{0}-x_{1})\\ 0&-x_{1}(x_{0}-x_{1})\\ x_{2}&0\\ -x_{3}&-x_{3}(x_{0}+x_{1})\end{pmatrix}$
gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ and $m(\sigma)=5$.
Proposition 55.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=4$, then $\sigma$ is nearly free, and we have two possible cases:
  • (a)
    ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=1$ (see Example 56);
  • (b)
    ${\mathcal{T}}_{\sigma}$ is unstable with $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 14);
Furthermore, these are the only two possibilities of numerical invariants for nearly free sequences with $d_{f}=1,d_{g}=2$.
Proof.
Using 50, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$. Using Remark 6 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples. To conclude the last claim, we note that if $m(\sigma)\leq 3$, then
$\operatorname{Bour}(\sigma)=7-m(\sigma)+e(e-3)\geq 4+e(e-3)\geq 2$
for $e\in\{1,2,3\}$. ∎
Example 56 (Nearly free sequence with $d_{f}=1,d_{g}=2$ and $e=2$).
We consider the following normal sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(x_{0}x_{1}-x_{2}x_{3},x_{1}x_{3}(x_{0}-x_{2})),$
with Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}x_{1}&x_{0}&-x_{3}&-x_{2}\\ x_{1}x_{3}&x_{3}(x_{0}-x_{2})&-x_{1}x_{3}&x_{1}(x_{0}-x_{2})\end{pmatrix}$.
The irreducible components of $\Xi_{\sigma}$ consist of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$ outside the three lines, so $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$. We may compute $m(\sigma)=4$, where the first line above occurs with length two and the other two occur with length one.
For $\mathfrak{p}_{1}=(x_{1},x_{3})$, we may consider $u=(x_{0})^{-1}$ and perform two elementary operations, namely $R_{2}^{\prime}=R_{2}-ux_{3}(x_{0}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite the matrix:
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{1}&1&-ux_{3}&-ux_{2}\\ x_{1}x_{3}(1-u(x_{0}-x_{2}))&0&-x_{3}(x_{1}-x_{3}(x_{0}-x_{2})u)&(x_{0}-x_{2})(x_{1}-x_{2}x_{3}u)\end{pmatrix}$.
Thus, it sends the second basis vector $e_{2}\in R_{\mathfrak{p}_{1}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$, and we may compute the cokernel from the second row of the remaining matrix:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{1}x_{3}(1-u(x_{0}-x_{2})),-x_{3}(x_{1}-x_{3}(x_{0}-x_{2})u),(x_{0}-x_{2})(x_{1}-x_{2}x_{3}u))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{1}x_{3},-x_{3}^{2},x_{1}-x_{2}x_{3}u)}$
$\simeq\kappa\langle 1,x_{3}\rangle,$
simplifying the invertible elements $(1-u(x_{0}-x_{2})),(x_{0}-x_{2})\notin\mathfrak{p}_{1}$, hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=2$.
For $\mathfrak{p}_{2}=(x_{1},x_{0}-x_{2})$, consider $u=(-x_{3})^{-1}\in R_{\mathfrak{p}_{2}}$ and perform two elementary operations $R_{2}^{\prime}=R_{2}+ux_{1}x_{3}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite $\nabla\sigma$ as
$(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}ux_{1}&ux_{0}&1&-ux_{2}\\ x_{1}x_{3}(1+ux_{1})&x_{3}((x_{0}-x_{2})+ux_{0}x_{1})&0&x_{1}(x_{0}-x_{2})-ux_{1}x_{2}x_{3}\end{pmatrix}$.
Since it sends $e_{3}$ to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, we may compute the cokernel using the second row and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I},$
where
$f_{1}$ $=x_{1}x_{3}(1+ux_{1})$
$f_{2}$ $=x_{3}((x_{0}-x_{2})+ux_{0}x_{1})$
$f_{3}$ $=x_{1}((x_{0}-x_{2})-ux_{2}x_{3}).$
Since $f_{1}\in I$ and $x_{3}(1+ux_{1})\in R_{\mathfrak{p}_{2}}^{\times}$, we obtain $x_{1}\in I$. From this and $f_{2}\in I$, we obtain:
$x_{3}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
and thus $I=(x_{0}-x_{2},x_{1})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
For $\mathfrak{p}_{3}=(x_{3},x_{0}-x_{2})$, take $u=(x_{1})^{-1}\in R_{\mathfrak{p}_{3}}$ and the elementary operations $R_{2}^{\prime}=R_{2}-ux_{1}x_{3}R_{1}$, $R_{1}^{\prime}=uR_{1}$ to obtain
$(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}1&ux_{0}&-ux_{3}&-ux_{2}\\ 0&x_{3}((x_{0}-x_{2})-ux_{0}x_{1})&x_{1}x_{3}(-1+ux_{3})&x_{1}(x_{0}-x_{2})+ux_{1}x_{2}x_{3}\end{pmatrix}$.
Since it sends $e_{1}$ to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, we may compute the cokernel using the second row and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I},$
where
$f_{1}$ $=x_{3}((x_{0}-x_{2})-ux_{1}x_{0})$
$f_{2}$ $=x_{1}x_{3}(ux_{3}-1)$
$f_{3}$ $=x_{1}((x_{0}-x_{2})+ux_{2}x_{3}).$
Since $f_{1}\in I$ and $(x_{0}-x_{2})-ux_{1}x_{0}\in R_{\mathfrak{p}_{2}}^{\times}$, we obtain $x_{3}\in I$. From this and $f_{3}\in I$, we obtain:
$x_{1}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
and thus $I=(x_{3},x_{0}-x_{2})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. From the associativity formula over $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$, we then obtain
$m(\sigma)=2\cdot 1+1\cdot 1+1\cdot 1=4.$
From the Bourbaki degree formula, once we know that $e=2$, we conclude that $\operatorname{Bour}(\sigma)=1$ and this is a nearly free sequence. The complete minimal free resolution of ${\mathcal{T}}_{\sigma}$ can be obtained computationally
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)\xrightarrow{\gamma}\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 3}\xrightarrow{M}{\mathcal{T}}_{\sigma}\rightarrow 0$
given by matrices
$M=\begin{pmatrix}x_{0}x_{1}+x_{3}(x_{2}-x_{0})&x_{0}^{2}&x_{0}x_{3}\\ -x_{1}^{2}+x_{1}x_{3}&-x_{0}x_{1}&0\\ x_{1}x_{2}&x_{2}^{2}&x_{0}x_{1}+x_{3}(x_{2}-x_{1})\\ 0&-x_{2}x_{3}&x_{1}x_{3}-x_{3}^{2}\end{pmatrix},\gamma=\begin{pmatrix}x_{0}\\ x_{3}-x_{1}\\ -x_{2}\end{pmatrix}$.
which corresponds to the nearly free resolution given in 16. We have $c_{3}({\mathcal{T}}_{\sigma})=1$, corresponding to the point $p$, the unique irreducible component of $\Xi_{\sigma}$ of codimension three.
Proposition 57.
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$, such that $m(\sigma)=3$ and ${\mathcal{T}}_{\sigma}$ is $\mu$-stable. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $\operatorname{Bour}(\sigma)=2$ and we may have $c_{3}=0,2,4$. We have examples for the cases $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
Proof.
We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$. Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$. This follows from [undefw, Proposition 1.1] for $c_{3}=0$, [undefb, Lemma 2.4] for $c_{3}=2$ and [undefn, Lemma 9.6] for $c_{3}=4$. ∎
The unstable case of the above Proposition is numerically described in 27.
Example 58 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=2$).
We consider the following sequence:
$\sigma=(x_{3}(x_{0}-x_{1}),x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{3}^{3})$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$.
Here, $(\Xi_{\sigma})_{1}$ has two irreducible components, the lines $V(x_{3},x_{0})$ and $V(x_{3},x_{0}-x_{1})$. Let the corresponding primes be denoted by $\mathfrak{p}_{1},\mathfrak{p}_{2}$.
Over the prime $\mathfrak{p}_{1}=(x_{0},x_{3})$, let us take $u\doteq(x_{0}-x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ and $Q\doteq x_{0}x_{1}+3x_{3}^{2}$. Consider the elementary operations $R_{2}=R_{2}-u(x_{0}x_{1}+3x_{3}^{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{3}&-ux_{3}&0&1\\ 2x_{0}x_{2}+x_{1}x_{3}-ux_{3}&x_{0}x_{3}+uQx_{3}&x_{0}^{2}&0\end{pmatrix}$,
so the matrix sends $e_{4}\in R_{\mathfrak{p}_{1}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$, so we may compute the cokernel from the second row and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{f_{1},f_{2},f_{3}}\doteq\frac{R_{\mathfrak{p}_{1}}}{I},$
where
$f_{1}$ $=2x_{0}x_{2}+x_{1}x_{3}-u(x_{0}x_{1}+3x_{3}^{2})x_{3}$
$f_{2}$ $=x_{3}(1+u(x_{0}x_{1}+3x_{3}^{2}))$
$f_{3}$ $=x_{0}^{2}.$
Therefore, since $f_{2}\in I$ and $u(x_{0}x_{1}+3x_{3}^{2})\in\mathfrak{p}_{1}$, we conclude $x_{3}\in I$. Rewriting $uf_{1}$ and removing terms with $x_{3}$, we obtain
$uf_{1}=2x_{0}^{3}x_{1}x_{2},$
and since $f_{3}=x_{0}^{2}\in I$, we obtain $I=(x_{0}^{2},x_{3})$, hence
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{0}^{2},x_{3})}\simeq\kappa\langle 1,x_{0}\rangle$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=2$.
Over $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, take $u\doteq(x_{0}^{2})^{-1}$ and consider the elementary operation and $R_{2}^{\prime}=uR_{2}$, to obtain the matrix
$(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$,
which sends $e_{3}\in R_{\mathfrak{p}_{2}}^{4}$ to $(0,1)\in R_{\mathfrak{p}_{2}}^{2}$, so we may compute the cokernel using the remaining columns of the first row, namely:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{0}-x_{1})}=\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
and therefore $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
Thus, from the additivity formula, we obtain $m(\sigma)=3$. Moreover, $e=2$, so that $\operatorname{Bour}(\sigma)=2$.
The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{4}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{4}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
so we get $c_{3}({\mathcal{T}}_{\sigma})=2$ and the Bourbaki scheme is a union of two skew lines.
Example 59 ($m(\sigma)=3$, $e=2$, $\operatorname{Bour}(\sigma)=2$, $c_{3}=4$).
We consider the following sequence:
$\sigma=(x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2},x_{3}(x_{2}-x_{3})(x_{0}-x_{1}))$
with corresponding Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}2x_{0}&2x_{1}&2x_{2}&2x_{3}\\ x_{3}(x_{2}-x_{3})&x_{3}(x_{3}-x_{2})&x_{3}(x_{0}-x_{1})&(x_{0}-x_{1})(x_{2}-2x_{3})\end{pmatrix}$.
Here, the irreducible components of $(\Xi_{\sigma})_{1}$ are three lines, given by the associated primes:
$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
$\mathfrak{p}_{2}$ $=(x_{0}-x_{1},x_{3})$
$\mathfrak{p}_{3}$ $=(x_{2}-x_{3},x_{0}-x_{1}).$
Over $\mathfrak{p}_{1}$, we may choose $u=(2x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}-ux_{3}(x_{3}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite:
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}2ux_{0}&1&2ux_{2}&2ux_{3}\\ x_{3}(x_{2}-x_{3})(1+ux_{0})&0&x_{3}((x_{0}-x_{1})-2ux_{2}(x_{3}-x_{2}))&(x_{0}-x_{1})(x_{2}-2x_{3})-2ux_{3}^{2}(x_{3}-x_{2})\end{pmatrix}$
Thus, it sends $e_{2}$ to $(1,0)$ and compute the cokernel by considering the remaining columns of the second row, namely
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{I}$
with $I=(f_{1},f_{2},f_{3})$ and
$f_{1}$ $=x_{3}(x_{2}-x_{3})$
$f_{2}$ $=x_{3}((x_{0}-x_{1})-2ux_{3}(x_{3}-x_{2}))$
$f_{3}$ $=(x_{0}-x_{1})(x_{2}-2x_{3})-2ux_{3}^{2}(x_{3}-x_{2}).$
Since $f_{1}\in I$, $x_{3}(x_{2}-x_{3})\in I$, and we may rewrite $f_{2}$ as
$f_{2}=x_{3}(x_{0}-x_{1})+2ux_{2}x_{3}(x_{2}-x_{3})\in I\Rightarrow x_{3}(x_{0}-x_{1})\in I,$
and therefore $x_{3}\in I$. Now, turning to $f_{3}\in I$ and removing terms with $x_{3}$, we obtain:
$f_{3}\equiv x_{2}(x_{0}-x_{1})\Rightarrow x_{2}\in I,$
since $x_{0}-x_{1}$ is invertible. Thus, $I\simeq\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
Over $\mathfrak{p}_{2}=((x_{0}-x_{1}),x_{3})$, we may choose the same elementary operations and thus
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{I}$
with $I=(f_{1},f_{2},f_{3})$, the polynomials as before. Since $f_{1}\in I$, we conclude:
$f_{3}=(x_{0}-x_{1})(x_{2}-2x_{3})+2ux_{3}^{2}(x_{2}-x_{3})\Rightarrow(x_{0}-x_{1})(x_{2}-2x_{3})\in I.$
Since $(x_{2}-2x_{3})$ is invertible in $R_{\mathfrak{p}_{2}}$, we obtain $(x_{0}-x_{1})\in I$. From this fact, since $x_{2}-x_{3}$ is invertible and $f_{1}\in R_{\mathfrak{p}_{2}}$, we obtain $x_{3}\in I$, and thus $I\simeq\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
Over the prime $\mathfrak{p}_{3}$, we may also do the same elementary operations and consider the same generators for the ideal $I$. From $f_{1}\in I$ we get $(x_{2}-x_{3})\in I$, since $x_{3}$ is now invertible. Moreover, from $f_{2}\in I$ we obtain:
$f_{2}=x_{3}(x_{0}-x_{1})+2ux_{2}x_{3}(x_{2}-x_{3})\Rightarrow x_{3}(x_{0}-x_{1})\in I\Rightarrow(x_{0}-x_{1})\in I,$
so that $I\simeq\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=1$. Thus, from the associativity formula $m(\sigma)=3$. Moreover, $e=2$ and thus $\operatorname{Bour}(\sigma)=2$.
Here, the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)^{2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $c_{3}({\mathcal{T}}_{\sigma})=4$, and the associated Bourbaki scheme is a plane smooth quadric.
Corollary 60.
Let $\sigma$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since $\mu({\mathcal{T}}_{\sigma})=-3/2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable if and only if $e>1$. For $m(\sigma)<3$, $\operatorname{Bour}(\sigma)>2$ and by Theorem B Theorem B, $(a)$, $e\neq 1$. ∎
Proposition 61.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)=2$, then $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, $c_{3}\in\{1,3,5,7\}$ and $c_{3}=7$ if and only if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, with $\operatorname{Bour}(\sigma)=3$ and $p_{a}(B)=0$ (see Example 62).
Proof.
If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [undefn, Section 9]. By the proof of [undefn, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
From Remark 6, we obtain $p_{a}(B)=1$, and thus $B$ must be a plane cubic curve, with a resolution of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
which by Lemma 12 yields a resolution
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [undefn, Lemma $9.3$]. If $c_{3}=7$, then ${\mathcal{T}}_{\sigma}(1)$ is stable with Chern classes (1,3,7)(-1,3,7). From [undefb, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$. ∎
Example 62 ($m(\sigma)=2,e=2$, $\operatorname{Bour}(\sigma)=3$ , $c_{3}=7$).
Consider the following sequence with $d_{f}=1$, $d_{g}=2$:
$\sigma=(-x_{0}x_{1}+x_{1}x_{2}-x_{2}x_{3},x_{0}x_{1}^{2}+x_{2}^{3}+x_{2}^{2}x_{3}),$
with corresponding Jacobian matrix given by:
$\nabla\sigma=\begin{pmatrix}-x_{1}&x_{2}-x_{0}&x_{1}-x_{3}&-x_{2}\\ x_{1}^{2}&2x_{0}x_{1}&3x_{2}^{2}+2x_{2}x_{3}&x_{2}^{2}\end{pmatrix}$.
The irreducible components of $(\Xi_{\sigma})_{1}$ consist only of a line $V(x_{1},x_{2})$. Consider $\mathfrak{p}=(x_{1},x_{2})$ the corresponding prime ideal in $R$, and let $u=(x_{2}-x_{0})^{-1}\in R_{\mathfrak{p}}$. The elementary operations $R_{2}^{\prime}=R_{2}-2ux_{0}x_{1}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ turn the matrix into
$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&1&u(x_{1}-x_{3})&-ux_{2}\\ x_{1}^{2}+2ux_{0}x_{1}^{2}&0&x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})&x_{2}^{2}+2ux_{0}x_{1}x_{2}\end{pmatrix}$,
which sends $e_{2}\in R_{\mathfrak{p}}^{4}$ to $(1,0)\in R_{\mathfrak{p}}^{2}$, so we may compute the cokernel $(\mathcal{Q}_{\sigma})_{\mathfrak{p}}$ by considering the second row and the remaining entries:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I},$
where
$f_{1}$ $=x_{1}^{2}(1+2ux_{0})$
$f_{2}$ $=x_{2}(3x_{2}+2x_{3})-2ux_{0}x_{1}(x_{1}-x_{3})$
$f_{3}$ $=x_{2}(x_{2}+2ux_{0}x_{1}).$
Since $1+2ux_{0}\in R_{\mathfrak{p}}^{\times}$ and $f_{1}\in I$, we conclude $x_{1}^{2}\in I$. This simplifies $f_{2}\in I$ to:
$x_{2}(3x_{2}+2x_{3})\equiv-(2ux_{0}x_{3})x_{1}\mod I,$
so there is a unit $v\in R_{\mathfrak{p}}^{\times}$ such that $x_{1}\equiv vx_{2}\mod I$, which together with $x_{1}^{2}\in I$ implies $x_{2}^{2}\in I$. From this, $f_{3}\in I$ simplifies to
$f\equiv 2ux_{0}x_{1}x_{2}\in I\Rightarrow x_{1}x_{2}\in I,$
and therefore we have
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{1}\rangle,$
with $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=2$. Thus, $m(\sigma)=2$ (from the associativity formula), and $e=2$, yielding $\operatorname{Bour}(\sigma)=3$. The following resolution may also be obtained computationally:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 2}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 3}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
describing the Bourbaki scheme as a twisted cubic and giving $c_{3}({\mathcal{T}}_{\sigma})=7$.
Example 63 ($m(\sigma)=2$, $e=3$, $\operatorname{Bour}(\sigma)=5$, $c_{3}=3$).
We consider the following sequence:
$\sigma=(x_{2}x_{3}-x_{0}x_{1},x_{0}^{2}x_{2}+x_{0}x_{1}x_{3}+x_{2}x_{3}^{2}+x_{3}^{3})$
with corresponding Jacobian matrix given by
$\nabla\sigma=\begin{pmatrix}-x_{1}&-x_{0}&x_{3}&x_{2}\\ 2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}&x_{0}^{2}+x_{3}^{2}&x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2}\end{pmatrix}$.
Here, there is a unique irreducible component of $(\Xi_{\sigma})_{1}$, the line $V(x_{0},x_{3})$. Considering $\mathfrak{p}=(x_{0},x_{3})$ the associated prime ideal, we compute $m(\sigma)=2$ using the associativity formula.
Take $u=(x_{2})^{-1}\in R_{\mathfrak{p}}$ and consider the elementary operations $R_{2}^{\prime}=R_{2}-u(x_{0}x_{1}+2x_{2}x_{3}+3x_{3}^{2})R_{1}\doteq R_{2}-uQR_{1}$, $R_{1}^{\prime}=uR_{1}$ so that
$(\nabla\sigma)_{\mathfrak{p}}\sim\begin{pmatrix}-ux_{1}&-ux_{0}&ux_{3}&1\\ 2x_{0}x_{2}+x_{1}x_{3}+uQx_{1}&x_{0}x_{3}+uQx_{0}&x_{0}^{2}+x_{3}^{2}-uQx_{3}&0\end{pmatrix}$.
Arguing as usual, we may compute the cokernel by
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\frac{R_{\mathfrak{p}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}}}{I}$
with
$f_{1}$ $=x_{0}(2x_{2}+ux_{1}^{2})+x_{1}x_{3}(2ux_{2}+3ux_{3}+1)$
$f_{2}$ $=x_{0}(x_{3}(2ux_{2}+3ux_{3}+1)+ux_{0}x_{1})$
$f_{3}$ $=x_{0}^{2}+x_{3}(x_{3}-ux_{0}x_{1}-2ux_{2}x_{3}-3ux_{3}^{2})$
Writing
$f_{2}-\frac{x_{0}}{x_{1}}f_{1}$ $=x_{0}\left(-\frac{x_{0}}{x_{1}}(2x_{2}+ux_{1}^{2})+ux_{0}x_{1}\right)$
$=x_{0}^{2}\left(-\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)\in I,$
and since the term inside parenthesis is invertible in $R_{\mathfrak{p}}$, we conclude $x_{0}^{2}\in I$.
On the other hand, we consider
$B\doteq f_{3}-x_{0}^{2}+\frac{x_{3}}{x_{1}}f_{1}$ $=-x_{3}\left(x_{0}\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{0}x_{1}-2x_{3}\right)$
$=-x_{3}x_{0}\left(\frac{(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right)+2x_{3}^{2}\in I.$
and moreover, let
$A\doteq x_{0}x_{3}(2ux_{2}+3ux_{3}+1)=f_{2}-ux_{0}^{2}x_{1}x_{3}\in I.$
Then
$A-\frac{3}{2}ux_{0}B$ $=x_{0}x_{3}(2ux_{2}+1)+3x_{3}^{2}-3x_{3}^{2}+\frac{3}{2}ux_{0}^{2}x_{3}\left(\frac{-(2x_{2}+ux_{1}^{2})}{x_{1}}+ux_{1}\right),$
and after removing the term with $x_{0}^{2}$ factor, we obtain
$x_{0}x_{3}(2ux_{2}+1)\in I\Rightarrow x_{0}x_{3}\in I,$
since $2ux_{2}+1$ is invertible in $R_{\mathfrak{p}}$. Now, removing the $x_{0}x_{3}$ term in $B$ we conclude $x_{3}^{2}\in I$. So, up until now, we have $x_{0}^{2},x_{0}x_{3}$ and $x_{3}^{2}\in I$. Finally, considering
$f_{1}-3ux_{1}x_{3}^{2}=x_{0}(2x_{2}+ux_{1}^{2})+x_{3}x_{1}(2ux_{2}+1)\in I,$
this means there are $a,b\in R_{\mathfrak{p}}^{\times}$ such that $ax_{0}+bx_{3}\in I$. Therefore, we conclude that
$I=(x_{0}^{2},x_{3}^{2},x_{0}x_{3},ax_{0}+bx_{3}),$
and these mean that we can write
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}}\simeq\kappa\langle 1,x_{0}\rangle\simeq\langle 1,x_{3}\rangle$
and obtain that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}}=1$. Moreover, $e=3$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-6)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 3}\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so we obtain $\operatorname{Bour}(\sigma)=5$ and $c_{3}({\mathcal{T}}_{\sigma})=3$.
We summarize the results of this subsection in the following theorem:
Theorem D.
Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
    $m(\sigma)\leq 7$ and equality holds if and only if $\sigma$ is compressible;
  • (b)
    The sequence $\sigma$ is free if and only if $m(\sigma)=7$ or $5$, and each corresponds to $e$ being 0 or $1$, respectively;
  • (c)
    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 56 and Example 14);
  • (d)
    If $m(\sigma)=3$ then $\operatorname{Bour}(\sigma)=2$, with a possible unstable case $(c_{3}=6)$ and stable cases with $c_{3}\in\{0,2,4\}$. Among these, we have examples for the stable cases with $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
  • (e)
    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Item $(a)$ is 48 and item $(b)$ follows with 53. Item $(c)$ is shown in $5$, and item $(d)$ is described in 57. The stability result in $(e)$ is in 60. ∎